### v1

# Are EU low-carbon structural funds efficient in reducing emissions?

Abstract
This paper investigates the effectiveness of the “low-carbon economy” expenditures from European Structural and Investment Funds in fostering reductions in greenhouse gas emissions within European regions, focusing on the 2007-2013 and 2014-2020 programme periods. By decomposing emissions time series into trend and cycle components and considering them within a panel data framework, our research highlights that the impacts of low-carbon economy expenditures vary, qualitatively and quantitatively, with the targeted regions’ development level. We find significant emissions reductions in developed and transition regions yet less favourable outcomes in less developed areas. Further analysis into specific greenhouse gas emissions types (CO2, CH4, and N2O) reveals inconsistent impacts, underscoring the complexity of achieving emissions reductions. Our findings emphasise the need for tailored environmental strategies that accommodate the economic disparities of regions in the European Union.
Keywords: Greenhouse gas emissions; Emission decomposition; Low-carbon economy; European Structural and Investment Funds; Regional development
1 Introduction
The European Union (EU) has the objective of achieving climate-neutrality by 2050. This implies large-scale investments in the energy transition (McCollum et al., 2018; Klaassen and Steffen, 2023). According to the European Commission, energy investments in the EU will have to reach €396 billion per year from 2021 to 2030 and €520-575 billion per year in the subsequent decades until 2050 (European Parliament, 2023). A substantial share of these ought to be public investments that foster the development of infrastructure, the deployment of technological innovations, and the scale-up of private capital (Carraro et al., 2012; Owen et al., 2018; Meckling et al., 2022). Accordingly, the EU has enacted climate spending targets amounting to 20 of its 2014-2020 budget and 30 of its 2021-2027 budget that ought to be devoted to climate action111See European Court of Auditors (2022), which also emphasises that climate spending over the 2014-2020 period has been over-estimated.. In this context, one can ask a necessary question about the efficiency of European climate policy: have these expenses effectively led to the reduction of greenhouse gas (GHG) emissions?
This paper exploits regional variations in European climate spending to address this issue. The main investment vehicles of the EU are the European Structural and Investment Funds (ESI funds), which accounted for €535 billion over the period 2014-2020 (European Commission, 2023). By construction, these funds are spent heterogeneously. EU regions are grouped into three categories based on their relative development level (least developed regions, transition regions, and developed regions), and this determines their eligibility and their financing rate for structural funds. From the 2014-2020 programming period onwards, the usage of these funds has been tagged with a “thematic objective”. In particular, the “Low-Carbon Economy” tag allows the identification of the funds that have been assigned to climate action. Although the programming period 2007-2013 did not define thematic objectives in the same way, the aims of the expenditures are also tagged by “priority codes”, from which it is possible to determine the funds related to climate action consistently with the low-carbon economy thematic objective from 2014-2020.
Consequently, we use the total amount spent on the low-carbon economy objective in the last two budgeting periods (2007-2013 and 2014-2020) as a measure of public climate investment. We then employ an econometric panel data model to investigate the influence of these investments on GHG emissions in European regions between 2007 and 2022 at the NUTS 2 level. Our methodology involves decomposing the emission time series of each region into trend and cycle components. This strategy gives us deeper insights into the nexus between economic growth and environmental sustainability. Due to the heterogeneity of European regions, it is clear that some have significantly advanced in terms of sustainability, achieving various degrees of decoupling, whereas many others have yet to do so. This disparity must be accounted for in the analysis to determine such expenditures’ short-term and long-term effects. Our analysis also accounts for the direct impacts of policy by considering the OECD Environmental Policy Stringency (EPS) index (OECD, 2016) and mean regional political opinions as covariates.
We find heterogeneous impacts of EU structural fund investments in low-carbon economies on GHG emissions reductions across European regions, with findings highlighting the variability of these effects by regional development levels, investment periods, and emission types. In developed and transition regions, low-carbon investments significantly lowered the long-term emission trends, in both per capita and relative to GDP, across the 2007-2013 and 2014-2020 programs. However, in the least developed regions, “low-carbon” investments had a positive (and significant) impact on emissions. Additionally, the effectiveness of reducing specific gas emissions, including CO2, CH4, and N2O, exhibits variability, indicating a complex interaction between investment efficiency and pollutant type. Our analysis also discerns significant but divergent impacts on the cyclical emissions components in transition economies between the two program periods. Hence, our analysis suggests a partial inefficiency of low-carbon public investments in Europe. It calls for a refined spatial and temporal allocation of these investments.
The heterogeneous impacts of the climate action efforts derived from the ESI funds can be attributed to several critical factors. These factors include the inherent complexities of reducing emissions, the intrinsic differences among European regions in terms of capabilities and resource endowments, and the relevance of development paths, as suggested by the Kuznets environmental curve (Bataille et al., 2021; Saraji and Streimikiene, 2023). For instance, it is possible that any additional funding may increase energy consumption and GHG emissions for investments below a certain level. This is rooted in how investments are targeted and their immediate impacts, which might not always be aligned with long-term sustainability goals. In many developing regions, the primary focus of initial investments may be on establishing basic infrastructure and boosting economic output quickly to meet immediate developmental needs. Additionally, low-carbon economy investments typically require substantial up-front costs and the deployment of heavy infrastructure, which can drive significant energy consumption and emissions. The transition to sustainable energy systems includes upgrading grids, developing renewable energy sources, and improving energy resilience. Though essential in the long term, these investments can increase short-term emissions due to the intensive energy and resources required during their implementation phase. On another front, there is a risk that investments have been used inefficiently or inappropriately, for example, as a countercyclical buffer to stabilise the economy during recessions or by investing in large infrastructure projects (aka “white elephants”) that do not generate the expected impacts. Arguments like these could align with the heterogeneous effects of public investment on climate outcomes.
Therefore, the differential impacts on GHG emissions have their origin in the cyclical nature of economic activities. Recent literature offers valuable insights into the interplay between environmental impacts and the business cycle. Doda (2014) finds that CO2 emissions increase during economic upturns, exhibit greater volatility compared to GDP, and show a procyclicality that is more pronounced in higher GDP per capita economies, suggesting wealthier economies have more stable emission patterns through economic cycles. Sheldon (2017) finds an asymmetrical response of emissions to economic changes attributed to the timing of capital investments and technological advancements, indicating the significant role of corporate decisions in environmental impacts. Moreover, Klarl (2020) finds CO2 emissions’ elasticity to GDP changes to be significantly higher during recessions in the United States, showcasing the dynamic elasticity of emissions in response to economic conditions.
In addition, climate and energy policy tools have proven some degree of effectiveness, but their success often hinges on countries’ stringency and the complementarity among strategies. Diakoulaki and Mandaraka (2007) find that despite considerable efforts, the expected acceleration in decoupling post-Kyoto Protocol was not uniformly achieved. For Europe, Naqvi and Zwickl (2017) and Ivanova et al. (2017) also observe significant disparities in decoupling across countries, sectors, pollutants, and regions. Papież et al. (2022, 2021) further report mixed results in the EU’s long-term decoupling efforts, pointing to disparities between older and newer member states and emphasising the role of energy intensity and structure in emissions reductions. For OECD countries, Chen et al. (2018), focusing on the key factors affecting CO2 emissions, conclude that while reductions in energy intensity are associated with decreased CO2 emissions, economic growth inevitably tends to increase CO2 emissions. This implies that emissions reductions are hard to accomplish unless economic growth is accompanied by significant improvements in energy efficiency and a shift towards less carbon-intensive energy sources.
Based on the literature discussed above, our paper contributes to the analysis of sustainable development and economic policy in European regions. By analysing how low-carbon investments from structural funds impact emissions reductions within business cycle fluctuations, we contribute to the understanding of decoupling economic growth from environmental degradation. Our results emphasise the importance of tailored investment strategies in achieving sustainable development goals, with a special focus on less developed regions to mitigate potential increases in emissions as they grow economically.
The remainder of this study is structured as follows. Section 2 discusses the relevant strands of literature and the conceptual framework of our study and situates our contribution. Section 3 presents the data and methods. Section 4 presents the empirical results of the econometric estimations. Finally, Section 5 concludes by discussing the policy implications of our results.
2 Decoupling emissions: related literature
Our study builds on the extensive literature analysing the relationship between economic growth and GHG emissions. The Environmental Kuznets Hypothesis indicates that there is a connection between a country’s income level and the quality of its environment: it can be said that an increase in wealth generates greater demand from the population for a cleaner and healthier environment. This demand, in turn, drives governments to implement and enforce stricter environmental laws and regulations to protect the environment. While wealthier countries tend to have the resources and political will to adopt and enforce these effective environmental policies, regions with more constrained economic resources may find it more challenging to establish and maintain high environmental standards, resulting in more significant environmental challenges (O’Connor and Turnham, 1992; Doyle and Stiglitz, 2014).
Certainly, economic development can lead to improved environmental outcomes by decoupling emissions from economic output. This is central to sustainable development and indicates that an economy can grow without a proportional increase in environmental degradation. However, the process is influenced by a multitude of factors, not only including income levels but also other aspects such as consumption patterns and regional characteristics. Shuai et al. (2017) analysed the country-by-country turning points of the carbon Kuznets curve, finding a positive correlation between a country’s income level and its likelihood to fit the Kuznets hypothesis, with higher income economies achieving more quick tipping points than their poorer counterparts. While economic development level is linked to faster environmental improvement transitions, other variables also play a relevant role. Ivanova et al. (2017) analyses household consumption across 177 regions in 27 EU countries, finding wide variations in per capita GHG emissions across regions (0.6 to 6.5 tCO2e/cap in 2007). Regarding the drivers behind carbon footprints, the authors identified the relevance of household size, whether an area is urban or rural, the population’s education level, expenditure patterns, local temperature, resource availability, and the carbon intensity of the local electricity mix.
Yet a significant strand of literature doubts that the downward-sloping relationships between higher-income countries and pollution levels – which might imply a decoupling of the economic activity with emissions – is enough to face climate change efficiently. Certainly, there are concerns about the possible shift in the production of pollution-intensive goods from developed to developing countries with less stringent environmental regulations (Hertwich and Peters, 2009; Levinson, 2023). This would weaken the Kuznets hypothesis since developed regions may reduce their at-home pollution levels by offshoring the production of environmentally harmful goods to nations where environmental regulations are less strict (Fischer and Heutel, 2013). While this points out that high-income countries should also look into reducing aggregate demand as a strategy for significant emissions reductions (Knight and Schor, 2014), it also requires the implementation of agreements at different scales: regional and global (Oberthür and Ott, 1999).
A growing body of evidence indicates that such policies and international agreements are increasingly impacting emission trends, as reflected in the economic practices within countries, contributing to our understanding of the decoupling mechanism (Fischer and Heutel, 2013; Hildén et al., 2014; Haberl et al., 2020). Decoupling analyses aim to categorise the patterns of how emissions respond to economic growth – whether they increase at a slower rate, decrease, or stabilise as the economy grows (relative and absolute decoupling) or increase faster than the economy (coupling).
Cohen et al. (2018) provide valuable quantitative analysis that helps recognise how major emitting countries are progressing in decoupling emissions. Using information from OECD countries for 2001-2015, Chen et al. (2018) analyse how economic growth can be decoupled from CO2 emissions, identifying three key forms: recessive, where emissions drop due to efficiency gains or energy mix shifts, not economic growth; weak negative, where efficiency improvements do not fully offset emissions from economic expansion; and strong, where economic and population growth do not lead to higher emissions, thanks to effective policies, technological progress, and structural economic changes.
Building on this understanding of the decoupling patterns, it is imperative to analyse emissions in the context of business cycles due to their substantial implications for environmental policy and economic management, particularly regarding the relationship between economic activity, regulation, emissions and energy consumption (Fischer and Heutel, 2013). In fact, emissions tend to be procyclical, increasing during economic booms and decreasing during recessions (Heutel, 2012). A fundamental question relates to the causal relationship between economic output and energy consumption, especially targeting business cycle fluctuations (Thoma, 2004; Narayan et al., 2011). Energy consumption notably responds to changes in industrial production and macroeconomic conditions, reinforcing the concept of the procyclicality of emissions relative to GDP observed in broader research. For instance, Thoma (2004) demonstrates that changes in macroeconomic conditions cause significant changes in electrical usage, particularly in the commercial and industrial sectors.
Therefore, the link between economic growth, energy consumption, and GHG emissions is clearly established in the short term. For example, Peters et al. (2012) claim that economic downturns typically lead to a reduction in energy-intensive activities and lower CO2 emissions. Notwithstanding, the authors discuss that the 2008-2009 financial crisis saw a quick dip in GDP followed by a swift rebound in emissions by 2010, driven by falling energy prices, government economic incentives, and prior economic growth in developing countries, quickly returning to high emission levels post-crisis. Similarly, Doda (2014) finds that CO2 emissions are procyclical and more pronounced in countries with higher GDP per capita. Hence, the procyclicality also implies that economic growth without specific interventions often leads to increased emissions. Another fact is that emissions exhibit greater volatility compared to GDP, responding more dramatically to economic changes, and the extent to which emissions volatility exceeds GDP volatility decreases with the income level of the economies (Doda, 2014). This implies that more stable emission patterns through economic cycles in developed regions could be linked to more efficient energy use, a greater share of services in the economy, and potentially more effective environmental regulations and policies. In addition, emissions decrease more significantly during economic contractions than during expansions, suggesting an elastic response to economic downturns and inelasticity to economic growth (Sheldon, 2017).
Likewise, Klarl (2020) examines the fluctuation of CO2 emissions’ elasticity to GDP in the United States from 1973 to 2015, revealing that this elasticity is not constant and varies with economic conditions. He showed that emissions respond more sensitively to GDP changes during recessions than periods of economic normality, with elasticity often exceeding one in downturns. This indicates that CO2 emissions decrease more rapidly than economic output in recessions, highlighting the impact of economic cycles on environmental outcomes. The finding that emissions exhibit greater sensitivity during economic downturns suggests that the capacity of an economy to influence environmental outcomes can be significantly conditioned by its current economic state.
In a similar vein, Naqvi (2021) finds that, on average, EU regions have managed to decouple emissions from economic output, with the most significant emission reductions occurring before the 2008 financial crisis. After 2008, the trend towards decoupling weakened, with instances of re-coupling between emissions and economic output observed. The author uses the EPS index, finding that stronger environmental policies effectively reduce emissions. However, the effectiveness varies significantly across different types of emissions and is also influenced by the income levels of the regions. Similarly, with the aim of monitoring the effectiveness of the EU’s environmental policies, Delgado Rodríguez et al. (2018) developed the EU-cyclical environmental performance index by applying dynamic factor analysis over the period 1950 to 2012. This index demonstrates that the implementation of stricter emission targets, initiated particularly since the adoption of the 5th Environmental Action Programme (EAP), has effectively contributed to restraining the growth of CO2 emissions across the EU. Furthermore, the index reveals heterogeneity among the member states, distinguishing between countries that are leading in emission linkage – thus enhancing their alignment with EU environmental standards – and those lagging behind, which may require additional incentives or policy measures to improve compliance with EU regulations. In addition, Papież et al. (2022) evaluate the EU’s energy policy effectiveness in achieving long-term decoupling and reports mixed findings. The study notes progress in detaching economic growth from emissions under the EU ETS and production-based accounting in both EU-15 and new member countries. However, they find no advancement in separating economic growth from consumption-based accounting emissions in either group.
In summary, environmental policies must confront the broader economy, influenced by external macroeconomic shocks and internal technological changes. Under Kuznets’ hypothesis, decoupling might appear problematic since countries would only begin to improve their environmental performance after reaching a certain level of wealth, implying a dependency between economic growth and environmental improvement. Therefore, seeking economic development without the corresponding environmental degradation is a great challenge for any policy framework.
3 Data and methods
We focus on emission data reported by the Emissions Database for Global Atmospheric Research (EDGAR), managed by the Joint Research Centre of the European Commission. EDGAR provides quantitative estimates of global GHG emissions by country and sector, including CO2, CH4, N2O, and F-gases, using a methodology aligned with Intergovernmental Panel on Climate Change (IPCC) guidelines (Crippa et al., 2023). Specifically, we use European regional data at the NUTS 2 level of aggregation, focusing on aggregated GHG emissions and their breakdown into CO2, CH4, and N2O emissions.
Our primary objective is to analyse the impacts on GHG emissions resulting from expenditure on the thematic objective related to the low-carbon economy in the 2007-2013 and 2014-2020 ESI funds programmes (a more detailed definition of the expenditure variables is presented below). In the econometric modelling setup, we use other relevant covariates to control for economic activity, sectoral composition, and local policy stringency that influence emissions across European regions.
3.1 Data
ESI funds and eligibility
We use data from the Cohesion Open Data Platform, which provides a granular view of the thematic content financed during the 2007-2013 and 2014-2020 programmes. The datasets for both periods encompass cumulative categorisation data on projects as reported in the European Regional Development Fund (ERDF), Cohesion Fund (CF), and European Social Fund (ESF) programme final implementation reports as presented in the closure documentation.222For 2007-2013 raw categorisation data see: https://cohesiondata.ec.europa.eu/2007-2013-Categorisation/2007-2013-Cohesion-categorisation-raw-data-from-cl/qt3c-wiyg/about_data.
For 2014-2020 raw categorisation data see: https://cohesiondata.ec.europa.eu/2014-2020-Categorisation/ESIF-2014-2020-ERDF-CF-ESF-raw-categorisation/xe4p-7b9q/about_data. These data contain the volume of EU support allocated to projects alongside a combination of categorisation dimension codes, including priority themes, the form of finance, the territorial context, location (using the Nomenclature of Territorial Units for Statistics, NUTS), and the economic dimension. We aggregate investments at NUTS level 2 (NUTS 2), as defined in version 2021. To do this, we consider all related investments, regardless of their original level of geographic specificity in the original database – be it national, NUTS-1, or even more granular at NUTS-3.
Figure 1 shows the calculated distribution of total ESI funds expenditures reported by region for the periods 2007-2013 and 2014-2020, along with the eligibility of the regions to access these funds. Eligibility is based on an income classification of regions, categorising them into three groups: LDR (less developed regions), TER (transition regions), and DER (developed regions). In all econometric exercises in this paper, we use this regional income categorisation fixed.333The details of the eligibility classification, according to the ESI funds assigned in 2014-2020, can be found here: https://eur-lex.europa.eu/legal-content/EN/TXT/?uri=CELEX:32014D0099.
The total expenditures reflect the targeted efforts to reduce economic disparities within the EU. Certain regions consistently receive higher funds according to ongoing economic development needs or projects requiring sustained investment. The disparities in the distribution of funds consistently with eligibility show that less developed regions in Southern and Eastern Europe generally receive higher per capita funding than Western and Northern Europe regions.
There are also variations in the distribution of funds between programming periods, which must be explained by changes in policy focus. Clearly, the EU’s structural funds programmes significantly evolved from their initial focus on procuring the catching-up of less developed regions towards a more strategic approach. In the 2007-2013 programming period, a transition towards integrating environmental and climatic considerations began, emphasising energy efficiency, renewable energies, and sustainable resource management. However, the transition to a more strategic approach was cemented during the 2014-2020 programming period, with a clear directive to channel funds towards smart, sustainable, and inclusive growth.
The 2014-2020 funding period introduced thematic objectives (TOs), including: Research and Innovation (TO-1), Information and Communication Technology (TO-2), Competitiveness of Small and Medium-Sized Enterprises (TO-3), Low-Carbon Economy (TO-4), Climate Change Adaptation, Risk Prevention, and Management (TO-5), Environmental Protection and Resource Efficiency (TO-6), Sustainable Transport and Removing Bottlenecks in Key Network Infrastructures (TO-7), Employment and Labour Mobility (TO-8), Social Inclusion and Combating Poverty (TO-9), Education, Training, and Vocational Training for Skills and Lifelong Learning (TO-10), and Institutional Capacity Building and Efficient Public Administration (TO-11).
Low-carbon economy funds
We define as “climate action funds” all expenditures aimed at reducing emissions in both programming periods. Despite thematic changes between the two periods, the raw categorisation data allows for the creation of a comparative set of expenditure objectives to ensure comparability and coherence in our analysis.
For the 2014-2020 period, the low-carbon expenditures are straightforward to determine since there is a categorical variable related to thematic objective 4 (TO-4), labelled as “low-carbon economy”. The aim of the expenditure can be tracked by the reported “intervention field” in the data. Specifically, related to TO-4, the intervention fields include renewable energy (wind/solar/biomass), energy efficiency renovation of public infrastructure, cycle tracks and footpaths, among others (see Table A1 in the Appendix).
For the 2007-2013 program, there is no categorical variable explicitly associated with the low-carbon economy as in the 2014-2020 program. However, the raw categorisation data describes the “priority” related to the reported expenditures, serving as a proxy for the aim of the expenditure. We use this information to create a low-carbon expenditures variable for this period. Priority labels related to this period include cycle tracks, intelligent transport systems, renewable energy (wind/solar/biomass/hydroelectric), energy efficiency, co-generation, energy management, and promotion of clean urban transport. The priority codes are reported in Table A1 of the Appendix.
A critical aspect of the ESI funds expenditure data is the absence of detailed annual expenditure information at the regional level. While reported expenditures can be easily tracked at the regional level, the reports correspond to the entire programming period. To our knowledge, there is no disaggregated, time-varying consolidated official information. Nevertheless, the cohesion data platform presents estimations of the expenditure by NUTS 2, which are aggregated for the most significant types of structural funds only. Therefore, to address our need for yearly data, we have opted for two approaches: one assumes that expenditures are distributed evenly over time (dividing the reported expenditures over the programming period), and the other assumes they follow the same pattern as the total aggregated expenditure reported by the cohesion data platform.444Data available at: https://cohesiondata.ec.europa.eu/Other/Historic-EU-payments-regionalised-and-modelled/tc55-7ysv/about_data. Since both strategies yielded qualitatively similar conclusions, we present in the main text the results obtained with the first approach, while those derived from the second are included in the Appendix as a robustness check (Tables A2 and A3).
Supplementary covariates
Economic variables were obtained from the ARDECO (the Annual Regional Database of the European Commission’s Directorate General for Regional and Urban Policy), which contains annual data at the regional level and is maintained and updated by the Joint Research Centre of the European Commission. We use variables such as the real gross domestic product (GDP), the population, and the real gross value added (GVA) for all industry sectors according to NACE Rev. 2 definitions. These sectors are A: Agriculture, Forestry and Fishing; B–E: Industry; F: Construction; G–J: Wholesale, Retail, Transport, Accommodation & Food Services, Information and Communication; K–N: Financial & Business Services; and O–U: Non-market Services. The choice of these variables is motivated by their relevance to understanding how economic activity is related to emissions. In particular, we include industry-level GVA to control direct sources of emissions and to better capture regional emission profiles, given that the participation of industries varies significantly across European regions.
To assess the effects of environmental policies on emissions, we use the EPS index (OECD, 2016). This index, developed by Brunel and Levinson (2013) and further refined by Botta and Koźluk (2014), provides a standardised composite measure incorporating both market and non-market-based policies, covering 27 OECD countries over the period from 1990 to 2020.
Finally, to determine the political ideology at the regional level, we calculate the regional weighted average of political ideologies using the results of European subnational election data. To measure political ideology, we use the “left/right” dimension variable reported by ParlGov (Döring et al., 2022) and complemented by CHES-Europe (Jolly et al., 2022), which takes values in the interval 0–10, with values close to zero representing extreme left ideology and values close to 10 representing extreme right ideology. We use EU-NED (Schraff et al., 2023), which reports sub-national election data across European countries over the last three decades at the party level. We assign each party at the regional level an ideology by combining information from ParlGov and CHES-Europe. These databases are integrated within the Party Facts platform, which facilitates the incorporation of party-level data.555The data is available at: https://www.chesdata.eu/ches-europe and https://www.parlgov.org/. Also see: https://partyfacts.herokuapp.com/. We include this variable to provide additional controls for political ideology. While our focus is not on analysing how politics affects emissions, this is a growing area of interest in political science. For instance, Jahn (2021) highlights regional variations in the impact of populist governance on environmental policies, showing that right-wing populist administrations in North Western and Eastern Europe are linked to heightened GHG emissions, whereas their left-wing counterparts in Southern Europe are associated with decreases in emissions.
Distribution of emissions and low-carbon economy funds
Figure 2 presents a set of maps illustrating different facets of the environmental impact and distribution of the Cohesion Funds. In the upper part, three maps show average GHG emissions in shades of blue to red, where blue denotes lower levels and red denotes higher levels. The first map, on the left, represents the average emissions for 2015-2019 in teragrams of CO2 equivalent per year (Tg CO2eq/year). The other two maps compare the average annual per capita emissions for 2000-2004 and 2015-2019, respectively. These maps show that, on average, towards the end of the 2010s, most European regions emitted less than 50 teragrams of CO2. However, there is significant heterogeneity in per capita emissions, highlighting areas of intense economic activity or low population density. Comparing the early 2000s and late 2010s, it is evident that many regions with initially high emissions tended to reduce them (see, for instance, the Castile region in Spain), while some regions maintained or increased their per capita emission levels (for instance, the Masovian region in Poland).
The lower maps in Figure 2 show the calculated funds allocated to promote a low-carbon economy during 2007-2013 and 2014-2020 programmes, with the amounts received per capita in colour scales ranging from less than 100 euros to almost 1,000 euros per person. In both periods, most of the funds went to LDRs, with significant differences in the allocated amount, in agreement with the results presented in Figure 1.
This pattern raises questions about the effectiveness of the allocation of funds relative to emission levels. In Figure 3, we plot the total GHG emissions per capita versus the total expenditure in low-carbon economy funds per capita. Emissions are considered at the beginning of the corresponding programming period. The plot exhibits great moderation in emission levels per capita for most European regions and shows that higher funds per capita were consistently allocated to less developed regions. However, a positive correlation pattern in the allocation is less clearly seen in both panels, suggesting that these funds are not necessarily allocated to the regions with the highest need to reduce emissions.
As a preliminary analysis, we estimate the nexus between expenditure on low-carbon economy funds and total GHG emissions using the following regional-level model:
where is an emission indicator for region at time , and and are the cumulative yearly programme expenditures per capita (in log) for the periods 2007-2013 and 2014-2020, respectively, with denoting lags to acknowledge potential delayed effects. We include fixed effects for location () and time () to control for time-invariant unobserved heterogeneity across regions as well as uniform temporal shocks, such as those associated with environmental directives instituted at the European Union level. Finally, the term is the estimation residual.
In the left-hand in Eq. (1), the term serves as a reference notation for different emission variables. Specifically, we use four different dependent variables based on the total emissions of region at time . These are the levels of emissions per capita (in log), defined as GHGEpci,t=; the emission intensity relative to the GDP, defined as GHGEii,t=; and the growth rates of GHGEpci,t and GHGEii,t.
The estimation results are presented in Table 1. We find that expenditures have a statistically significant positive relationship with per capita emissions, as seen in models 1 and 2, indicating a perverse effect against the intended impact of such investments. When examining the growth rates of emissions per capita, a negative relationship appears with the three-year lag in spending (model 3). However, in model 4, which accounts for a five-year lag, this relationship is not statistically significant. Regarding emission intensity, we find that low-carbon expenditures are linked to decreases in this measure, as demonstrated in models 5 and 6. At the same time, the growth rate of emission intensity does not show a significant effect related to the expenditures.
| GHGEpc | GHGEi | |||||||
| Levels (log) | Growth rates | Levels (log) | Growth rates | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | |
| Low Carbon Exp. pc (log) | 0.004 | 0.081∗∗ | 0.130∗∗ | 0.042 | ||||
| (0.063) | (0.036) | (0.064) | (0.038) | |||||
| Low Carbon Exp. pc (log) | 1.169∗∗∗ | 0.008 | 0.438∗∗∗ | 0.009 | ||||
| (0.091) | (0.053) | (0.093) | (0.055) | |||||
| Low Carbon Exp. pc (log) | 0.149∗∗ | 0.052 | 0.123∗∗ | 0.028 | ||||
| (0.062) | (0.035) | (0.063) | (0.037) | |||||
| Low Carbon Exp. pc (log) | 1.633∗∗∗ | 0.067 | 0.601∗∗∗ | 0.091 | ||||
| (0.140) | (0.080) | (0.142) | (0.084) | |||||
| Observations | 3,872 | 3,872 | 3,872 | 3,872 | 3,872 | 3,872 | 3,872 | 3,872 |
| Adjusted R2 | 0.382 | 0.375 | 0.179 | 0.178 | 0.580 | 0.580 | 0.039 | 0.039 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||||
This preliminary exercise demonstrates that the relationship between low-carbon expenditures and GHG levels, as well as the growth rates of emissions per capita (GHGEpc) and emission intensity (GHGEi), can be mixed. These mixed results suggest that the econometric specification might be masking underlying factors that could restrict the effectiveness of low-carbon funds. One relevant limiting factor is the persistent disparities among European regions. Indeed, the development level of regions is a key factor influencing the effectiveness of cohesion funds (Boldrin and Canova, 2001; Rodríguez-Pose and Fratesi, 2004). However, many aspects related to the heterogeneity of European regions can explain the mixed results. For instance, less developed regions might use these funds to build sustainable infrastructure to reduce emissions, while more developed regions, having already made progress in emissions efficiency, might experience a more moderate impact from these funds. Therefore, the empirical strategy must consider that short-term outcomes may differ from longer-term impacts. This consideration is crucial for environmental policy because it suggests that sustainability investments may be misallocated, leading to unintended or unfavourable effects on emissions.
3.2 Methodology
The research on decoupling emissions from economic activity primarily hinges on analysing the emissions-GDP elasticity, which measures how emissions change with GDP and helps understand the link between economic growth and environmental impact. The central model is as follows:
where represents the log of total emissions at time , denotes the log of real GDP, denotes these first differences operator, and represents the error term. The parameters and capture the intercept and the elasticity of emissions with respect to GDP, respectively.
However, as Cohen et al. (2018) pointed out, this approach has some important caveats. Looking at the growth of emissions in relation to real GDP growth for the world’s 20 largest emitters, the authors showed a positive emissions-output elasticity across all countries, suggesting a strong correlation between economic activity and emissions. However, by decomposing emissions and real GDP into their trend and cyclical components, they show clearer evidence of decoupling in richer nations, particularly in European countries. The trend components of this analysis, which focus on long-term movements rather than short-term fluctuations, indicate that while cyclical components are still closely tied to GDP growth, the trend components are not. For several countries, including Italy, the United Kingdom, and Germany, the trend elasticities for emissions are either close to zero or negative, suggesting that the long-term component of emissions growth has decoupled from GDP growth (Cohen et al., 2018).
Our methodology shifts from this type of analysis, which primarily assesses the impact of GDP growth on emissions, to a comparative panel data analysis focusing on the effects of expenditure in the low-carbon economy. We specifically highlight the role of cohesion fund investments in fostering low-carbon economies to curb emissions. Therefore, it is central in our methodology to decompose time series variables, such as emissions, GDP, and GVA, into their cyclical and trend components for each region independently. Then we consider these time series in two different panel data model frameworks: regional trends and regional cycles of emissions. In this way, we analyse the difference between long-term trends and cyclical effects of the impact of low-carbon economy funds on GHG emissions.
In the first phase of our methodology, we apply the Hodrick-Prescott (HP) filter to each region’s emissions per capita and emissions intensity time series to separate them into their corresponding cyclical and trend components. This method allows us to distinguish short-term fluctuations associated with the economic cycle from the long-term trajectory that reflects structural changes and persistent trends in emissions. Therefore, the total emissions time series of the region , , is decomposed in its trend component and its cycle component .
After decomposing the GHGEpc and GHGEi time series, we independently analyse the panels of regional emission trends and cycles, considering the regional development level differences. The model for emissions trends is defined as follows:
where the superscript specifies that we conduct the estimation on time series trend components. This way, is the trend of the logarithm of emissions per capita for NUTS 2 region at time , similarly is the trend component of the GDP (in log) of region at time , and and are the yearly cumulative programme expenditures (in log) at time . We include the development level of region using the dummy variable , where . The term consider additional control variables such as the dummy variables capturing the effect of stronger stringency of environmental policies, denoted by 666Following Naqvi (2021), we measure the impact of stronger environmental policies employing a dummy variable that takes the value of one when EPS index is greater or equal than 2.5., and the term , which indicates the average political ideology of the region, determined by the outcomes of local elections (see data section for details). This model incorporates fixed effects for both location () and time (), removing biases related to unobserved fixed regional heterogeneity and factors that vary over time but impact all regions equally.
Additionally to Eq. (3) focusing on the trend of regional emissions, we also consider the impact of low-carbon investments on the cyclical components of emissions using the following specification:
Note that the superscript now represents the emissions’ cyclical components, therefore, is the cycle component of the GDP (in log) of region at time .
In our specifications, Eq. (3) and (4), the parameters of interest are , , and , which capture how a region’s development level can influence structural funds’ effect in emissions components. The aim is to explore whether the impact of these expenditures varies according to the economic development level of the region, affecting the cyclical and trend components of emissions.
As a robust check, in our specifications, we replace GDP with the set of GVA variables for all economic sectors. This allows us to better control how regional productive structures affect GHG emissions.
4 Results
4.1 Effect on emissions trend
Table 2 presents the estimation results of Eq (3) using as the dependent variable the trends for GHGEpc or GHGEi. The results related to the expenditures for the different groups of regions are presented with respect to the most developed regions (DER). The table reveals mixed patterns that underline the importance of regional development level in the influence of expenditure.
| Dependent variable: | ||||||
| GHGEpc | GHGEi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.158∗∗ | 0.138∗∗ | 0.012 | 0.159∗∗ | 0.171∗∗∗ | 0.108∗∗ |
| (0.069) | (0.057) | (0.054) | (0.062) | (0.053) | (0.050) | |
| Low Carbon Exp. pc | 0.558∗∗ | 0.817∗∗ | 0.049 | 0.806∗∗∗ | 1.321∗∗∗ | 0.802∗∗ |
| (0.226) | (0.415) | (0.398) | (0.204) | (0.389) | (0.371) | |
| Low Carbon Exp. pcTER | 0.029 | 0.111 | 0.127 | 0.111 | 0.059 | 0.004 |
| (0.132) | (0.116) | (0.110) | (0.119) | (0.108) | (0.103) | |
| Low Carbon Exp. pcTER | 2.156∗∗∗ | 1.177 | 3.539∗∗∗ | 2.539∗∗∗ | 1.974∗∗ | 2.895∗∗∗ |
| (0.444) | (0.818) | (0.784) | (0.400) | (0.766) | (0.732) | |
| Low Carbon Exp. pcLDR | 0.730∗∗∗ | 0.620∗∗∗ | 0.489∗∗∗ | 0.361∗∗∗ | 0.356∗∗∗ | 0.315∗∗∗ |
| (0.085) | (0.071) | (0.069) | (0.077) | (0.067) | (0.064) | |
| Low Carbon Exp. pcLDR | 2.470∗∗∗ | 3.329∗∗∗ | 2.468∗∗∗ | 1.382∗∗∗ | 2.154∗∗∗ | 1.666∗∗∗ |
| (0.226) | (0.427) | (0.408) | (0.204) | (0.399) | (0.381) | |
| GDP | 0.209∗∗∗ | 0.297∗∗∗ | 0.631∗∗∗ | 0.525∗∗∗ | ||
| (0.017) | (0.019) | (0.015) | (0.018) | |||
| EPS | 0.009∗∗ | 0.011∗∗ | 0.005 | 0.010∗∗ | ||
| (0.005) | (0.004) | (0.004) | (0.004) | |||
| Political Ideology | 0.016∗∗∗ | 0.001 | 0.002 | 0.007∗∗ | ||
| (0.004) | (0.004) | (0.003) | (0.003) | |||
| GVAA | 0.044∗∗∗ | 0.074∗∗∗ | ||||
| (0.013) | (0.012) | |||||
| GVAB-E | 0.210∗∗∗ | 0.089∗∗∗ | ||||
| (0.015) | (0.014) | |||||
| GVAF | 0.016 | 0.095∗∗∗ | ||||
| (0.014) | (0.013) | |||||
| GVAG-J | 0.0001 | 0.111∗∗∗ | ||||
| (0.029) | (0.027) | |||||
| GVAK-N | 0.327∗∗∗ | 0.506∗∗∗ | ||||
| (0.027) | (0.025) | |||||
| GVAO-U | 0.324∗∗∗ | 0.330∗∗∗ | ||||
| (0.032) | (0.030) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.534 | 0.626 | 0.664 | 0.791 | 0.798 | 0.820 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. | ||||||
Therefore, for developed regions, the reference group, the effect of expenditures in both periods indicates a negative and significant impact in all specifications, except in model 3, where the effect is not significant for both periods. Similarly, the effect of the interaction term for transition regions is negative in all specifications, except in model 2, being non-significant for expenditures associated with the period 2007-2013 and significant for expenditures in the period 2014-2020. This implies that the net impact for transition regions is negative, meaning that, for both transition and developed regions, the long-term effect of expenditures is expected to result in a decrease in emissions, both per capita and in relation to GDP.
In contrast, a significant positive effect is observed for less developed regions, reversing the trend seen in the other development groups. This is because the estimates of the expenditures interaction terms in developing regions are not only positive and significant in all specifications but also cancel any negative effect estimated for the reference group. While explaining the mechanism behind this perverse outcome is far from the scope of this paper, this could indicate that, in less developed regions, expenditures have effectively driven economic development, albeit possibly at the cost of an increase in emissions in the long term. It is important to note that as these regions are developing, the increase in emissions resulting from factors like industrialisation, which is often carbon-intensive, cannot be ignored.
Regarding the control variables, GDP is positively related to GHGEpc and GHGEi, implying that economic growth can lead to an increase in per capita emissions in the long term and to a more efficient economy in terms of emissions relative to GDP.
The EPS dummy has a small but positive impact on per capita emissions and negative emission intensity. Although this positive impact is unexpected, it may also indicate that the effect of more stringent policies may be observed thereafter. Indeed, Naqvi (2021) analysed the cumulative effects of EPS dummy on different emission components, showing that while CH4 and N2O emissions responded strongly to policies and declined continuously across regions, CO2 emissions declined two years after implementing the policies. In the Appendix, we show that the EPS dummy has a negative impact on CH4 and N2O while positive for CO2, the latter being the main component of the total GHG emissions used in this paper (Tables A4 to A6).
The effect of the average political ideology is not robust; while it is found positive for GHGEpc, except in model 3, for GHGEi, it is positive and not significant in model 5 and negative and significant in model 6. These mixed results connect to the complex relationship between the effectiveness of sustainable policies and political ideologies. In this respect, for instance, Garmann (2014) finds that in 19 OECD countries from 1992 to 2008, right-wing governments were associated with fewer emission reductions than the centre and left-wing governments. However, he also observes that emissions were higher in governments with more participating parties, although the distinction between majority and minority governments did not significantly influence emissions.
The effect related to the gross value added by sectors reveals a differentiated impact on GHGEpc and GHGEi. For instance, the GVA of the agriculture sector, GVAA, has a negative impact on both GHGEpc and GHGEi, which suggests that the trend of the agricultural sector production contributes to a reduction of the emissions trends. While the GVA of the manufacturing industry, GVAB-E has a significant positive effect on GHGEpc and a significant negative effect on GHGEi, indicating that greater production in the manufacturing industry implies higher emissions per capita but lower emission intensity.
4.2 Effect in the emissions cycle
Table 3 presents the results of the regressions for the cyclical part of per capita emissions and emissions intensity. The results show mixed effects associated with the expenditures of both periods and the development level interacted variables. These effects are presented in reference to the most developed regions. Therefore, for developed regions, we find that the effect of expenditure in 2007-2013 is negative, although only significant in model 1. The effect of expenditure in 2014-2020 is not significant, but in specifications (2-3 and 5-6), the estimated effect is positive. These results hint that expenditure in the low-carbon economy has no robust effect on the emission cycle for the most developed regions.
| Dependent variable: | ||||||
| GHGEpc | GHGEi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.081∗ | 0.027 | 0.034 | 0.068 | 0.025 | 0.022 |
| (0.045) | (0.045) | (0.045) | (0.045) | (0.045) | (0.046) | |
| Low Carbon Exp. pc | 0.143 | 0.291 | 0.237 | 0.133 | 0.273 | 0.159 |
| (0.147) | (0.331) | (0.330) | (0.146) | (0.331) | (0.335) | |
| Low Carbon Exp. pcTER | 0.120 | 0.279∗∗∗ | 0.323∗∗∗ | 0.113 | 0.282∗∗∗ | 0.326∗∗∗ |
| (0.086) | (0.090) | (0.090) | (0.085) | (0.090) | (0.091) | |
| Low Carbon Exp. pcTER | 0.247 | 1.620∗∗ | 1.807∗∗∗ | 0.266 | 1.746∗∗∗ | 1.558∗∗ |
| (0.285) | (0.638) | (0.637) | (0.283) | (0.637) | (0.645) | |
| Low Carbon Exp. pcLDR | 0.010 | 0.003 | 0.012 | 0.018 | 0.018 | 0.016 |
| (0.056) | (0.057) | (0.057) | (0.055) | (0.057) | (0.058) | |
| Low Carbon Exp. pcLDR | 0.029 | 0.473 | 0.481 | 0.035 | 0.467 | 0.253 |
| (0.148) | (0.341) | (0.341) | (0.147) | (0.341) | (0.345) | |
| GDP | 0.458∗∗∗ | 0.519∗∗∗ | 0.538∗∗∗ | 0.463∗∗∗ | ||
| (0.028) | (0.035) | (0.028) | (0.035) | |||
| EPS | 0.025∗∗∗ | 0.030∗∗∗ | 0.017∗∗∗ | 0.016∗∗∗ | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| Political Ideology | 0.003 | 0.002 | 0.001 | 0.0001 | ||
| (0.003) | (0.003) | (0.003) | (0.003) | |||
| GVAA | 0.011 | 0.006 | ||||
| (0.008) | (0.008) | |||||
| GVAB-E | 0.112∗∗∗ | 0.082∗∗∗ | ||||
| (0.016) | (0.016) | |||||
| GVAF | 0.086∗∗∗ | 0.030∗∗ | ||||
| (0.013) | (0.013) | |||||
| GVAG-J | 0.058∗∗ | 0.199∗∗∗ | ||||
| (0.025) | (0.025) | |||||
| GVAK-N | 0.069∗∗ | 0.143∗∗∗ | ||||
| (0.033) | (0.033) | |||||
| GVAO-U | 0.187∗∗∗ | 0.0003 | ||||
| (0.037) | (0.037) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| R2 | 0.248 | 0.295 | 0.302 | 0.166 | 0.150 | 0.136 |
| Adjusted R2 | 0.193 | 0.235 | 0.241 | 0.105 | 0.078 | 0.060 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
For transition regions, the estimated effect is relatively clearer; nonetheless, determining the net effect related to each program has opposite signs. The interaction with expenditures in the 2007-2013 programme shows a positive and significant effect (except in models 1 and 4) even after discounting the effect of the reference group. The net effect of expenditure in this period suggests an increase in the intensity of emission cycles, both per capita and concerning GDP. In contrast, the effect associated with expenditure in the 2014-2020 programme is consistently negative and significant (except in models 1 and 4). Then, considering the effect of the reference group, the net effect remains negative and comparatively much stronger. Although the estimated effects for both periods are opposing, the expected effect of these expenditures is to reduce the intensity of emission cycles. For less developed regions, the effect of expenditures associated with emission cycles is negative in almost all specifications for both periods, although not significant. Thus, considering the estimated effects in the reference group, the effect of expenditures on emission cycles in less developed regions is not significant.
Regarding the other control variables, the GDP cycle has a strong positive impact on the per capita emission cycle and a negative impact on the emission intensity cycle, indicating that economic growth cycles are associated with higher per capita emissions but with a lower intensity of emissions intensity. The country-level EPS dummy shows a significant negative effect on emissions, both per capita and intensity, suggesting that stricter environmental policies are associated with lower emissions in their cycles. The region’s weighted average political ideology does not show a significant impact on emission cycles, indicating that local political outcomes, whether left or right, may not have a direct relationship with the cyclical variability of emissions.
Finally, the sectoral gross value added coefficients indicate that different sectors have varying impacts on emission cycles. For example, the manufacturing sector (B-E) has a significant positive impact on per capita emissions and a negative impact on emission intensity. Therefore, most industry sectors have a significant influence on the cyclical variations in emissions.
4.3 Effect by type of gases
We break down emissions by specific gases, considering CO2, CH4, and N2O, to assess low-carbon investments’ impacts, using the same methodology. Detailed numerical result tables can be found in the Appendix (Tables A4 to A9), allowing us to concentrate on qualitative insights.
Disaggregated GHG emissions data at the regional level allow for a better understanding of the heterogeneous impacts of the investments and can help identify specific expenditure gaps to improve policy responses. While CO2 contributes significantly to GHG emissions and changes in CO2 emissions primarily explain the impact of low-carbon funding, it is essential to consider a broader range of gases, which have different sources and economic activity patterns. In addition, decomposing GHG emissions into leading pollutant gases is relevant as European regions vary significantly in emission types, which are linked to their socioeconomic composition (Crippa et al., 2023). CO2 results from burning fossil fuels, chemical processes like cement production, and natural emissions from trees and biological materials. CH4 is emitted during fossil fuel extraction and transportation and significantly from agricultural activities such as livestock and organic waste decomposition in landfills. N2O is produced by agricultural and industrial activities, fossil fuel combustion, and wastewater treatment.
The fixed effects estimations results on CO2, CH4, and N2O emissions are presented on Tables A4 to A6 in the Appendix.
For CO2 emissions, low-carbon expenditures for the 2007-2013 and 2014-2020 programmes show significant negative impacts on emissions per capita and relative to GDP in developed and transition regions, indicating effective reductions. However, in less developed regions, these expenditures result in positive significant impacts, suggesting an increase in emissions or less effective reductions. These results are similar to those observed for overall GHG emissions (Table 2).
For CH4 emissions, low-carbon expenditures for the 2007-2013 programme generally show negative but non-significant impacts across most specifications (including emissions per capita and relative to GDP). For the 2014-2020 programme, there are significant negative impacts in transition regions but positive or non-significant impacts in less developed regions, indicating inconsistent effectiveness. Regarding N2O emissions, low-carbon expenditures for both the 2007-2013 and 2014-2020 programmes have negative impacts but are not significant in developed regions. In transition regions, the impacts are negative and significant, while in less developed regions, the impacts are positive and significant, indicating less effective reductions or increases in emissions.
These results show that low-carbon expenditures have varying impacts depending on the type of gas and the level of regional development. In developed and transition regions, these investments are generally effective in reducing CO2 emissions but exhibit mixed and often non-significant results for CH4 and N2O. In less developed regions, low-carbon expenditures frequently result in increased emissions or less effective reductions across all gases.
Tables A7 to A9 in the Appendix present the fixed effect estimation results on the cyclical components of CO2, CH4, and N2O emissions, respectively. For CO2 emissions, low-carbon expenditures for the 2007-2013 and 2014-2020 programmes show generally non-significant impacts on the cyclical component of emissions per capita and relative to GDP. There are mixed results in transition regions, with significant positive impacts for 2007-2013 expenditures and significant negative impacts for 2014-2020 expenditures. In less developed regions, the impacts are primarily non-significant, with some negative impacts on the cyclical emission component for 2014-2020 expenditures in certain specifications.
For CH4 emissions, low-carbon expenditures for the 2007-2013 programme show consistently significant negative impacts for all regions, indicating effective reductions on the cyclical emissions component. The 2014-2020 expenditures yield mixed results, with generally non-significant impacts except for significant negative impacts in transition regions. Regarding N2O emissions, low-carbon expenditures for both the 2007-2013 and 2014-2020 programmes typically show non-significant impacts. In transition regions, there are some negative differential impacts for 2014-2020 expenditures, but these are not consistently significant. In less developed regions, the differential impacts are mixed, with positive and significant for the 2007-2013 period and non-significant for 2014-2020.
The estimations for the cyclical emissions components reveal that low-carbon expenditures have mixed results, which vary depending on the type of gas used and the level of regional development. This is observed for both per capita and per unit of GDP.
5 Concluding remarks
This paper examines the impact of climate action through one of the key vehicles for European regions: the “Low-Carbon Economy” funds of the EU Cohesion and Structural Policy for the 2007-2013 and 2014-2020 programmes. Our analysis reveals that the impact of spending on low-carbon economies in reducing GHG emissions varies significantly depending on the level of regional development, according to the EU classification that distinguishes between less developed regions, regions in transition, and developed regions. We observe that, for both analysed programmes, investments in low-carbon economies are associated with a statistically significant decrease in the long-term trend of total emissions, both per capita and relative to GDP, in developed and transition regions. However, in less developed regions, these investments have a statistically significant positive impact, indicating either an increase in emissions or a less effective reduction. Furthermore, estimations based on specific gas types also demonstrate differential impacts.
These results indicate a clear dichotomy: investments yield significant reductions in long-term emissions trends in developed and transition regions but appear to have the opposite effect in less developed regions. This aligns with insights from Doda (2014) and Sheldon (2017), who emphasise the variability of emission responses to economic cycles, particularly noting the stabilising effect of wealthier economies on emission patterns. Conversely, in less developed regions, the counterintuitive increase in emissions suggests a subtle interplay between economic development and environmental sustainability efforts, reminiscent of the disparities identified in the existing literature (Naqvi and Zwickl, 2017; Ivanova et al., 2017).
In addition, our analysis showed that European regions’ responses to low-carbon economies funds vary significantly among programmes, emphasising the importance of taking a nuanced approach when evaluating the impact of these investments on environmental sustainability. This involves considering the particularities of each program, the intrinsic differences between regions and types of emissions.
Our findings support the idea of a refined approach to the implementation of climate policy objectives through the EU cohesion policy. We agree with other studies claiming that a diverse set of guidelines must be implemented to effectively reduce emissions rather than relying on individual initiatives (Brinkley, 2014). Consequently, economy-wide assessments, sector-specific trends, planning expertise in implementing place-based approaches, and extensive case studies can help underscore the complexity of decoupling economic growth from pollutant emissions (Brinkley, 2014; Naqvi and Zwickl, 2017).
In general, the need for differentiated policies to achieve economic development is an issue that has been widely studied since the creation of the Cohesion Policy for European regions, underscoring the importance of addressing persistent inequalities between EU regions. Very early on, Boldrin and Canova (2001) highlighted that these policies tend to be more redistributive than growth-oriented, reflecting political compromises over economic efficiency. However, investments in education and human capital, which constitute a significant portion of these funds, have shown medium-term positive impacts, emphasising the critical role of human capital development in fostering long-term regional growth and convergence (Rodríguez-Pose and Fratesi, 2004). Moreover, Rodríguez-Pose and Garcilazo (2015) and Charron et al. (2014) argue that the quality of local and regional governance directly affects economic growth and the effectiveness of structural funds. They find significant variations in governance quality within countries, such as levels of corruption and efficiency of public services, which have substantial implications for policy design. Iammarino et al. (2019) emphasise the need for tailored policies that are sensitive to the structural prospects of different regions, aligning with the Barca report’s advocacy for place-based strategies (Barca, 2009). These insights align with the broader literature on the Environmental Kuznets Hypothesis, which suggests that economic development can lead to improved environmental outcomes, but only if accompanied by effective governance and tailored policy interventions.
Our research underscores the pressing need for tailored environmental policies to optimise interventions across European regions. The differential impacts of cohesion policy investments on regional emission trends and cycles highlight the necessity of moving away from one-size-fits-all solutions towards development-sensitive and emission-specific strategies. Special attention should be given to less developed regions, which, although not major emitters per capita currently, risk becoming significant pollution producers as they develop economically. Aligning investment strategies with regional profiles can enhance the effectiveness of emissions reduction efforts, contributing to a more nuanced policy dialogue on sustainable development. This underscores the critical intersection of regional growth, sustainability investments, and environmental outcomes within the EU’s policy framework.
Acknowledgements
The authors acknowledge the financial support of the DECIPHER (Decision-making framework and processes for holistic evaluation of environmental and climate policies) project, funded by the Horizon Europe: Research and Innovation programme under grant agreement No. 101056898.
Appendix
Low-carbon economy definitions by ESI programmes
| 2014-2020 funding program | |
| Intervention field code | Intervention field description |
| 009 | Renewable energy: wind |
| 010 | Renewable energy: solar |
| 011 | Renewable energy: biomass |
| 012 | Other renewable energy (including hydroelectric, geothermal and marine energy) and renewable energy integration (including storage, power to gas and renewable hydrogen infrastructure) |
| 013 | Energy efficiency renovation of public infrastructure, demonstration projects and supporting measures |
| 014 | Energy efficiency renovation of existing housing stock, demonstration projects and supporting measures |
| 015 | Intelligent Energy Distribution Systems at medium and low voltage levels (including smart grids and ICT systems) |
| 016 | High efficiency co-generation and district heating |
| 023 | Environmental measures aimed at reducing and / or avoiding greenhouse gas emissions (including treatment and storage of methane gas and composting) |
| 043 | Clean urban transport infrastructure and promotion (including equipment and rolling stock) |
| 068 | Energy efficiency and demonstration projects in SMEs and supporting measures |
| 070 | Promotion of energy efficiency in large enterprises |
| 090 | Cycle tracks and footpaths |
| 2007-2013 funding program | |
| Priority code | Priority description |
| 06 | Assistance to small and medium enterprises (SMEs) for the promotion of environmentally-friendly products and production processes (…) |
| 24 | Cycle tracks |
| 28 | Intelligent transport systems |
| 39 | Renewable energy: wind |
| 40 | Renewable energy: solar |
| 41 | Renewable energy: biomass |
| 42 | Renewable energy: hydroelectric, geothermal and other |
| 43 | Energy efficiency, co-generation, energy management |
| 25-52 | Promotion of clean urban transport |
Estimation results modelled expenditures
| Dependent variable: | ||||||
| GHGEpc | GHGEi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.182∗ | 0.167∗ | 0.004 | 0.185∗∗ | 0.213∗∗∗ | 0.134∗ |
| (0.105) | (0.085) | (0.081) | (0.093) | (0.079) | (0.075) | |
| Low Carbon Exp. pc | 1.239∗ | 5.703 | 2.736 | 1.786∗∗∗ | 9.205∗∗ | 3.117 |
| (0.715) | (3.972) | (3.798) | (0.635) | (3.684) | (3.512) | |
| Low Carbon Exp. pcTER | 0.051 | 0.151 | 0.178 | 0.198 | 0.107 | 0.046 |
| (0.200) | (0.169) | (0.161) | (0.177) | (0.157) | (0.149) | |
| Low Carbon Exp. pcTER | 5.945∗∗∗ | 13.126 | 34.828∗∗∗ | 7.612∗∗∗ | 23.200∗∗ | 29.144∗∗∗ |
| (1.559) | (12.028) | (11.466) | (1.383) | (11.155) | (10.604) | |
| Low Carbon Exp. pcLDR | 0.792∗∗∗ | 0.768∗∗∗ | 0.588∗∗∗ | 0.394∗∗∗ | 0.439∗∗∗ | 0.387∗∗∗ |
| (0.129) | (0.108) | (0.103) | (0.114) | (0.100) | (0.095) | |
| Low Carbon Exp. pcLDR | 6.004∗∗∗ | 36.630∗∗∗ | 29.030∗∗∗ | 3.151∗∗∗ | 18.037∗∗∗ | 13.218∗∗∗ |
| (0.736) | (5.048) | (4.862) | (0.653) | (4.682) | (4.497) | |
| GDP | 0.246∗∗∗ | 0.321∗∗∗ | 0.615∗∗∗ | 0.510∗∗∗ | ||
| (0.017) | (0.019) | (0.015) | (0.017) | |||
| EPS | 0.010∗∗ | 0.013∗∗∗ | 0.005 | 0.010∗∗ | ||
| (0.005) | (0.005) | (0.004) | (0.004) | |||
| Political Ideology | 0.017∗∗∗ | 0.001 | 0.002 | 0.007∗∗ | ||
| (0.004) | (0.004) | (0.003) | (0.003) | |||
| GVAA | 0.045∗∗∗ | 0.072∗∗∗ | ||||
| (0.013) | (0.012) | |||||
| GVAB-E | 0.222∗∗∗ | 0.084∗∗∗ | ||||
| (0.015) | (0.014) | |||||
| GVAF | 0.023 | 0.095∗∗∗ | ||||
| (0.014) | (0.013) | |||||
| GVAG-J | 0.0004 | 0.110∗∗∗ | ||||
| (0.030) | (0.027) | |||||
| GVAK-N | 0.311∗∗∗ | 0.498∗∗∗ | ||||
| (0.027) | (0.025) | |||||
| GVAO-U | 0.345∗∗∗ | 0.334∗∗∗ | ||||
| (0.033) | (0.030) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.510 | 0.614 | 0.653 | 0.787 | 0.796 | 0.817 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
| Dependent variable: | ||||||
| GHGEpc | GHGEi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.049 | 0.020 | 0.013 | 0.031 | 0.024 | 0.030 |
| (0.067) | (0.067) | (0.067) | (0.067) | (0.067) | (0.068) | |
| Low Carbon Exp. pc | 0.378 | 6.016∗ | 5.025 | 0.306 | 5.767∗ | 4.607 |
| (0.456) | (3.122) | (3.120) | (0.453) | (3.120) | (3.160) | |
| Low Carbon Exp. pcTER | 0.163 | 0.376∗∗∗ | 0.426∗∗∗ | 0.162 | 0.385∗∗∗ | 0.453∗∗∗ |
| (0.127) | (0.131) | (0.131) | (0.126) | (0.130) | (0.132) | |
| Low Carbon Exp. pcTER | 0.076 | 25.584∗∗∗ | 30.159∗∗∗ | 0.072 | 25.857∗∗∗ | 23.884∗∗ |
| (0.984) | (9.360) | (9.338) | (0.978) | (9.354) | (9.457) | |
| Low Carbon Exp. pcLDR | 0.028 | 0.001 | 0.009 | 0.044 | 0.025 | 0.013 |
| (0.082) | (0.085) | (0.085) | (0.082) | (0.085) | (0.086) | |
| Low Carbon Exp. pcLDR | 0.225 | 5.407 | 5.581 | 0.230 | 5.182 | 2.318 |
| (0.471) | (3.963) | (3.987) | (0.468) | (3.961) | (4.038) | |
| GDP | 0.460∗∗∗ | 0.519∗∗∗ | 0.535∗∗∗ | 0.464∗∗∗ | ||
| (0.028) | (0.035) | (0.028) | (0.035) | |||
| EPS | 0.026∗∗∗ | 0.031∗∗∗ | 0.018∗∗∗ | 0.018∗∗∗ | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| Political Ideology | 0.003 | 0.003 | 0.0004 | 0.0004 | ||
| (0.003) | (0.003) | (0.003) | (0.003) | |||
| GVAA | 0.010 | 0.007 | ||||
| (0.008) | (0.008) | |||||
| GVAB-E | 0.115∗∗∗ | 0.078∗∗∗ | ||||
| (0.016) | (0.016) | |||||
| GVAF | 0.086∗∗∗ | 0.030∗∗ | ||||
| (0.012) | (0.013) | |||||
| GVAG-J | 0.054∗∗ | 0.203∗∗∗ | ||||
| (0.025) | (0.025) | |||||
| GVAK-N | 0.076∗∗ | 0.134∗∗∗ | ||||
| (0.032) | (0.033) | |||||
| GVAO-U | 0.176∗∗∗ | 0.013 | ||||
| (0.037) | (0.037) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.191 | 0.235 | 0.241 | 0.103 | 0.078 | 0.061 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
Regression by type of polluting gas: trend components
Trend components
| Dependent variable: | ||||||
| CO2pc | CO2i | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.170∗∗ | 0.161∗∗ | 0.002 | 0.172∗∗ | 0.193∗∗∗ | 0.094 |
| (0.078) | (0.066) | (0.062) | (0.073) | (0.062) | (0.058) | |
| Low Carbon Exp. pc | 0.678∗∗∗ | 0.833∗ | 0.234 | 0.926∗∗∗ | 1.337∗∗∗ | 0.617 |
| (0.255) | (0.481) | (0.453) | (0.239) | (0.456) | (0.427) | |
| Low Carbon Exp. pcTER | 0.003 | 0.268∗∗ | 0.013 | 0.084 | 0.097 | 0.110 |
| (0.149) | (0.134) | (0.126) | (0.140) | (0.127) | (0.118) | |
| Low Carbon Exp. pcTER | 1.718∗∗∗ | 0.049 | 2.866∗∗∗ | 2.102∗∗∗ | 0.847 | 2.222∗∗∗ |
| (0.501) | (0.947) | (0.894) | (0.470) | (0.899) | (0.842) | |
| Low Carbon Exp. pcLDR | 0.739∗∗∗ | 0.647∗∗∗ | 0.490∗∗∗ | 0.370∗∗∗ | 0.383∗∗∗ | 0.315∗∗∗ |
| (0.096) | (0.083) | (0.078) | (0.090) | (0.078) | (0.074) | |
| Low Carbon Exp. pcLDR | 2.603∗∗∗ | 3.441∗∗∗ | 2.423∗∗∗ | 1.514∗∗∗ | 2.267∗∗∗ | 1.621∗∗∗ |
| (0.255) | (0.494) | (0.466) | (0.239) | (0.469) | (0.439) | |
| GDP | 0.225∗∗∗ | 0.375∗∗∗ | 0.615∗∗∗ | 0.446∗∗∗ | ||
| (0.019) | (0.022) | (0.018) | (0.021) | |||
| EPS | 0.008 | 0.010∗∗ | 0.004 | 0.009∗ | ||
| (0.005) | (0.005) | (0.005) | (0.005) | |||
| Political Ideology | 0.022∗∗∗ | 0.002 | 0.008∗ | 0.006 | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| GVAA | 0.065∗∗∗ | 0.095∗∗∗ | ||||
| (0.014) | (0.014) | |||||
| GVAB-E | 0.244∗∗∗ | 0.055∗∗∗ | ||||
| (0.017) | (0.016) | |||||
| GVAF | 0.005 | 0.074∗∗∗ | ||||
| (0.016) | (0.015) | |||||
| GVAG-J | 0.032 | 0.143∗∗∗ | ||||
| (0.033) | (0.031) | |||||
| GVAK-N | 0.406∗∗∗ | 0.585∗∗∗ | ||||
| (0.031) | (0.029) | |||||
| GVAO-U | 0.430∗∗∗ | 0.435∗∗∗ | ||||
| (0.037) | (0.035) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.531 | 0.609 | 0.659 | 0.760 | 0.763 | 0.796 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
| Dependent variable: | ||||||
| CH4pc | CH4i | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.134∗ | 0.100 | 0.103 | 0.136∗ | 0.132∗∗ | 0.199∗∗∗ |
| (0.079) | (0.064) | (0.064) | (0.075) | (0.065) | (0.066) | |
| Low Carbon Exp. pc | 0.378 | 0.077 | 0.114 | 0.130 | 0.580 | 0.965∗∗ |
| (0.258) | (0.471) | (0.468) | (0.247) | (0.473) | (0.483) | |
| Low Carbon Exp. pcTER | 0.037 | 0.033 | 0.044 | 0.044 | 0.138 | 0.080 |
| (0.151) | (0.131) | (0.130) | (0.145) | (0.132) | (0.134) | |
| Low Carbon Exp. pcTER | 2.751∗∗∗ | 2.460∗∗∗ | 3.132∗∗∗ | 3.133∗∗∗ | 3.258∗∗∗ | 2.488∗∗∗ |
| (0.507) | (0.928) | (0.923) | (0.487) | (0.933) | (0.952) | |
| Low Carbon Exp. pcLDR | 0.406∗∗∗ | 0.429∗∗∗ | 0.352∗∗∗ | 0.038 | 0.165∗∗ | 0.177∗∗ |
| (0.097) | (0.081) | (0.081) | (0.093) | (0.081) | (0.083) | |
| Low Carbon Exp. pcLDR | 0.404 | 1.339∗∗∗ | 0.836∗ | 0.683∗∗∗ | 0.163 | 0.034 |
| (0.258) | (0.484) | (0.481) | (0.248) | (0.486) | (0.496) | |
| GDP | 0.172∗∗∗ | 0.100∗∗∗ | 0.668∗∗∗ | 0.722∗∗∗ | ||
| (0.020) | (0.021) | (0.019) | (0.021) | |||
| EPS | 0.026∗∗∗ | 0.019∗∗∗ | 0.030∗∗∗ | 0.021∗∗∗ | ||
| (0.005) | (0.005) | (0.005) | (0.005) | |||
| Political Ideology | 0.003 | 0.003 | 0.017∗∗∗ | 0.011∗∗ | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| GVAA | 0.023 | 0.007 | ||||
| (0.015) | (0.015) | |||||
| GVAB-E | 0.108∗∗∗ | 0.191∗∗∗ | ||||
| (0.018) | (0.018) | |||||
| GVAF | 0.136∗∗∗ | 0.214∗∗∗ | ||||
| (0.016) | (0.017) | |||||
| GVAG-J | 0.112∗∗∗ | 0.001 | ||||
| (0.034) | (0.036) | |||||
| GVAK-N | 0.010 | 0.189∗∗∗ | ||||
| (0.032) | (0.033) | |||||
| GVAO-U | 0.170∗∗∗ | 0.175∗∗∗ | ||||
| (0.038) | (0.039) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.345 | 0.473 | 0.490 | 0.683 | 0.716 | 0.711 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
| Dependent variable: | ||||||
| N2Opc | N2Oi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.108 | 0.060 | 0.013 | 0.102∗ | 0.082 | 0.111∗∗ |
| (0.071) | (0.055) | (0.054) | (0.061) | (0.053) | (0.055) | |
| Low Carbon Exp. pc | 0.189 | 0.052 | 0.553 | 0.366∗ | 0.452 | 0.291 |
| (0.233) | (0.403) | (0.397) | (0.199) | (0.386) | (0.402) | |
| Low Carbon Exp. pcTER | 0.254∗ | 0.036 | 0.104 | 0.353∗∗∗ | 0.240∗∗ | 0.021 |
| (0.136) | (0.113) | (0.110) | (0.116) | (0.108) | (0.111) | |
| Low Carbon Exp. pcTER | 1.804∗∗∗ | 1.333∗ | 2.032∗∗∗ | 2.275∗∗∗ | 2.352∗∗∗ | 1.364∗ |
| (0.457) | (0.795) | (0.781) | (0.391) | (0.760) | (0.792) | |
| Low Carbon Exp. pcLDR | 0.693∗∗∗ | 0.344∗∗∗ | 0.290∗∗∗ | 0.331∗∗∗ | 0.085 | 0.116∗ |
| (0.087) | (0.069) | (0.068) | (0.075) | (0.066) | (0.069) | |
| Low Carbon Exp. pcLDR | 2.399∗∗∗ | 1.722∗∗∗ | 1.273∗∗∗ | 1.297∗∗∗ | 0.534 | 0.465 |
| (0.233) | (0.414) | (0.407) | (0.199) | (0.396) | (0.412) | |
| GDP | 0.103∗∗∗ | 0.135∗∗∗ | 0.759∗∗∗ | 0.714∗∗∗ | ||
| (0.018) | (0.019) | (0.015) | (0.018) | |||
| EPS | 0.006 | 0.010∗∗ | 0.010∗∗ | 0.012∗∗ | ||
| (0.004) | (0.004) | (0.004) | (0.005) | |||
| Political Ideology | 0.016∗∗∗ | 0.010∗∗∗ | 0.003 | 0.001 | ||
| (0.004) | (0.004) | (0.003) | (0.004) | |||
| GVAA | 0.091∗∗∗ | 0.062∗∗∗ | ||||
| (0.013) | (0.013) | |||||
| GVAB-E | 0.113∗∗∗ | 0.186∗∗∗ | ||||
| (0.015) | (0.015) | |||||
| GVAF | 0.040∗∗∗ | 0.039∗∗∗ | ||||
| (0.014) | (0.014) | |||||
| GVAG-J | 0.009 | 0.101∗∗∗ | ||||
| (0.029) | (0.030) | |||||
| GVAK-N | 0.188∗∗∗ | 0.369∗∗∗ | ||||
| (0.027) | (0.027) | |||||
| GVAO-U | 0.032 | 0.042 | ||||
| (0.036) | (0.036) | |||||
| Observations | 3,856 | 2,982 | 2,982 | 3,856 | 2,982 | 2,982 |
| Adjusted R2 | 0.109 | 0.159 | 0.203 | 0.594 | 0.619 | 0.595 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
Cycle components
| Dependent variable: | ||||||
| CO2pc | CO2i | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.072 | 0.003 | 0.008 | 0.059 | 0.001 | 0.004 |
| (0.054) | (0.054) | (0.054) | (0.054) | (0.054) | (0.054) | |
| Low Carbon Exp. pc | 0.162 | 0.482 | 0.394 | 0.152 | 0.463 | 0.315 |
| (0.177) | (0.395) | (0.394) | (0.177) | (0.396) | (0.397) | |
| Low Carbon Exp. pcTER | 0.096 | 0.308∗∗∗ | 0.374∗∗∗ | 0.089 | 0.311∗∗∗ | 0.377∗∗∗ |
| (0.104) | (0.108) | (0.108) | (0.103) | (0.108) | (0.108) | |
| Low Carbon Exp. pcTER | 0.115 | 1.745∗∗ | 1.918∗∗ | 0.134 | 1.871∗∗ | 1.668∗∗ |
| (0.344) | (0.763) | (0.760) | (0.342) | (0.763) | (0.765) | |
| Low Carbon Exp. pcLDR | 0.062 | 0.052 | 0.063 | 0.069 | 0.067 | 0.067 |
| (0.067) | (0.068) | (0.068) | (0.067) | (0.068) | (0.068) | |
| Low Carbon Exp. pcLDR | 0.024 | 0.681∗ | 0.721∗ | 0.030 | 0.674∗ | 0.494 |
| (0.179) | (0.407) | (0.407) | (0.178) | (0.408) | (0.410) | |
| GDP | 0.567∗∗∗ | 0.660∗∗∗ | 0.428∗∗∗ | 0.323∗∗∗ | ||
| (0.034) | (0.042) | (0.034) | (0.042) | |||
| EPS | 0.036∗∗∗ | 0.041∗∗∗ | 0.028∗∗∗ | 0.028∗∗∗ | ||
| (0.004) | (0.005) | (0.004) | (0.005) | |||
| Political Ideology | 0.001 | 0.0004 | 0.003 | 0.002 | ||
| (0.003) | (0.003) | (0.003) | (0.003) | |||
| GVAA | 0.006 | 0.010 | ||||
| (0.010) | (0.010) | |||||
| GVAB-E | 0.125∗∗∗ | 0.069∗∗∗ | ||||
| (0.019) | (0.019) | |||||
| GVAF | 0.100∗∗∗ | 0.044∗∗∗ | ||||
| (0.015) | (0.015) | |||||
| GVAG-J | 0.078∗∗∗ | 0.179∗∗∗ | ||||
| (0.030) | (0.030) | |||||
| GVAK-N | 0.125∗∗∗ | 0.087∗∗ | ||||
| (0.039) | (0.039) | |||||
| GVAO-U | 0.291∗∗∗ | 0.103∗∗ | ||||
| (0.044) | (0.044) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.206 | 0.261 | 0.270 | 0.044 | 0.042 | 0.041 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
| Dependent variable: | ||||||
| CH4pc | CH4i | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.198∗∗∗ | 0.174∗∗∗ | 0.184∗∗∗ | 0.185∗∗∗ | 0.173∗∗∗ | 0.172∗∗∗ |
| (0.047) | (0.048) | (0.048) | (0.047) | (0.047) | (0.048) | |
| Low Carbon Exp. pc | 0.075 | 0.188 | 0.168 | 0.085 | 0.207 | 0.247 |
| (0.152) | (0.348) | (0.348) | (0.152) | (0.346) | (0.353) | |
| Low Carbon Exp. pcTER | 0.152∗ | 0.237∗∗ | 0.256∗∗∗ | 0.146 | 0.240∗∗ | 0.259∗∗∗ |
| (0.089) | (0.095) | (0.095) | (0.089) | (0.094) | (0.096) | |
| Low Carbon Exp. pcTER | 0.699∗∗ | 1.353∗∗ | 1.455∗∗ | 0.718∗∗ | 1.479∗∗ | 1.206∗ |
| (0.295) | (0.671) | (0.670) | (0.294) | (0.667) | (0.681) | |
| Low Carbon Exp. pcLDR | 0.094 | 0.066 | 0.069 | 0.086 | 0.052 | 0.066 |
| (0.058) | (0.060) | (0.060) | (0.057) | (0.060) | (0.061) | |
| Low Carbon Exp. pcLDR | 0.199 | 0.198 | 0.124 | 0.194 | 0.190 | 0.105 |
| (0.153) | (0.359) | (0.359) | (0.153) | (0.356) | (0.364) | |
| GDP | 0.163∗∗∗ | 0.181∗∗∗ | 0.831∗∗∗ | 0.802∗∗∗ | ||
| (0.029) | (0.037) | (0.029) | (0.037) | |||
| EPS | 0.004 | 0.008∗∗ | 0.004 | 0.006 | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| Political Ideology | 0.002 | 0.001 | 0.002 | 0.001 | ||
| (0.003) | (0.003) | (0.003) | (0.003) | |||
| GVAA | 0.009 | 0.007 | ||||
| (0.009) | (0.009) | |||||
| GVAB-E | 0.048∗∗∗ | 0.146∗∗∗ | ||||
| (0.016) | (0.017) | |||||
| GVAF | 0.074∗∗∗ | 0.018 | ||||
| (0.013) | (0.013) | |||||
| GVAG-J | 0.002 | 0.260∗∗∗ | ||||
| (0.026) | (0.027) | |||||
| GVAK-N | 0.011 | 0.223∗∗∗ | ||||
| (0.034) | (0.035) | |||||
| GVAO-U | 0.045 | 0.233∗∗∗ | ||||
| (0.039) | (0.039) | |||||
| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
| Adjusted R2 | 0.034 | 0.040 | 0.032 | 0.258 | 0.243 | 0.216 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
| Dependent variable: | ||||||
| N2Opc | N2Oi | |||||
| (1) | (2) | (3) | (4) | (5) | (6) | |
| Low Carbon Exp. pc | 0.059 | 0.035 | 0.062 | 0.045 | 0.032 | 0.049 |
| (0.056) | (0.063) | (0.063) | (0.055) | (0.063) | (0.064) | |
| Low Carbon Exp. pc | 0.050 | 0.160 | 0.157 | 0.062 | 0.145 | 0.076 |
| (0.181) | (0.462) | (0.461) | (0.180) | (0.458) | (0.465) | |
| Low Carbon Exp. pcTER | 0.072 | 0.076 | 0.101 | 0.080 | 0.074 | 0.098 |
| (0.106) | (0.126) | (0.126) | (0.105) | (0.125) | (0.127) | |
| Low Carbon Exp. pcTER | 0.245 | 0.718 | 1.139 | 0.263 | 0.844 | 0.893 |
| (0.351) | (0.891) | (0.889) | (0.348) | (0.883) | (0.896) | |
| Low Carbon Exp. pcLDR | 0.191∗∗∗ | 0.214∗∗∗ | 0.193∗∗ | 0.184∗∗∗ | 0.200∗∗ | 0.190∗∗ |
| (0.069) | (0.080) | (0.080) | (0.068) | (0.079) | (0.080) | |
| Low Carbon Exp. pcLDR | 0.065 | 0.050 | 0.016 | 0.070 | 0.057 | 0.250 |
| (0.183) | (0.476) | (0.476) | (0.181) | (0.472) | (0.480) | |
| GDP | 0.140∗∗∗ | 0.157∗∗∗ | 0.856∗∗∗ | 0.826∗∗∗ | ||
| (0.035) | (0.049) | (0.035) | (0.049) | |||
| EPS | 0.0002 | 0.004 | 0.007 | 0.017∗∗∗ | ||
| (0.005) | (0.005) | (0.005) | (0.005) | |||
| Political Ideology | 0.009∗∗ | 0.007∗ | 0.005 | 0.005 | ||
| (0.004) | (0.004) | (0.004) | (0.004) | |||
| GVAA | 0.029∗∗ | 0.012 | ||||
| (0.012) | (0.012) | |||||
| GVAB-E | 0.039∗ | 0.156∗∗∗ | ||||
| (0.022) | (0.022) | |||||
| GVAF | 0.019 | 0.038∗∗ | ||||
| (0.017) | (0.018) | |||||
| GVAG-J | 0.114∗∗∗ | 0.147∗∗∗ | ||||
| (0.035) | (0.035) | |||||
| GVAK-N | 0.166∗∗∗ | 0.383∗∗∗ | ||||
| (0.046) | (0.046) | |||||
| GVAO-U | 0.031 | 0.154∗∗∗ | ||||
| (0.052) | (0.052) | |||||
| Observations | 3,856 | 2,982 | 2,982 | 3,856 | 2,982 | 2,982 |
| Adjusted R2 | 0.018 | 0.011 | 0.020 | 0.241 | 0.175 | 0.156 |
| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||

### v2

@@
 Abstract
-This paper investigates the effectiveness of the “low-carbon economy” expenditures from European Structural and Investment Funds in fostering reductions in greenhouse gas emissions within European regions, focusing on the 2007-2013 and 2014-2020 programme periods.
-By decomposing emissions time series into trend and cycle components and considering them within a panel data framework, our research highlights that the impacts of low-carbon economy expenditures vary, qualitatively and quantitatively, with the targeted regions’ development level.
-We find significant emissions reductions in developed and transition regions yet less favourable outcomes in less developed areas.
-Further analysis into specific greenhouse gas emissions types (CO2, CH4, and N2O) reveals inconsistent impacts, underscoring the complexity of achieving emissions reductions.
-Our findings emphasise the need for tailored environmental strategies that accommodate the economic disparities of regions in the European Union.
-Keywords: Greenhouse gas emissions; Emission decomposition; Low-carbon economy; European Structural and Investment Funds; Regional development
+We investigate the effectiveness of low-carbon expenditures from the European Structural and Investment Funds in reducing greenhouse gas emissions across EU regions.
+Using trend and cycle decomposition of two indicators – per capita emissions and emissions intensity – we find highly heterogeneous effects.
+In less developed regions, investments are linked to long-term increases in per capita emissions, while in transition and developed regions, effects are weak or not significant.
+Instrumental variable estimates confirm this pattern.
+When disaggregated by gas type, results remain inconsistent.
+Overall, our findings highlight the limited climate effectiveness of EU funding and the need for region-specific strategies.
+Keywords: Greenhouse gas emissions; Green transition; Climate action; Low-carbon economy; European Structural and Investment Funds; Regional development
 1 Introduction
-The European Union (EU) has the objective of achieving climate-neutrality by 2050.
+The European Union (EU) has the objective of achieving climate neutrality by 2050.
 This implies large-scale investments in the energy transition (McCollum et al., 2018; Klaassen and Steffen, 2023).
@@
 Accordingly, the EU has enacted climate spending targets amounting to 20 of its 2014-2020 budget and 30 of its 2021-2027 budget that ought to be devoted to climate action111See European Court of Auditors (2022), which also emphasises that climate spending over the 2014-2020 period has been over-estimated..
-In this context, one can ask a necessary question about the efficiency of European climate policy: have these expenses effectively led to the reduction of greenhouse gas (GHG) emissions?
+Therefore, it is reasonable to question the efficiency of European climate policy to determine if these expenses effectively lead to reducing greenhouse gas (GHG) emissions.
 This paper exploits regional variations in European climate spending to address this issue.
@@
 Although the programming period 2007-2013 did not define thematic objectives in the same way, the aims of the expenditures are also tagged by “priority codes”, from which it is possible to determine the funds related to climate action consistently with the low-carbon economy thematic objective from 2014-2020.
-Consequently, we use the total amount spent on the low-carbon economy objective in the last two budgeting periods (2007-2013 and 2014-2020) as a measure of public climate investment.
+Consequently, we use the total amount spent on the low-carbon economy objective for the 2007-2013 and 2014-2020 budgeting periods as a measure of public climate investment.
 We then employ an econometric panel data model to investigate the influence of these investments on GHG emissions in European regions between 2007 and 2022 at the NUTS 2 level.
-Our methodology involves decomposing the emission time series of each region into trend and cycle components.
-This strategy gives us deeper insights into the nexus between economic growth and environmental sustainability.
-Due to the heterogeneity of European regions, it is clear that some have significantly advanced in terms of sustainability, achieving various degrees of decoupling, whereas many others have yet to do so.
-This disparity must be accounted for in the analysis to determine such expenditures’ short-term and long-term effects.
-Our analysis also accounts for the direct impacts of policy by considering the OECD Environmental Policy Stringency (EPS) index (OECD, 2016) and mean regional political opinions as covariates.
-We find heterogeneous impacts of EU structural fund investments in low-carbon economies on GHG emissions reductions across European regions, with findings highlighting the variability of these effects by regional development levels, investment periods, and emission types.
-In developed and transition regions, low-carbon investments significantly lowered the long-term emission trends, in both per capita and relative to GDP, across the 2007-2013 and 2014-2020 programs.
-However, in the least developed regions, “low-carbon” investments had a positive (and significant) impact on emissions.
-Additionally, the effectiveness of reducing specific gas emissions, including CO2, CH4, and N2O, exhibits variability, indicating a complex interaction between investment efficiency and pollutant type.
-Our analysis also discerns significant but divergent impacts on the cyclical emissions components in transition economies between the two program periods.
-Hence, our analysis suggests a partial inefficiency of low-carbon public investments in Europe.
-It calls for a refined spatial and temporal allocation of these investments.
-The heterogeneous impacts of the climate action efforts derived from the ESI funds can be attributed to several critical factors.
-These factors include the inherent complexities of reducing emissions, the intrinsic differences among European regions in terms of capabilities and resource endowments, and the relevance of development paths, as suggested by the Kuznets environmental curve (Bataille et al., 2021; Saraji and Streimikiene, 2023).
-For instance, it is possible that any additional funding may increase energy consumption and GHG emissions for investments below a certain level.
-This is rooted in how investments are targeted and their immediate impacts, which might not always be aligned with long-term sustainability goals.
-In many developing regions, the primary focus of initial investments may be on establishing basic infrastructure and boosting economic output quickly to meet immediate developmental needs.
-Additionally, low-carbon economy investments typically require substantial up-front costs and the deployment of heavy infrastructure, which can drive significant energy consumption and emissions.
-The transition to sustainable energy systems includes upgrading grids, developing renewable energy sources, and improving energy resilience.
+Our analysis centres on two emission indicators: per capita emissions and emissions intensity (relative to regional gross domestic product (GDP)).
+We focus on the relationship between economic growth and environmental sustainability, and, to that end, we decompose the emission time series of each region into trend and cycle components.
+This allows us to determine expenditures’ short-term and long-term effects, considering that some regions have significantly advanced in terms of sustainability – achieving a given degree of decoupling – whereas many others have yet to do so.
+Our analysis also accounts for the direct impacts of policy by considering the OECD Environmental Policy Stringency (EPS) index (OECD, 2016) and average regional political opinions as covariates.
+We find heterogeneous effects of low-carbon expenditures on GHG emissions across European regions.
+These impacts vary depending on the regional development level, investment period, and emission indicator.
+In the least developed regions, low-carbon expenditures are consistently associated with increases in long-term emissions per capita and, to a lesser extent, in emissions intensity.
+In contrast, results for developed and transition regions do not reveal significant reductions in long-term emissions trends, although there is some evidence of short-term mitigating effects, particularly for the 2007–2013 programme.
+To address potential endogeneity, we implement a two-stage least squares (2SLS) model using lagged structural indicators and regional development categories as instruments.
+The instrumental variables estimates reinforce the main findings: low-carbon investments have heterogeneous effects across regions, with increases in per capita emissions in less developed regions and short-term reductions in developed regions.
+In transition regions, distinct patterns emerge that ultimately underscore the importance of the regional context.
+When disaggregating emissions by gas type, we consistently find marked regional heterogeneity.
+The results reveal no significant reductions in emissions for developed regions, while in less developed regions, emissions – especially CO2 and N2O per capita – increase significantly.
+Effects in transition regions are mixed.
+Short-term dynamics also show variation by gas and region: investments under the 2007–2013 programme help reduce emissions volatility, particularly for CO2 in less developed and developed regions, and for N2O in transition regions.
+Overall, our findings point to important challenges in the effectiveness and coherence of current EU climate investment strategies, particularly those guided by the ESI funds, in supporting long-term decarbonisation in less developed regions.
+However, rather than signalling a failure of climate investment itself, the uneven performance may reflect the persistent structural regional disparities, including differences in institutional capacity, infrastructure, and development trajectories, as suggested by the Kuznets environmental curve (Bataille et al., 2021; Saraji and Streimikiene, 2023).
+Investments’ immediate impacts are not necessarily always aligned with long-term sustainability goals.
+For instance, it could be that investments below a certain level may increase energy consumption and raise GHG emissions.
+In many developing regions, investments are focused on establishing basic infrastructure to boost economic output and meet developmental needs.
+In addition, low-carbon economy investments might require substantial up-front costs (deploying heavy infrastructure), and sustainable energy systems demand upgrading grids, developing renewable energy sources, and improving energy resilience.
 Though essential in the long term, these investments can increase short-term emissions due to the intensive energy and resources required during their implementation phase.
-On another front, there is a risk that investments have been used inefficiently or inappropriately, for example, as a countercyclical buffer to stabilise the economy during recessions or by investing in large infrastructure projects (aka “white elephants”) that do not generate the expected impacts.
-Arguments like these could align with the heterogeneous effects of public investment on climate outcomes.
-Therefore, the differential impacts on GHG emissions have their origin in the cyclical nature of economic activities.
-Recent literature offers valuable insights into the interplay between environmental impacts and the business cycle.
-Doda (2014) finds that CO2 emissions increase during economic upturns, exhibit greater volatility compared to GDP, and show a procyclicality that is more pronounced in higher GDP per capita economies, suggesting wealthier economies have more stable emission patterns through economic cycles.
-Sheldon (2017) finds an asymmetrical response of emissions to economic changes attributed to the timing of capital investments and technological advancements, indicating the significant role of corporate decisions in environmental impacts.
-Moreover, Klarl (2020) finds CO2 emissions’ elasticity to GDP changes to be significantly higher during recessions in the United States, showcasing the dynamic elasticity of emissions in response to economic conditions.
-In addition, climate and energy policy tools have proven some degree of effectiveness, but their success often hinges on countries’ stringency and the complementarity among strategies.
+On another front, investments can be misused, for example, as a countercyclical buffer to stabilise the economy during recessions or in implementing large and inefficient infrastructures (aka “white elephants”).
+While arguments behind the heterogeneous impacts vary across regions, they are certainly shaped by economic conditions, particularly the economic cycle.
+A recent strand of literature offers insights into the interplay between environmental impacts and the business cycle.
+Doda (2014) finds that CO2 emissions increase during economic upturns, exhibit greater volatility compared to GDP, and show a procyclicality that is more pronounced in higher GDP per capita economies, suggesting that wealthier economies have more consistent emission patterns through economic cycles.
+Sheldon (2017) finds an asymmetrical response of emissions to the business cycle, which is attributed to the timing of capital investments and technological advancements.
+Moreover, Klarl (2020) finds CO2 emissions’ elasticity to GDP changes to be significantly higher during recessions in the United States.
+These results demonstrate the critical role of investment decisions in shaping environmental impacts.
+In addition, climate and energy policy tools have shown some degree of effectiveness, however, their success often relies on countries’ stringency and the complementarity among strategies.
 Diakoulaki and Mandaraka (2007) find that despite considerable efforts, the expected acceleration in decoupling post-Kyoto Protocol was not uniformly achieved.
 For Europe, Naqvi and Zwickl (2017) and Ivanova et al.
-(2017) also observe significant disparities in decoupling across countries, sectors, pollutants, and regions.
+(2017) find significant disparities in decoupling across countries, sectors, pollutants, and regions.
 Papież et al.
 (2022, 2021) further report mixed results in the EU’s long-term decoupling efforts, pointing to disparities between older and newer member states and emphasising the role of energy intensity and structure in emissions reductions.
-For OECD countries, Chen et al.
-(2018), focusing on the key factors affecting CO2 emissions, conclude that while reductions in energy intensity are associated with decreased CO2 emissions, economic growth inevitably tends to increase CO2 emissions.
-This implies that emissions reductions are hard to accomplish unless economic growth is accompanied by significant improvements in energy efficiency and a shift towards less carbon-intensive energy sources.
-Based on the literature discussed above, our paper contributes to the analysis of sustainable development and economic policy in European regions.
-By analysing how low-carbon investments from structural funds impact emissions reductions within business cycle fluctuations, we contribute to the understanding of decoupling economic growth from environmental degradation.
-Our results emphasise the importance of tailored investment strategies in achieving sustainable development goals, with a special focus on less developed regions to mitigate potential increases in emissions as they grow economically.
+In light of these findings, our paper contributes to the understanding of decoupling economic growth from emissions by analysing how low-carbon investments from the ESI funds impact emissions reductions within the business cycle.
 The remainder of this study is structured as follows.
-Section 2 discusses the relevant strands of literature and the conceptual framework of our study and situates our contribution.
+Section 2 discusses the relevant literature and the conceptual framework of our study and situates our contribution.
 Section 3 presents the data and methods.
 Section 4 presents the empirical results of the econometric estimations.
-Finally, Section 5 concludes by discussing the policy implications of our results.
+Section 5 discusses the results and provides policy implications.
+Finally, Section 6 concludes.
 2 Decoupling emissions: related literature
-Our study builds on the extensive literature analysing the relationship between economic growth and GHG emissions.
-The Environmental Kuznets Hypothesis indicates that there is a connection between a country’s income level and the quality of its environment: it can be said that an increase in wealth generates greater demand from the population for a cleaner and healthier environment.
-This demand, in turn, drives governments to implement and enforce stricter environmental laws and regulations to protect the environment.
-While wealthier countries tend to have the resources and political will to adopt and enforce these effective environmental policies, regions with more constrained economic resources may find it more challenging to establish and maintain high environmental standards, resulting in more significant environmental challenges (O’Connor and Turnham, 1992; Doyle and Stiglitz, 2014).
+Our study builds on the extensive literature on the connection between economic growth and GHG emissions.
+Drawing from the Environmental Kuznets Hypothesis, it is assumed that an increase in wealth generates greater demand from the population for a cleaner and healthier environment.
+This demand drives governments to implement stricter environmental regulations to protect the environment.
+In this way, while wealthier countries tend to have the resources and political will to adopt these policies, regions with more constrained economic resources may find it more challenging to establish high environmental standards (O’Connor and Turnham, 1992; Doyle and Stiglitz, 2014).
 Certainly, economic development can lead to improved environmental outcomes by decoupling emissions from economic output.
@@
 Shuai et al.
-(2017) analysed the country-by-country turning points of the carbon Kuznets curve, finding a positive correlation between a country’s income level and its likelihood to fit the Kuznets hypothesis, with higher income economies achieving more quick tipping points than their poorer counterparts.
-While economic development level is linked to faster environmental improvement transitions, other variables also play a relevant role.
+(2017) analyse the country-by-country turning points of the carbon Kuznets curve, finding a positive correlation between a country’s income level and its likelihood to fit the Kuznets hypothesis, with higher income economies achieving quicker tipping points than their poorer counterparts.
+While economic development level is linked to faster environmental improvement transitions, other variables such as consumption patterns and regional characteristics also play a relevant role.
 Ivanova et al.
-(2017) analyses household consumption across 177 regions in 27 EU countries, finding wide variations in per capita GHG emissions across regions (0.6 to 6.5 tCO2e/cap in 2007).
+(2017) analyse household consumption across 177 regions in 27 EU countries, finding wide variations in per capita GHG emissions across regions (0.6 to 6.5 tCO2e/cap in 2007).
 Regarding the drivers behind carbon footprints, the authors identified the relevance of household size, whether an area is urban or rural, the population’s education level, expenditure patterns, local temperature, resource availability, and the carbon intensity of the local electricity mix.
-Yet a significant strand of literature doubts that the downward-sloping relationships between higher-income countries and pollution levels – which might imply a decoupling of the economic activity with emissions – is enough to face climate change efficiently.
+Yet a significant strand of literature doubts that the downward-sloping relationships between higher-income countries and pollution levels – which might imply a decoupling of economic activity with emissions – are enough to face climate change efficiently.
 Certainly, there are concerns about the possible shift in the production of pollution-intensive goods from developed to developing countries with less stringent environmental regulations (Hertwich and Peters, 2009; Levinson, 2023).
@@
 While this points out that high-income countries should also look into reducing aggregate demand as a strategy for significant emissions reductions (Knight and Schor, 2014), it also requires the implementation of agreements at different scales: regional and global (Oberthür and Ott, 1999).
-A growing body of evidence indicates that such policies and international agreements are increasingly impacting emission trends, as reflected in the economic practices within countries, contributing to our understanding of the decoupling mechanism (Fischer and Heutel, 2013; Hildén et al., 2014; Haberl et al., 2020).
+In this context, a growing body of literature indicates that international environmental agreements and climate policies are increasingly impacting emission trends, as reflected in the economic practices within countries, contributing to our understanding of the decoupling mechanism (Fischer and Heutel, 2013; Hildén et al., 2014; Haberl et al., 2020).
 Decoupling analyses aim to categorise the patterns of how emissions respond to economic growth – whether they increase at a slower rate, decrease, or stabilise as the economy grows (relative and absolute decoupling) or increase faster than the economy (coupling).
@@
 Using information from OECD countries for 2001-2015, Chen et al.
-(2018) analyse how economic growth can be decoupled from CO2 emissions, identifying three key forms: recessive, where emissions drop due to efficiency gains or energy mix shifts, not economic growth; weak negative, where efficiency improvements do not fully offset emissions from economic expansion; and strong, where economic and population growth do not lead to higher emissions, thanks to effective policies, technological progress, and structural economic changes.
+(2018) analyse how economic growth can be decoupled from CO2 emissions, identifying three key forms: recessive, where emissions drop due to efficiency gains or energy mix shifts, not from economic growth; weak negative, where efficiency improvements do not fully offset emissions from economic expansion; and strong, where economic and population growth do not lead to higher emissions, thanks to effective policies, technological progress, and structural economic changes.
 Building on this understanding of the decoupling patterns, it is imperative to analyse emissions in the context of business cycles due to their substantial implications for environmental policy and economic management, particularly regarding the relationship between economic activity, regulation, emissions and energy consumption (Fischer and Heutel, 2013).
 In fact, emissions tend to be procyclical, increasing during economic booms and decreasing during recessions (Heutel, 2012).
-A fundamental question relates to the causal relationship between economic output and energy consumption, especially targeting business cycle fluctuations (Thoma, 2004; Narayan et al., 2011).
-Energy consumption notably responds to changes in industrial production and macroeconomic conditions, reinforcing the concept of the procyclicality of emissions relative to GDP observed in broader research.
-For instance, Thoma (2004) demonstrates that changes in macroeconomic conditions cause significant changes in electrical usage, particularly in the commercial and industrial sectors.
-Therefore, the link between economic growth, energy consumption, and GHG emissions is clearly established in the short term.
+A fundamental question relates to the causal relationship between economic output and energy consumption, especially targeting business cycle fluctuations (Narayan et al., 2011).
+Thoma (2004) shows that changes in macroeconomic conditions cause significant changes in electrical usage, particularly in the commercial and industrial sectors.
+These energy consumption responses reinforce the procyclicality of emissions relative to GDP.
+Therefore, the connection among economic growth, energy consumption, and GHG emissions is well established in the short term.
 For example, Peters et al.
@@
 Another fact is that emissions exhibit greater volatility compared to GDP, responding more dramatically to economic changes, and the extent to which emissions volatility exceeds GDP volatility decreases with the income level of the economies (Doda, 2014).
+These patterns are also observed at the subnational level: Dueñas and Mandel (2025), analysing NUTS 2 regions in Europe, further show that CO2 emissions growth rates follow an asymmetric and fat-tailed distribution and, similarly, that emissions volatility scales inversely with economic size of regions.
 This implies that more stable emission patterns through economic cycles in developed regions could be linked to more efficient energy use, a greater share of services in the economy, and potentially more effective environmental regulations and policies.
@@
 He showed that emissions respond more sensitively to GDP changes during recessions than periods of economic normality, with elasticity often exceeding one in downturns.
-This indicates that CO2 emissions decrease more rapidly than economic output in recessions, highlighting the impact of economic cycles on environmental outcomes.
-The finding that emissions exhibit greater sensitivity during economic downturns suggests that the capacity of an economy to influence environmental outcomes can be significantly conditioned by its current economic state.
-In a similar vein, Naqvi (2021) finds that, on average, EU regions have managed to decouple emissions from economic output, with the most significant emission reductions occurring before the 2008 financial crisis.
+This indicates that CO2 emissions decrease more rapidly than economic output in recessions and also suggests that the capacity of an economy to influence environmental outcomes can be significantly conditioned by its current economic state.
+Thus, the literature has established the relevance of economic policy in addressing both the cyclical sensitivity of emissions to economic activity and the long-term decoupling of emissions from economic growth.
+This is of particular interest in the context of the European Union, where studies have started to examine the role of policy in shaping both short-term dynamics and long-term environmental trends.
+For example, Naqvi (2021) finds that, on average, EU regions have managed to decouple emissions from economic output, with the most significant emission reductions occurring before the 2008 financial crisis.
 After 2008, the trend towards decoupling weakened, with instances of re-coupling between emissions and economic output observed.
-The author uses the EPS index, finding that stronger environmental policies effectively reduce emissions.
-However, the effectiveness varies significantly across different types of emissions and is also influenced by the income levels of the regions.
+In addition, the author finds that environmental policy effectiveness varies significantly across different types of emissions and is influenced by the income levels of the regions.
 Similarly, with the aim of monitoring the effectiveness of the EU’s environmental policies, Delgado Rodríguez et al.
 (2018) developed the EU-cyclical environmental performance index by applying dynamic factor analysis over the period 1950 to 2012.
-This index demonstrates that the implementation of stricter emission targets, initiated particularly since the adoption of the 5th Environmental Action Programme (EAP), has effectively contributed to restraining the growth of CO2 emissions across the EU.
-Furthermore, the index reveals heterogeneity among the member states, distinguishing between countries that are leading in emission linkage – thus enhancing their alignment with EU environmental standards – and those lagging behind, which may require additional incentives or policy measures to improve compliance with EU regulations.
+This index shows that the implementation of stricter emission targets, initiated particularly since the adoption of the 5th Environmental Action Programme (EAP), has effectively contributed to restraining the growth of CO2 emissions across the EU.
+Furthermore, the index is substantially heterogeneous among the member states, distinguishing between countries that lead to emission linkage and those lagging behind.
 In addition, Papież et al.
 (2022) evaluate the EU’s energy policy effectiveness in achieving long-term decoupling and reports mixed findings.
-The study notes progress in detaching economic growth from emissions under the EU ETS and production-based accounting in both EU-15 and new member countries.
-However, they find no advancement in separating economic growth from consumption-based accounting emissions in either group.
-In summary, environmental policies must confront the broader economy, influenced by external macroeconomic shocks and internal technological changes.
-Under Kuznets’ hypothesis, decoupling might appear problematic since countries would only begin to improve their environmental performance after reaching a certain level of wealth, implying a dependency between economic growth and environmental improvement.
+The study notes progress in detaching economic growth from emissions under the EU ETS and production-based accounting, but they find no advancement in separating economic growth from consumption-based accounting emissions in either group.
+More directly related to the impacts of the EU Cohesion Policy, much of the literature has focused primarily on understanding its impact on regional development, specifically whether the funds have been effective in reducing regional disparities.
+For instance, very early on, Boldrin and Canova (2001) pointed out that these policies tend to be more redistributive than growth-oriented, reflecting political compromises over economic efficiency.
+However, investments in education and human capital, which have gained a significant portion of these funds, have shown medium-term positive impacts, emphasising the critical role of human capital development in fostering long-term regional growth and convergence (Rodríguez-Pose and Fratesi, 2004).
+Institutional quality plays a relevant role in the effectiveness of the funds.
+Evidence shown that local and regional governance directly affects economic growth and the effectiveness of structural funds (Rodríguez-Pose and Garcilazo, 2015; Charron et al., 2014).
+Although relatively few studies have explicitly focused on the effect of EU structural funds on reducing emissions, some anticipate that many of the same structural patterns that determine growth performance may also shape the trajectory of sustainable growth.
+Rodríguez-Pose and Bartalucci (2024) warn that while the green transition may benefit certain regions, it can also exacerbate existing territorial inequalities, particularly affecting less-developed regions.
+Nichifor et al.
+(2025) find that investments in renewables and environmental expenditures have positive effects on emissions reduction, but caution that these effects may exhibit diminishing returns beyond certain thresholds.
+In summary, in light of these findings, the effectiveness of the ESI funds in driving a low-carbon transition is conditional not only on structural factors such as institutional quality and regional development levels, but also on the phase of the economic cycle.
+Under Kuznets’ hypothesis, decoupling might appear problematic since regions would only begin to improve their environmental performance after reaching a certain level of wealth, implying a dependency between economic growth and environmental improvement.
 Therefore, seeking economic development without the corresponding environmental degradation is a great challenge for any policy framework.
+Therefore, the central hypothesis of this paper is as follows:
+Hypothesis: Low‑carbon economy expenditures from the European Structural and Investment Funds have heterogeneous impacts on GHG emissions across regions at different stages of development, influencing both the long‑run trend and the magnitude of the business‑cycle component of emissions.
+To test this hypothesis, rather than focusing on total emissions, we concentrate on per capita emissions and emissions intensity (relative to regional GDP) .
+These indicators enhance comparability across regions by accounting for differences in population and economic scale, while also isolating environmental efficiency more effectively.
 3 Data and methods
-We focus on emission data reported by the Emissions Database for Global Atmospheric Research (EDGAR), managed by the Joint Research Centre of the European Commission.
+We aim to analyse the impact of expenditure on GHG emissions resulting from expenditure on the thematic objective related to the low-carbon economy in the 2007-2013 and 2014-2020 ESI funds programmes (a more detailed definition of the expenditure variables is presented below).
+We shall use relevant covariates to control economic activity, sectoral composition, and environmental policy stringency.
+3.1 Data
+Emissions data
+We focus on emissions data reported by the Emissions Database for Global Atmospheric Research (EDGAR), managed by the Joint Research Centre of the European Commission.
 EDGAR provides quantitative estimates of global GHG emissions by country and sector, including CO2, CH4, N2O, and F-gases, using a methodology aligned with Intergovernmental Panel on Climate Change (IPCC) guidelines (Crippa et al., 2023).
 Specifically, we use European regional data at the NUTS 2 level of aggregation, focusing on aggregated GHG emissions and their breakdown into CO2, CH4, and N2O emissions.
-Our primary objective is to analyse the impacts on GHG emissions resulting from expenditure on the thematic objective related to the low-carbon economy in the 2007-2013 and 2014-2020 ESI funds programmes (a more detailed definition of the expenditure variables is presented below).
-In the econometric modelling setup, we use other relevant covariates to control for economic activity, sectoral composition, and local policy stringency that influence emissions across European regions.
-3.1 Data
 ESI funds and eligibility
@@
 Certain regions consistently receive higher funds according to ongoing economic development needs or projects requiring sustained investment.
-The disparities in the distribution of funds consistently with eligibility show that less developed regions in Southern and Eastern Europe generally receive higher per capita funding than Western and Northern Europe regions.
+The disparities in the distribution of funds, consistent with eligibility, show that less developed regions in Southern and Eastern Europe generally receive higher per capita funding than Western and Northern Europe regions.
 There are also variations in the distribution of funds between programming periods, which must be explained by changes in policy focus.
@@
 Despite thematic changes between the two periods, the raw categorisation data allows for the creation of a comparative set of expenditure objectives to ensure comparability and coherence in our analysis.
-For the 2014-2020 period, the low-carbon expenditures are straightforward to determine since there is a categorical variable related to thematic objective 4 (TO-4), labelled as “low-carbon economy”.
+For the 2014-2020 programme, the low-carbon expenditures are straightforward to determine since there is a categorical variable related to thematic objective 4 (TO-4), labelled as “low-carbon economy”.
 The aim of the expenditure can be tracked by the reported “intervention field” in the data.
 Specifically, related to TO-4, the intervention fields include renewable energy (wind/solar/biomass), energy efficiency renovation of public infrastructure, cycle tracks and footpaths, among others (see Table A1 in the Appendix).
-For the 2007-2013 program, there is no categorical variable explicitly associated with the low-carbon economy as in the 2014-2020 program.
-However, the raw categorisation data describes the “priority” related to the reported expenditures, serving as a proxy for the aim of the expenditure.
+For the 2007-2013 programme, there is no categorical variable explicitly associated with the low-carbon economy as in the 2014-2020 programme.
+However, the raw categorisation data describe the “priority” related to the reported expenditures, serving as a proxy for the aim of the expenditure.
 We use this information to create a low-carbon expenditures variable for this period.
@@
 While reported expenditures can be easily tracked at the regional level, the reports correspond to the entire programming period.
-To our knowledge, there is no disaggregated, time-varying consolidated official information.
-Nevertheless, the cohesion data platform presents estimations of the expenditure by NUTS 2, which are aggregated for the most significant types of structural funds only.
-Therefore, to address our need for yearly data, we have opted for two approaches: one assumes that expenditures are distributed evenly over time (dividing the reported expenditures over the programming period), and the other assumes they follow the same pattern as the total aggregated expenditure reported by the cohesion data platform.444Data available at: https://cohesiondata.ec.europa.eu/Other/Historic-EU-payments-regionalised-and-modelled/tc55-7ysv/about_data.
-Since both strategies yielded qualitatively similar conclusions, we present in the main text the results obtained with the first approach, while those derived from the second are included in the Appendix as a robustness check (Tables A2 and A3).
+To our knowledge, there is no disaggregated time-varying consolidated official information available.
+Given this constraint, we built an annually growing cumulative investment series for each region and programme.
+This implies that, within each programming period, the total per capita expenditure is allocated progressively across the corresponding years.
+While this introduces a smoothing assumption, it allows us to preserve the annual variation in emissions and covariates.
+Therefore, considering the capital-intensive nature of the funded projects, we argue that their environmental effects are plausibly associated with the cumulative stock of investments.
+We further validate this modelling choice through robustness checks, including alternative cross-sectional estimations presented in Appendix Tables A2 and A3.
 Supplementary covariates
 Economic variables were obtained from the ARDECO (the Annual Regional Database of the European Commission’s Directorate General for Regional and Urban Policy), which contains annual data at the regional level and is maintained and updated by the Joint Research Centre of the European Commission.
-We use variables such as the real gross domestic product (GDP), the population, and the real gross value added (GVA) for all industry sectors according to NACE Rev.
+We use variables such as the real gross domestic product, the population, and the real gross value added (GVA) for all industry sectors according to NACE Rev.
 2 definitions.
@@
 To assess the effects of environmental policies on emissions, we use the EPS index (OECD, 2016).
-This index, developed by Brunel and Levinson (2013) and further refined by Botta and Koźluk (2014), provides a standardised composite measure incorporating both market and non-market-based policies, covering 27 OECD countries over the period from 1990 to 2020.
+This index, developed by Brunel and Levinson (2013) and further refined by Botta and Koźluk (2014), provides a standardised composite measure incorporating both market and non-market-based policies, covering 27 OECD countries between 1990 and 2020.
 Finally, to determine the political ideology at the regional level, we calculate the regional weighted average of political ideologies using the results of European subnational election data.
-To measure political ideology, we use the “left/right” dimension variable reported by ParlGov (Döring et al., 2022) and complemented by CHES-Europe (Jolly et al., 2022), which takes values in the interval 0–10, with values close to zero representing extreme left ideology and values close to 10 representing extreme right ideology.
+To measure political ideology, we use the “left/right” dimension variable reported by ParlGov (Döring et al., 2022) and complemented by CHES-Europe (Jolly et al., 2022), which takes values in the interval 0-10, with values close to zero representing extreme left ideology and values close to 10 representing extreme right ideology.
 We use EU-NED (Schraff et al., 2023), which reports sub-national election data across European countries over the last three decades at the party level.
 We assign each party at the regional level an ideology by combining information from ParlGov and CHES-Europe.
-These databases are integrated within the Party Facts platform, which facilitates the incorporation of party-level data.555The data is available at: https://www.chesdata.eu/ches-europe and https://www.parlgov.org/.
+These databases are integrated within the Party Facts platform, which facilitates the incorporation of party-level data.444The data are available at: https://www.chesdata.eu/ches-europe and https://www.parlgov.org/.
 Also see: https://partyfacts.herokuapp.com/.
@@
 Distribution of emissions and low-carbon economy funds
-Figure 2 presents a set of maps illustrating different facets of the environmental impact and distribution of the Cohesion Funds.
+Figure 2 presents a set of maps illustrating different facets of the environmental impact and distribution of the ESI funds.
 In the upper part, three maps show average GHG emissions in shades of blue to red, where blue denotes lower levels and red denotes higher levels.
@@
 These maps show that, on average, towards the end of the 2010s, most European regions emitted less than 50 teragrams of CO2.
-However, there is significant heterogeneity in per capita emissions, highlighting areas of intense economic activity or low population density.
-Comparing the early 2000s and late 2010s, it is evident that many regions with initially high emissions tended to reduce them (see, for instance, the Castile region in Spain), while some regions maintained or increased their per capita emission levels (for instance, the Masovian region in Poland).
+However, there is significant heterogeneity in per capita emissions, highlighting areas of intense economic activity or low population density.555Sparsely populated regions may have high per capita emissions despite relatively moderate total emissions.
+This is due to emission-intensive activities, like agriculture and energy production, concentrated in some low-density areas.
+Comparing the early 2000s and late 2010s, it is evident that many regions with initially high emissions tended to reduce them (for instance, the Castile region in Spain), while some regions maintained or increased their per capita emission levels (for instance, the Masovian region in Poland).
 The lower maps in Figure 2 show the calculated funds allocated to promote a low-carbon economy during 2007-2013 and 2014-2020 programmes, with the amounts received per capita in colour scales ranging from less than 100 euros to almost 1,000 euros per person.
 In both periods, most of the funds went to LDRs, with significant differences in the allocated amount, in agreement with the results presented in Figure 1.
-This pattern raises questions about the effectiveness of the allocation of funds relative to emission levels.
-In Figure 3, we plot the total GHG emissions per capita versus the total expenditure in low-carbon economy funds per capita.
-Emissions are considered at the beginning of the corresponding programming period.
-The plot exhibits great moderation in emission levels per capita for most European regions and shows that higher funds per capita were consistently allocated to less developed regions.
-However, a positive correlation pattern in the allocation is less clearly seen in both panels, suggesting that these funds are not necessarily allocated to the regions with the highest need to reduce emissions.
 As a preliminary analysis, we estimate the nexus between expenditure on low-carbon economy funds and total GHG emissions using the following regional-level model:
-where is an emission indicator for region at time , and and are the cumulative yearly programme expenditures per capita (in log) for the periods 2007-2013 and 2014-2020, respectively, with denoting lags to acknowledge potential delayed effects.
+where is an emission indicator for region at time , and and are the cumulative yearly programme expenditures per capita for the periods 2007-2013 and 2014-2020, respectively, with denoting lags to acknowledge potential delayed effects.
 We include fixed effects for location () and time () to control for time-invariant unobserved heterogeneity across regions as well as uniform temporal shocks, such as those associated with environmental directives instituted at the European Union level.
 Finally, the term is the estimation residual.
-In the left-hand in Eq.
+In the left-hand side of Eq.
 (1), the term serves as a reference notation for different emission variables.
@@
 | (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | |
-| Low Carbon Exp. pc (log) | 0.004 | 0.081∗∗ | 0.130∗∗ | 0.042 | ||||
+| Low Carbon Exp. pc | 0.004 | 0.081∗∗ | 0.130∗∗ | 0.042 | ||||
 | (0.063) | (0.036) | (0.064) | (0.038) | |||||
-| Low Carbon Exp. pc (log) | 1.169∗∗∗ | 0.008 | 0.438∗∗∗ | 0.009 | ||||
+| Low Carbon Exp. pc | 1.169∗∗∗ | 0.008 | 0.438∗∗∗ | 0.009 | ||||
 | (0.091) | (0.053) | (0.093) | (0.055) | |||||
-| Low Carbon Exp. pc (log) | 0.149∗∗ | 0.052 | 0.123∗∗ | 0.028 | ||||
+| Low Carbon Exp. pc | 0.149∗∗ | 0.052 | 0.123∗∗ | 0.028 | ||||
 | (0.062) | (0.035) | (0.063) | (0.037) | |||||
-| Low Carbon Exp. pc (log) | 1.633∗∗∗ | 0.067 | 0.601∗∗∗ | 0.091 | ||||
+| Low Carbon Exp. pc | 1.633∗∗∗ | 0.067 | 0.601∗∗∗ | 0.091 | ||||
 | (0.140) | (0.080) | (0.142) | (0.084) | |||||
 | Observations | 3,872 | 3,872 | 3,872 | 3,872 | 3,872 | 3,872 | 3,872 | 3,872 |
-| Adjusted R2 | 0.382 | 0.375 | 0.179 | 0.178 | 0.580 | 0.580 | 0.039 | 0.039 |
 | Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||||
-This preliminary exercise demonstrates that the relationship between low-carbon expenditures and GHG levels, as well as the growth rates of emissions per capita (GHGEpc) and emission intensity (GHGEi), can be mixed.
 These mixed results suggest that the econometric specification might be masking underlying factors that could restrict the effectiveness of low-carbon funds.
 One relevant limiting factor is the persistent disparities among European regions.
-Indeed, the development level of regions is a key factor influencing the effectiveness of cohesion funds (Boldrin and Canova, 2001; Rodríguez-Pose and Fratesi, 2004).
+Indeed, the development level of regions is a key factor influencing the effectiveness of the ESI funds (Boldrin and Canova, 2001; Rodríguez-Pose and Fratesi, 2004).
 However, many aspects related to the heterogeneity of European regions can explain the mixed results.
@@
 Our methodology shifts from this type of analysis, which primarily assesses the impact of GDP growth on emissions, to a comparative panel data analysis focusing on the effects of expenditure in the low-carbon economy.
-We specifically highlight the role of cohesion fund investments in fostering low-carbon economies to curb emissions.
+We specifically highlight the role of the ESI funds in fostering low-carbon economies to curb emissions.
 Therefore, it is central in our methodology to decompose time series variables, such as emissions, GDP, and GVA, into their cyclical and trend components for each region independently.
-Then we consider these time series in two different panel data model frameworks: regional trends and regional cycles of emissions.
+Then, we consider these time series in two different panel data model frameworks: regional trends and regional cycles of emissions.
 In this way, we analyse the difference between long-term trends and cyclical effects of the impact of low-carbon economy funds on GHG emissions.
@@
 Therefore, the total emissions time series of the region , , is decomposed in its trend component and its cycle component .
-After decomposing the GHGEpc and GHGEi time series, we independently analyse the panels of regional emission trends and cycles, considering the regional development level differences.
-The model for emissions trends is defined as follows:
-where the superscript specifies that we conduct the estimation on time series trend components.
-This way, is the trend of the logarithm of emissions per capita for NUTS 2 region at time , similarly is the trend component of the GDP (in log) of region at time , and and are the yearly cumulative programme expenditures (in log) at time .
-We include the development level of region using the dummy variable , where .
-The term consider additional control variables such as the dummy variables capturing the effect of stronger stringency of environmental policies, denoted by 666Following Naqvi (2021), we measure the impact of stronger environmental policies employing a dummy variable that takes the value of one when EPS index is greater or equal than 2.5., and the term , which indicates the average political ideology of the region, determined by the outcomes of local elections (see data section for details).
-This model incorporates fixed effects for both location () and time (), removing biases related to unobserved fixed regional heterogeneity and factors that vary over time but impact all regions equally.
-Additionally to Eq. (3) focusing on the trend of regional emissions, we also consider the impact of low-carbon investments on the cyclical components of emissions using the following specification:
-Note that the superscript now represents the emissions’ cyclical components, therefore, is the cycle component of the GDP (in log) of region at time .
-In our specifications, Eq.
-(3) and (4), the parameters of interest are , , and , which capture how a region’s development level can influence structural funds’ effect in emissions components.
-The aim is to explore whether the impact of these expenditures varies according to the economic development level of the region, affecting the cyclical and trend components of emissions.
-As a robust check, in our specifications, we replace GDP with the set of GVA variables for all economic sectors.
-This allows us to better control how regional productive structures affect GHG emissions.
+After decomposing the GHGEpc and GHGEi time series, we independently analyse the panel data of regional emission trends and cycles, taking into account differences in regional development levels.
+The model for the trend and cycle components of emissions is defined as follows:
+where the superscript indicates that we estimate the model separately for the trend and cycle components of the series.
+Specifically, denotes either the trend or the cycle component of the logarithm of an emissions indicator (GHGEpc or GHGEi)666 Using these relative measures offers several advantages: first, per‑capita emissions normalise for population size, allowing us to compare each region’s carbon footprint regardless of its scale; second, emissions intensity controls for economic scale, highlighting how efficiently regions produce output per unit of carbon.
+This enhances comparability across diverse regions and reduces heteroskedasticity in panel regressions.
+More importantly, it allows us to better isolate the impact of low‑carbon investments on environmental efficiency and avoid confounding with demographic or economic differences.
+for region at time , and is the corresponding component of the logarithm of GDP.
+The terms and refer to the cumulative yearly expenditures of the 2007–2013 and 2014–2020 programmes, respectively, lagged by years.
+Regional development levels are introduced via the dummy variable , with representing less developed, transition, and developed regions, respectively.
+The vector includes additional controls, such as a dummy for environmental policy stringency ()777Following Naqvi (2021), we use a dummy equal to one when the EPS index is greater than or equal to 2.5.
+and an indicator of average political ideology (), based on local election outcomes.
+The model also incorporates region fixed effects () and year fixed effects () to control for time-invariant regional characteristics and common shocks across all regions.
+As a robustness check, we replace GDP with the set of GVA variables for all economic sectors to better account for the role of regional productive structures in explaining GHG emissions.
+While our specification includes key relevant variables, we acknowledge that endogeneity may still pose concerns.
+For instance, one might argue that environmental regulations or improvements in energy efficiency could, in principle, influence regional output and investment decisions.
+For these reasons, we have included a five-year lag for all explanatory variables, ensuring that the predictors precede the observed emissions outcomes by a meaningful temporal distance.
+Therefore, while our specification mitigates (but does not fully eliminate) simultaneity concerns, it relies on the timing assumption that lagged expenditures impact emissions.
+Moreover, the allocation of ESI funds is determined based on regional eligibility criteria related to development levels, making the investment variable exogenous to current emission levels.
+By interacting expenditures with regional development levels, we account for heterogeneous effects while controlling for potential confounding factors.
+Nonetheless, one could still argue that regions with higher emissions may have received more funds, representing a case of reverse causality or omitted variable bias.
+To test the robustness of our identification, we implement a 2SLS estimation using structural and institutional instruments plausibly uncorrelated with current emissions.
 4 Results
 4.1 Effect on emissions trend
-Table 2 presents the estimation results of Eq (3) using as the dependent variable the trends for GHGEpc or GHGEi.
-The results related to the expenditures for the different groups of regions are presented with respect to the most developed regions (DER).
-The table reveals mixed patterns that underline the importance of regional development level in the influence of expenditure.
+Table 2 presents the estimation results of Eq.
+(3) using as the dependent variable the trends for GHGE per capita or GHGE intensity.
+The key independent variables are low-carbon expenditures per capita from the 2007-2013 and 2014-2020 programmes, each introduced with a five-year lag.
+The heterogeneous impacts of regional development levels are captured by interacting these investment variables with dummy variables for transition (TER) and less-developed regions (LDR), using developed regions (DER) as the reference category.
 | Dependent variable: | ||||||
-| GHGEpc | GHGEi | |||||
-| (1) | (2) | (3) | (4) | (5) | (6) | |
-| Low Carbon Exp. pc | 0.158∗∗ | 0.138∗∗ | 0.012 | 0.159∗∗ | 0.171∗∗∗ | 0.108∗∗ |
+| GHGE per capita | GHGE intensity | |||||
+| (1t) | (2t) | (3t) | (4t) | (5t) | (6t) | |
+| Low Carbon Exp. pc | 0.039 | -0.024 | 0.065 | -0.383*** | -0.073 | -0.037 |
 | (0.177) | (0.143) | (0.102) | (0.122) | (0.154) | (0.126) | |
-| Low Carbon Exp. pc | 0.558∗∗ | 0.817∗∗ | 0.049 | 0.806∗∗∗ | 1.321∗∗∗ | 0.802∗∗ |
+| Low Carbon Exp. pc | -0.182 | -0.446 | 0.020 | -1.117 | -0.763 | -0.458 |
 | (0.586) | (0.733) | (0.680) | (0.848) | (0.673) | (0.641) | |
-| Low Carbon Exp. pcTER | 0.029 | 0.111 | 0.127 | 0.111 | 0.059 | 0.004 |
+| Low Carbon Exp. pcTER | -1.081* | -0.491 | -0.638 | 0.065 | -0.337 | -0.375 |
 | (0.644) | (0.790) | (0.862) | (0.604) | (0.517) | (0.567) | |
-| Low Carbon Exp. pcTER | 2.156∗∗∗ | 1.177 | 3.539∗∗∗ | 2.539∗∗∗ | 1.974∗∗ | 2.895∗∗∗ |
+| Low Carbon Exp. pcTER | -1.200 | -0.409 | -1.371 | 0.426 | -0.608 | -1.005 |
 | (1.613) | (1.585) | (1.561) | (1.461) | (1.510) | (1.400) | |
-| Low Carbon Exp. pcLDR | 0.730∗∗∗ | 0.620∗∗∗ | 0.489∗∗∗ | 0.361∗∗∗ | 0.356∗∗∗ | 0.315∗∗∗ |
+| Low Carbon Exp. pcLDR | 0.611*** | 0.548*** | 0.439** | 0.458** | 0.383 | 0.327 |
 | (0.226) | (0.200) | (0.185) | (0.219) | (0.259) | (0.248) | |
-| Low Carbon Exp. pcLDR | 2.470∗∗∗ | 3.329∗∗∗ | 2.468∗∗∗ | 1.382∗∗∗ | 2.154∗∗∗ | 1.666∗∗∗ |
+| Low Carbon Exp. pcLDR | 1.315** | 1.160* | 0.728 | 0.668 | 0.784 | 0.497 |
 | (0.587) | (0.681) | (0.644) | (0.810) | (0.632) | (0.611) | |
-| GDP | 0.209∗∗∗ | 0.297∗∗∗ | 0.631∗∗∗ | 0.525∗∗∗ | ||
+| GDPt-5 | 0.238*** | -0.462*** | ||||
+| (0.065) | (0.102) | |||||
+| EPSt-5 | 0.000 | 0.001 | 0.001 | 0.001 | ||
 | (0.007) | (0.006) | (0.007) | (0.007) | |||
-| EPS | 0.009∗∗ | 0.011∗∗ | 0.005 | 0.010∗∗ | ||
+| Political Ideologyt-5 | 0.014 | 0.007 | -0.011 | -0.009 | ||
 | (0.010) | (0.009) | (0.016) | (0.014) | |||
-| Political Ideology | 0.016∗∗∗ | 0.001 | 0.002 | 0.007∗∗ | ||
+| GVAA,t-5 | 0.065* | 0.034 | ||||
+| (0.034) | (0.038) | |||||
+| GVAB-E,t-5 | 0.242*** | -0.042 | ||||
+| (0.047) | (0.075) | |||||
+| GVAF,t-5 | -0.135*** | -0.153*** | ||||
+| (0.041) | (0.057) | |||||
+| GVAG-J,t-5 | 0.129 | -0.200 | ||||
+| (0.091) | (0.132) | |||||
+| GVAK-N,t-5 | -0.362*** | -0.391*** | ||||
+| (0.084) | (0.100) | |||||
+| GVAO-U,t-5 | 0.170 | 0.310** | ||||
+| (0.108) | (0.134) | |||||
+| Observations | 3,872 | 3,408 | 3,408 | 3,872 | 3,408 | 3,408 |
+| Notes: Robust standard errors clustered at the NUTS 2 level are reported in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. | ||||||
+The estimates for per capita emissions (columns 1t–3t) show that in developed regions, the reference group, the coefficients on low-carbon expenditures are small and not statistically significant across all models.
+In contrast, the interaction terms for less developed regions are consistently positive and significant for the 2007–2013 programme, suggesting that greater expenditure is associated with increases in long-run emissions per capita.
+The interaction terms for the 2014–2020 period are also positive, although they are significant only in models 1t and 2t.
+For transition regions, the interaction coefficients are negative in most cases, but not statistically significant, except marginally in model 1t for 2007–2013 expenditures.
+Regarding emissions intensity (columns 4t-6t), we observe a statistically significant negative effect of 2007–2013 expenditures in developed regions only in model 4t, suggesting improved emissions efficiency.
+However, for less developed regions, the corresponding interaction terms are again positive and significant in model 4t.
+As in the GHGEpc models, results for transition regions remain statistically not significant across all models.
+Control variables, particularly GDP and sectoral GVA, exhibit expected significant relationships with emissions trends.
+Higher GDP is linked positively with emissions per capita but negatively with emissions intensity, highlighting the scale versus efficiency effects inherent in emissions analysis.
+Sectoral GVA variables reveal heterogeneity across economic activities.
+For instance, GVA in sectors B–E (industry) is positively correlated with per capita emissions, while GVA in services (sectors K–N) shows a significant negative association with both emissions indicators.
+These results underline the importance of regional economic structure in shaping emissions trajectories.
+As regards institutional and political controls, we find that environmental policy stringency and political ideology show no statistically significant effects.
+The small positive coefficients on EPS suggest that national-level measures may not strongly influence regional emissions trends in per capita or intensity terms.
+Similarly, political orientation, as proxied by the left-right index, does not exhibit consistent effects.
+4.2 Effect in the emissions cycle
+Table 3 presents the estimation results of Eq. (3), using the cyclical components of GHG emissions per capita and emissions intensity as dependent variables.
+| Dependent variable: | ||||||
+| GHGE per capita | GHGE intensity | |||||
+| (1c) | (2c) | (3c) | (4c) | (5c) | (6c) | |
+| Low Carbon Exp. pc | -0.017 | -0.055** | -0.065** | 0.027 | 0.004 | -0.002 |
+| (0.021) | (0.028) | (0.027) | (0.039) | (0.030) | (0.030) | |
+| Low Carbon Exp. pc | -0.128 | -0.189 | -0.200 | -0.092 | -0.168 | -0.180 |
+| (0.230) | (0.235) | (0.206) | (0.249) | (0.234) | (0.224) | |
+| Low Carbon Exp. pcTER | -0.166 | -0.130 | -0.131 | -0.107 | 0.048 | 0.043 |
+| (0.139) | (0.117) | (0.114) | (0.173) | (0.088) | (0.093) | |
+| Low Carbon Exp. pcTER | 0.079 | -0.096 | -0.077 | -0.358 | -0.664 | -0.662 |
+| (0.746) | (0.738) | (0.704) | (0.817) | (0.720) | (0.725) | |
+| Low Carbon Exp. pcLDR | -0.100** | -0.110** | -0.130*** | -0.078 | -0.096** | -0.104** |
+| (0.050) | (0.048) | (0.047) | (0.054) | (0.042) | (0.041) | |
+| Low Carbon Exp. pcLDR | 0.198 | 0.292 | 0.333 | 0.084 | 0.221 | 0.239 |
+| (0.238) | (0.240) | (0.217) | (0.253) | (0.234) | (0.224) | |
+| GDPt-5 | -0.251*** | 0.082** | ||||
+| (0.037) | (0.035) | |||||
+| EPSt-5 | 0.010*** | 0.011*** | 0.012*** | 0.012*** | ||
 | (0.003) | (0.003) | (0.003) | (0.003) | |||
-| GVAA | 0.044∗∗∗ | 0.074∗∗∗ | ||||
+| Political Ideologyt-5 | 0.005 | 0.003 | 0.003 | 0.003 | ||
+| (0.003) | (0.003) | (0.003) | (0.003) | |||
+| GVAA,t-5 | -0.012 | -0.012 | ||||
 | (0.009) | (0.007) | |||||
-| GVAB-E | 0.210∗∗∗ | 0.089∗∗∗ | ||||
+| GVAB-E,t-5 | -0.011 | -0.003 | ||||
 | (0.018) | (0.018) | |||||
-| GVAF | 0.016 | 0.095∗∗∗ | ||||
+| GVAF,t-5 | -0.044*** | 0.017 | ||||
 | (0.014) | (0.017) | |||||
-| GVAG-J | 0.0001 | 0.111∗∗∗ | ||||
+| GVAG-J,t-5 | 0.030 | 0.080** | ||||
 | (0.033) | (0.032) | |||||
-| GVAK-N | 0.327∗∗∗ | 0.506∗∗∗ | ||||
+| GVAK-N,t-5 | -0.044 | 0.024 | ||||
 | (0.041) | (0.038) | |||||
-| GVAO-U | 0.324∗∗∗ | 0.330∗∗∗ | ||||
+| GVAO-U,t-5 | -0.203*** | -0.058 | ||||
 | (0.041) | (0.042) | |||||
-| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
-| Adjusted R2 | 0.534 | 0.626 | 0.664 | 0.791 | 0.798 | 0.820 |
-| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. | ||||||
-Therefore, for developed regions, the reference group, the effect of expenditures in both periods indicates a negative and significant impact in all specifications, except in model 3, where the effect is not significant for both periods.
-Similarly, the effect of the interaction term for transition regions is negative in all specifications, except in model 2, being non-significant for expenditures associated with the period 2007-2013 and significant for expenditures in the period 2014-2020.
-This implies that the net impact for transition regions is negative, meaning that, for both transition and developed regions, the long-term effect of expenditures is expected to result in a decrease in emissions, both per capita and in relation to GDP.
-In contrast, a significant positive effect is observed for less developed regions, reversing the trend seen in the other development groups.
-This is because the estimates of the expenditures interaction terms in developing regions are not only positive and significant in all specifications but also cancel any negative effect estimated for the reference group.
-While explaining the mechanism behind this perverse outcome is far from the scope of this paper, this could indicate that, in less developed regions, expenditures have effectively driven economic development, albeit possibly at the cost of an increase in emissions in the long term.
-It is important to note that as these regions are developing, the increase in emissions resulting from factors like industrialisation, which is often carbon-intensive, cannot be ignored.
-Regarding the control variables, GDP is positively related to GHGEpc and GHGEi, implying that economic growth can lead to an increase in per capita emissions in the long term and to a more efficient economy in terms of emissions relative to GDP.
-The EPS dummy has a small but positive impact on per capita emissions and negative emission intensity.
-Although this positive impact is unexpected, it may also indicate that the effect of more stringent policies may be observed thereafter.
-Indeed, Naqvi (2021) analysed the cumulative effects of EPS dummy on different emission components, showing that while CH4 and N2O emissions responded strongly to policies and declined continuously across regions, CO2 emissions declined two years after implementing the policies.
-In the Appendix, we show that the EPS dummy has a negative impact on CH4 and N2O while positive for CO2, the latter being the main component of the total GHG emissions used in this paper (Tables A4 to A6).
-The effect of the average political ideology is not robust; while it is found positive for GHGEpc, except in model 3, for GHGEi, it is positive and not significant in model 5 and negative and significant in model 6.
-These mixed results connect to the complex relationship between the effectiveness of sustainable policies and political ideologies.
-In this respect, for instance, Garmann (2014) finds that in 19 OECD countries from 1992 to 2008, right-wing governments were associated with fewer emission reductions than the centre and left-wing governments.
-However, he also observes that emissions were higher in governments with more participating parties, although the distinction between majority and minority governments did not significantly influence emissions.
-The effect related to the gross value added by sectors reveals a differentiated impact on GHGEpc and GHGEi.
-For instance, the GVA of the agriculture sector, GVAA, has a negative impact on both GHGEpc and GHGEi, which suggests that the trend of the agricultural sector production contributes to a reduction of the emissions trends.
-While the GVA of the manufacturing industry, GVAB-E has a significant positive effect on GHGEpc and a significant negative effect on GHGEi, indicating that greater production in the manufacturing industry implies higher emissions per capita but lower emission intensity.
-4.2 Effect in the emissions cycle
-Table 3 presents the results of the regressions for the cyclical part of per capita emissions and emissions intensity.
-The results show mixed effects associated with the expenditures of both periods and the development level interacted variables.
-These effects are presented in reference to the most developed regions.
-Therefore, for developed regions, we find that the effect of expenditure in 2007-2013 is negative, although only significant in model 1.
-The effect of expenditure in 2014-2020 is not significant, but in specifications (2-3 and 5-6), the estimated effect is positive.
-These results hint that expenditure in the low-carbon economy has no robust effect on the emission cycle for the most developed regions.
-| Dependent variable: | ||||||
-| GHGEpc | GHGEi | |||||
-| (1) | (2) | (3) | (4) | (5) | (6) | |
-| Low Carbon Exp. pc | 0.081∗ | 0.027 | 0.034 | 0.068 | 0.025 | 0.022 |
+| Observations | 3,872 | 3,408 | 3,408 | 3,872 | 3,408 | 3,408 |
+| Notes: Robust standard errors clustered at the NUTS 2 level are reported in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
+The estimates for per capita emissions cycles (columns 1c–3c) show that in developed regions, the direct effect of low-carbon expenditures is negative and statistically significant for the 2007–2013 programme in models 2c and 3c, suggesting that these investments helped smooth short-term fluctuations.
+Notably, the interaction terms for less-developed regions are also negative and statistically significant across all three models for the same programming period.
+This indicates that, despite the positive long-term trend effect observed in less-developed regions (see Table 2), low-carbon investments during the 2007–2013 programme contributed to dampening short-term emissions volatility.
+The coefficients for transition regions are generally negative but not statistically significant.
+As regards emissions intensity cycles (columns 4c–6c), there is no significant direct effect of low-carbon spending for developed or transition regions.
+However, for less-developed regions, interaction terms for the 2007–2013 programme are again negative and statistically significant (in models 5c and 6c).
+As in the per capita models, the effects of the 2014–2020 programme are not statistically significant for any regional group.
+Lagged GDP is negatively associated with per capita emissions cycles and positively with emissions intensity cycles, consistent with short-run adjustments in the emissions-output relationship.
+Sectoral GVA variables display heterogeneous effects: GVA in construction (sector F) is associated with a reduction in emissions per capita, while services (G–J) appear positively associated with emissions intensity in model 6c.
+Other sectors are not significantly correlated with the cyclical components of emissions.
+Political ideology remains statistically not significant across specifications.
+However, environmental policy stringency shows a small but statistically significant positive effect in all models where it is included.
+This suggests that the cyclical component of emissions may co-move with the timing or enforcement of regulatory frameworks.
+These results are in line with recent findings for European regions, which report that emissions growth rates are highly volatile and sensitive to policy phases, particularly within the EU ETS framework (Dueñas and Mandel, 2025).
+4.3 Instrumental variables estimations
+To address potential endogeneity concerns related to low-carbon investments, we complement our baseline fixed-effects specification with an instrumental variables (2SLS) approach.
+While the use of lagged explanatory variables in the baseline models helps mitigate simultaneity bias, endogeneity may persist due to omitted variables or reverse causality.
+For instance, structurally more proactive regions might both attract higher low-carbon funding and implement complementary environmental measures not fully captured by our controls, potentially biasing the estimated coefficients.
+We instrument per capita low-carbon expenditures from both programming periods (2007–2013 and 2014–2020) using regional structural indicators from the ARDECO database: (i) labour productivity per employee (log), (ii) consumption of fixed capital (log), (iii) hours worked per capita, and (iv) compensation of employees (log).
+These instruments are lagged by three years, not only to ensure temporal separation from emissions outcomes, but also based on the assumption that the allocation of funds is determined by pre-existing structural conditions.
+Additionally, we include regional development level dummies (TER and LDR, with DER as the reference category), which reflect the eligibility and allocation logic of the ESI funds.
+These variables are plausibly exogenous to short- and medium-term emissions dynamics.
+First-stage results, presented in Table 4, confirm the strength of our instruments, with F-statistics exceeding conventional thresholds and statistically significant coefficients in the expected direction.
+| Low Carbon Exp. pc07-13 | Low Carbon Exp. pc14-20 | |
+| Labour productivity ppet-3 | -1.033*** | -0.683*** |
+| (0.114) | (0.071) | |
+| Hours worked per capitat-3 | 0.003*** | 0.001*** |
+| (0.000) | (0.000) | |
+| Compensation of Employeest-3 | -1.836*** | 0.422*** |
+| (0.237) | (0.109) | |
+| Consumption of fixed capitalt-3 | 1.990*** | -0.585*** |
+| (0.237) | (0.110) | |
+| TER | 0.970*** | 0.559*** |
+| (0.117) | (0.066) | |
+| LDR | 1.527*** | 1.329*** |
+| (0.145) | (0.080) | |
+| Time Dummies | Yes | Yes |
+| F-statistic | 114.670 | 265.060 |
+| Prob F | 0.000 | 0.000 |
+| Observations | 1,694 | 1,687 |
+| Notes: Standard errors in parentheses. Significance levels: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. TER and LDR refer to transition and less developed regions, respectively; DER is the reference group. | ||
+For the second stage, we focus on the specifications that include the most comprehensive set of controls, those considering GVA disaggregated by sector.
+The column structure is intentionally aligned with the baseline results to facilitate direct comparison.
+Table 5 presents the second-stage results for both emissions performance indicators.
+Overall, the IV estimates reveal patterns that are qualitatively consistent with those obtained from the fixed-effects specification.
+To further support the validity of our instruments, we regress the second-stage residuals on the full set of instruments; the absence of significant coefficients suggests that the instruments are uncorrelated with the structural error term.
+Results of this test are reported in Appendix Table A4.
+| Dependent variable: | ||||
+| Trend | Cycle | |||
+| (3t) | (6t) | (3c) | (6c) | |
+| GHGEpc | GHGEi | GHGEpc | GHGEi | |
+| Low Carbon Exp. pc | -0.587 | -0.224 | -0.601* | -0.144 |
+| (0.393) | (0.389) | (0.345) | (0.201) | |
+| Low Carbon Exp. pc | 5.692 | 4.207 | 0.460 | 0.340 |
+| (3.468) | (3.847) | (0.815) | (0.846) | |
+| Low Carbon Exp. pcTER | 1.399* | 1.949*** | -0.300 | 0.092 |
+| (0.732) | (0.695) | (0.484) | (0.399) | |
+| Low Carbon Exp. pcTER | -0.779 | 0.392 | -0.551 | -0.639 |
+| (3.041) | (3.005) | (0.665) | (0.702) | |
+| Low Carbon Exp. pcLDR | 1.126*** | 0.539 | 0.357 | 0.116 |
+| (0.408) | (0.387) | (0.339) | (0.204) | |
+| Low Carbon Exp. pcLDR | -2.350 | -2.487 | -0.556 | -0.360 |
+| (3.189) | (3.462) | (0.740) | (0.772) | |
+| EPSt-5 | 0.001 | -0.001 | 0.009*** | 0.011*** |
+| (0.007) | (0.007) | (0.003) | (0.003) | |
+| Political Ideologyt-5 | 0.006 | -0.010 | 0.003 | 0.002 |
+| (0.010) | (0.013) | (0.003) | (0.003) | |
+| GVAA,t-5 | 0.076** | 0.044 | -0.012 | -0.011 |
+| (0.037) | (0.038) | (0.009) | (0.007) | |
+| GVAB-E,t-5 | 0.213*** | -0.064 | -0.008 | -0.003 |
+| (0.057) | (0.078) | (0.018) | (0.018) | |
+| GVAF,t-5 | -0.110** | -0.134** | -0.046*** | 0.015 |
+| (0.048) | (0.058) | (0.015) | (0.017) | |
+| GVAG-J,t-5 | 0.205* | -0.129 | 0.018 | 0.078** |
+| (0.113) | (0.129) | (0.035) | (0.032) | |
+| GVAK-N,t-5 | -0.389*** | -0.436*** | -0.027 | 0.030 |
+| (0.079) | (0.094) | (0.039) | (0.038) | |
+| GVAO-U,t-5 | 0.147 | 0.303** | -0.180*** | -0.045 |
+| (0.102) | (0.125) | (0.042) | (0.041) | |
+| Observations | 3,408 | 3,408 | 3,408 | 3,408 |
+| NUTS 2 fixed effects | Yes | Yes | Yes | Yes |
+| Time fixed effects | Yes | Yes | Yes | Yes |
+| Notes: Robust standard errors clustered at the NUTS 2 level are reported in parentheses.
+Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01.
+Low Carbon Exp.
+pc07-13 and Low Carbon Exp.
+pc14-20 are instrumented.
+| ||||
+For emissions per capita trends (Column 3t), low-carbon expenditures from the 2007–2013 programme are not statistically significant in developed regions, but the interaction term for less-developed regions remains positive and significant.
+This confirms that these regions experienced increases in long-term emissions per capita following higher investments.
+A key difference from the baseline is that the TER interaction becomes positive and significant, whereas it was previously negative and not significant.
+This suggests potential divergence in the response of transition regions under the IV approach.
+The 2014–2020 effects remain statistically not significant across development levels.
+Regarding emissions intensity trends (Column 6t), we find a significant positive effect for the 2007-2013 expenditures in transition regions.
+This result departs from the baseline findings, where all interaction effects were not statistically significant.
+Turning to the cyclical components, for emissions per capita (Column 3c), the 2007–2013 programme has a negative and statistically significant effect in developed regions, consistent with a smoothing effect on short-term volatility.
+However, the previously significant negative interaction effects for LDR are no longer significant under the IV specification.
+Similarly, for emissions intensity cycles (Column 6c), none of the coefficients related to expenditures are statistically significant, contrasting with the baseline results where some mitigating effects were observed.
+In summary, the IV results reinforce the main conclusion of our baseline specification: low-carbon investments have different effects on emissions across different development levels.
+While some coefficients change in significance or sign, the overarching pattern, particularly the long-term positive association in less-developed regions, remains robust.
+These findings enhance confidence in the causal interpretation of the estimated effects and underscore the importance of considering regional structural differences when assessing climate policy outcomes.
+4.4 Effect by type of gases
+We break down emissions by specific gases, considering CO2, CH4, and N2O, to assess the differentiated impact of low-carbon investments.
+We adopt the same methodological approach used for total GHG emissions: decomposing the per capita and intensity time series for each gas into trend and cycle components, and estimating the effects of expenditures on these indicators following the specification in Eq.(3).
+Results are reported in Tables 6 and 7, corresponding to the trend and cycle components, respectively.
+For readability, the estimated coefficients of control variables are omitted from the tables.
+Regarding the trend of CO2 emissions (upper panel of Table 6), we find no statistically significant effects of the 2007–2013 programme for the reference group (developed regions) in terms of per capita emissions.
+However, the interaction terms with less developed regions are consistently positive and significant across specifications.
+For CO2 intensity, a significant negative effect is observed in column 4 for the reference group, indicating possible efficiency gains, yet these are counteracted by a positive and significant interaction with LDR in the same column.
+For the 2014–2020 programme, effects on CO2 per capita and intensity are generally not significant.
+Nonetheless, the interaction with LDR remains positive and significant in columns 1 and 2, reinforcing the earlier programming effect pattern.
+| Dependent variables by gas type | ||||||
+| CO2 per capita | CO2 intensity | |||||
+| Model | (1) | (2) | (3) | (4) | (5) | (6) |
+| Low Carbon Exp. pc | 0.066 | -0.061 | 0.077 | -0.356*** | -0.110 | -0.025 |
 | (0.199) | (0.154) | (0.102) | (0.125) | (0.159) | (0.123) | |
-| Low Carbon Exp. pc | 0.143 | 0.291 | 0.237 | 0.133 | 0.273 | 0.159 |
+| Low Carbon Exp. pc | -0.300 | -0.424 | 0.172 | -1.235 | -0.740 | -0.307 |
 | (0.523) | (0.708) | (0.707) | (0.802) | (0.643) | (0.658) | |
-| Low Carbon Exp. pcTER | 0.120 | 0.279∗∗∗ | 0.323∗∗∗ | 0.113 | 0.282∗∗∗ | 0.326∗∗∗ |
+| Low Carbon Exp. pcTER | -1.164 | -0.371 | -0.561 | -0.018 | -0.217 | -0.298 |
 | (0.715) | (0.868) | (0.958) | (0.691) | (0.591) | (0.660) | |
-| Low Carbon Exp. pcTER | 0.247 | 1.620∗∗ | 1.807∗∗∗ | 0.266 | 1.746∗∗∗ | 1.558∗∗ |
+| Low Carbon Exp. pcTER | -0.726 | 0.094 | -1.216 | 0.900 | -0.105 | -0.850 |
 | (1.728) | (1.663) | (1.675) | (1.559) | (1.592) | (1.531) | |
-| Low Carbon Exp. pcLDR | 0.010 | 0.003 | 0.012 | 0.018 | 0.018 | 0.016 |
+| Low Carbon Exp. pcLDR | 0.621** | 0.580** | 0.420** | 0.468* | 0.415 | 0.308 |
 | (0.258) | (0.224) | (0.202) | (0.238) | (0.282) | (0.268) | |
-| Low Carbon Exp. pcLDR | 0.029 | 0.473 | 0.481 | 0.035 | 0.467 | 0.253 |
+| Low Carbon Exp. pcLDR | 1.430*** | 1.189* | 0.643 | 0.783 | 0.813 | 0.411 |
 | (0.538) | (0.667) | (0.675) | (0.778) | (0.616) | (0.633) | |
-| GDP | 0.458∗∗∗ | 0.519∗∗∗ | 0.538∗∗∗ | 0.463∗∗∗ | ||
-| (0.028) | (0.035) | (0.028) | (0.035) | |||
-| EPS | 0.025∗∗∗ | 0.030∗∗∗ | 0.017∗∗∗ | 0.016∗∗∗ | ||
-| (0.004) | (0.004) | (0.004) | (0.004) | |||
-| Political Ideology | 0.003 | 0.002 | 0.001 | 0.0001 | ||
-| (0.003) | (0.003) | (0.003) | (0.003) | |||
-| GVAA | 0.011 | 0.006 | ||||
-| (0.008) | (0.008) | |||||
-| GVAB-E | 0.112∗∗∗ | 0.082∗∗∗ | ||||
-| (0.016) | (0.016) | |||||
-| GVAF | 0.086∗∗∗ | 0.030∗∗ | ||||
-| (0.013) | (0.013) | |||||
-| GVAG-J | 0.058∗∗ | 0.199∗∗∗ | ||||
-| (0.025) | (0.025) | |||||
-| GVAK-N | 0.069∗∗ | 0.143∗∗∗ | ||||
-| (0.033) | (0.033) | |||||
-| GVAO-U | 0.187∗∗∗ | 0.0003 | ||||
-| (0.037) | (0.037) | |||||
-| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
-| R2 | 0.248 | 0.295 | 0.302 | 0.166 | 0.150 | 0.136 |
-| Adjusted R2 | 0.193 | 0.235 | 0.241 | 0.105 | 0.078 | 0.060 |
-| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
-For transition regions, the estimated effect is relatively clearer; nonetheless, determining the net effect related to each program has opposite signs.
-The interaction with expenditures in the 2007-2013 programme shows a positive and significant effect (except in models 1 and 4) even after discounting the effect of the reference group.
-The net effect of expenditure in this period suggests an increase in the intensity of emission cycles, both per capita and concerning GDP.
-In contrast, the effect associated with expenditure in the 2014-2020 programme is consistently negative and significant (except in models 1 and 4).
-Then, considering the effect of the reference group, the net effect remains negative and comparatively much stronger.
-Although the estimated effects for both periods are opposing, the expected effect of these expenditures is to reduce the intensity of emission cycles.
-For less developed regions, the effect of expenditures associated with emission cycles is negative in almost all specifications for both periods, although not significant.
-Thus, considering the estimated effects in the reference group, the effect of expenditures on emission cycles in less developed regions is not significant.
-Regarding the other control variables, the GDP cycle has a strong positive impact on the per capita emission cycle and a negative impact on the emission intensity cycle, indicating that economic growth cycles are associated with higher per capita emissions but with a lower intensity of emissions intensity.
-The country-level EPS dummy shows a significant negative effect on emissions, both per capita and intensity, suggesting that stricter environmental policies are associated with lower emissions in their cycles.
-The region’s weighted average political ideology does not show a significant impact on emission cycles, indicating that local political outcomes, whether left or right, may not have a direct relationship with the cyclical variability of emissions.
-Finally, the sectoral gross value added coefficients indicate that different sectors have varying impacts on emission cycles.
-For example, the manufacturing sector (B-E) has a significant positive impact on per capita emissions and a negative impact on emission intensity.
-Therefore, most industry sectors have a significant influence on the cyclical variations in emissions.
-4.3 Effect by type of gases
-We break down emissions by specific gases, considering CO2, CH4, and N2O, to assess low-carbon investments’ impacts, using the same methodology.
-Detailed numerical result tables can be found in the Appendix (Tables A4 to A9), allowing us to concentrate on qualitative insights.
-Disaggregated GHG emissions data at the regional level allow for a better understanding of the heterogeneous impacts of the investments and can help identify specific expenditure gaps to improve policy responses.
-While CO2 contributes significantly to GHG emissions and changes in CO2 emissions primarily explain the impact of low-carbon funding, it is essential to consider a broader range of gases, which have different sources and economic activity patterns.
-In addition, decomposing GHG emissions into leading pollutant gases is relevant as European regions vary significantly in emission types, which are linked to their socioeconomic composition (Crippa et al., 2023).
-CO2 results from burning fossil fuels, chemical processes like cement production, and natural emissions from trees and biological materials.
-CH4 is emitted during fossil fuel extraction and transportation and significantly from agricultural activities such as livestock and organic waste decomposition in landfills.
-N2O is produced by agricultural and industrial activities, fossil fuel combustion, and wastewater treatment.
-The fixed effects estimations results on CO2, CH4, and N2O emissions are presented on Tables A4 to A6 in the Appendix.
-For CO2 emissions, low-carbon expenditures for the 2007-2013 and 2014-2020 programmes show significant negative impacts on emissions per capita and relative to GDP in developed and transition regions, indicating effective reductions.
-However, in less developed regions, these expenditures result in positive significant impacts, suggesting an increase in emissions or less effective reductions.
-These results are similar to those observed for overall GHG emissions (Table 2).
-For CH4 emissions, low-carbon expenditures for the 2007-2013 programme generally show negative but non-significant impacts across most specifications (including emissions per capita and relative to GDP).
-For the 2014-2020 programme, there are significant negative impacts in transition regions but positive or non-significant impacts in less developed regions, indicating inconsistent effectiveness.
-Regarding N2O emissions, low-carbon expenditures for both the 2007-2013 and 2014-2020 programmes have negative impacts but are not significant in developed regions.
-In transition regions, the impacts are negative and significant, while in less developed regions, the impacts are positive and significant, indicating less effective reductions or increases in emissions.
-These results show that low-carbon expenditures have varying impacts depending on the type of gas and the level of regional development.
-In developed and transition regions, these investments are generally effective in reducing CO2 emissions but exhibit mixed and often non-significant results for CH4 and N2O.
-In less developed regions, low-carbon expenditures frequently result in increased emissions or less effective reductions across all gases.
-Tables A7 to A9 in the Appendix present the fixed effect estimation results on the cyclical components of CO2, CH4, and N2O emissions, respectively.
-For CO2 emissions, low-carbon expenditures for the 2007-2013 and 2014-2020 programmes show generally non-significant impacts on the cyclical component of emissions per capita and relative to GDP.
-There are mixed results in transition regions, with significant positive impacts for 2007-2013 expenditures and significant negative impacts for 2014-2020 expenditures.
-In less developed regions, the impacts are primarily non-significant, with some negative impacts on the cyclical emission component for 2014-2020 expenditures in certain specifications.
-For CH4 emissions, low-carbon expenditures for the 2007-2013 programme show consistently significant negative impacts for all regions, indicating effective reductions on the cyclical emissions component.
-The 2014-2020 expenditures yield mixed results, with generally non-significant impacts except for significant negative impacts in transition regions.
-Regarding N2O emissions, low-carbon expenditures for both the 2007-2013 and 2014-2020 programmes typically show non-significant impacts.
-In transition regions, there are some negative differential impacts for 2014-2020 expenditures, but these are not consistently significant.
-In less developed regions, the differential impacts are mixed, with positive and significant for the 2007-2013 period and non-significant for 2014-2020.
-The estimations for the cyclical emissions components reveal that low-carbon expenditures have mixed results, which vary depending on the type of gas used and the level of regional development.
-This is observed for both per capita and per unit of GDP.
-5 Concluding remarks
-This paper examines the impact of climate action through one of the key vehicles for European regions: the “Low-Carbon Economy” funds of the EU Cohesion and Structural Policy for the 2007-2013 and 2014-2020 programmes.
-Our analysis reveals that the impact of spending on low-carbon economies in reducing GHG emissions varies significantly depending on the level of regional development, according to the EU classification that distinguishes between less developed regions, regions in transition, and developed regions.
-We observe that, for both analysed programmes, investments in low-carbon economies are associated with a statistically significant decrease in the long-term trend of total emissions, both per capita and relative to GDP, in developed and transition regions.
-However, in less developed regions, these investments have a statistically significant positive impact, indicating either an increase in emissions or a less effective reduction.
-Furthermore, estimations based on specific gas types also demonstrate differential impacts.
-These results indicate a clear dichotomy: investments yield significant reductions in long-term emissions trends in developed and transition regions but appear to have the opposite effect in less developed regions.
-This aligns with insights from Doda (2014) and Sheldon (2017), who emphasise the variability of emission responses to economic cycles, particularly noting the stabilising effect of wealthier economies on emission patterns.
-Conversely, in less developed regions, the counterintuitive increase in emissions suggests a subtle interplay between economic development and environmental sustainability efforts, reminiscent of the disparities identified in the existing literature (Naqvi and Zwickl, 2017; Ivanova et al., 2017).
-In addition, our analysis showed that European regions’ responses to low-carbon economies funds vary significantly among programmes, emphasising the importance of taking a nuanced approach when evaluating the impact of these investments on environmental sustainability.
-This involves considering the particularities of each program, the intrinsic differences between regions and types of emissions.
-Our findings support the idea of a refined approach to the implementation of climate policy objectives through the EU cohesion policy.
-We agree with other studies claiming that a diverse set of guidelines must be implemented to effectively reduce emissions rather than relying on individual initiatives (Brinkley, 2014).
-Consequently, economy-wide assessments, sector-specific trends, planning expertise in implementing place-based approaches, and extensive case studies can help underscore the complexity of decoupling economic growth from pollutant emissions (Brinkley, 2014; Naqvi and Zwickl, 2017).
-In general, the need for differentiated policies to achieve economic development is an issue that has been widely studied since the creation of the Cohesion Policy for European regions, underscoring the importance of addressing persistent inequalities between EU regions.
-Very early on, Boldrin and Canova (2001) highlighted that these policies tend to be more redistributive than growth-oriented, reflecting political compromises over economic efficiency.
-However, investments in education and human capital, which constitute a significant portion of these funds, have shown medium-term positive impacts, emphasising the critical role of human capital development in fostering long-term regional growth and convergence (Rodríguez-Pose and Fratesi, 2004).
-Moreover, Rodríguez-Pose and Garcilazo (2015) and Charron et al.
-(2014) argue that the quality of local and regional governance directly affects economic growth and the effectiveness of structural funds.
-They find significant variations in governance quality within countries, such as levels of corruption and efficiency of public services, which have substantial implications for policy design.
-Iammarino et al.
+| CH4 per capita | CH4 intensity | |||||
+| Low Carbon Exp. pc | -0.115 | -0.005 | -0.045 | -0.537*** | -0.054 | -0.147 |
+| (0.181) | (0.182) | (0.175) | (0.168) | (0.215) | (0.210) | |
+| Low Carbon Exp. pc | 0.826 | 0.042 | -0.095 | -0.109 | -0.275 | -0.573 |
+| (0.909) | (0.801) | (0.685) | (1.084) | (0.771) | (0.699) | |
+| Low Carbon Exp. pcTER | -0.718** | -0.507 | -0.588 | 0.429 | -0.354 | -0.324 |
+| (0.333) | (0.426) | (0.468) | (0.367) | (0.280) | (0.290) | |
+| Low Carbon Exp. pcTER | -2.247** | -1.191 | -1.261 | -0.621 | -1.390 | -0.894 |
+| (1.136) | (1.103) | (1.103) | (1.215) | (1.108) | (1.087) | |
+| Low Carbon Exp. pcLDR | 0.515** | 0.504** | 0.523** | 0.364 | 0.339 | 0.410 |
+| (0.201) | (0.199) | (0.211) | (0.238) | (0.271) | (0.275) | |
+| Low Carbon Exp. pcLDR | -0.435 | 0.148 | 0.119 | -1.083 | -0.227 | -0.112 |
+| (0.887) | (0.777) | (0.679) | (1.040) | (0.764) | (0.700) | |
+| N2O per capita | N2O intensity | |||||
+| Low Carbon Exp. pc | -0.115 | -0.059 | -0.058 | -0.543*** | -0.092 | -0.163 |
+| (0.100) | (0.091) | (0.070) | (0.133) | (0.131) | (0.108) | |
+| Low Carbon Exp. pc | 0.180 | 0.142 | 0.288 | -0.774 | -0.118 | -0.196 |
+| (0.340) | (0.382) | (0.339) | (0.631) | (0.312) | (0.299) | |
+| Low Carbon Exp. pcTER | -0.574*** | -0.187 | -0.140 | 0.570 | -0.088 | 0.134 |
+| (0.206) | (0.166) | (0.160) | (0.583) | (0.364) | (0.361) | |
+| Low Carbon Exp. pcTER | -0.969*** | -0.764* | -0.829 | 0.647 | -1.049** | -0.442 |
+| (0.369) | (0.440) | (0.527) | (0.756) | (0.411) | (0.659) | |
+| Low Carbon Exp. pcLDR | 0.536*** | 0.347** | 0.374** | 0.380** | 0.172 | 0.267 |
+| (0.178) | (0.144) | (0.150) | (0.193) | (0.191) | (0.176) | |
+| Low Carbon Exp. pcLDR | 1.358*** | 0.491 | 0.315 | 0.726 | 0.089 | 0.089 |
+| (0.424) | (0.377) | (0.355) | (0.619) | (0.309) | (0.305) | |
+| Notes: Estimated coefficients of control variables are omitted.
+Model specifications correspond to those in Table 2.
+Robust standard errors clustered at the NUTS 2 level are reported in parentheses.
+Significance levels: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01.
+| ||||||
+Turning to CH4 emissions (middle panel of Table 6), the 2007–2013 programme does not significantly affect per capita emissions in the control group.
+There is a negative and significant interaction with transition regions in column 1.
+In contrast, interactions with LDR are consistently positive and significant.
+For CH4 intensity, a significant negative impact is observed for the reference group in column 4, but not for the interactions with TER or LDR.
+Under the 2014–2020 programme, we find no statistically significant effects on either CH4 per capita or intensity, across all regional groups.
+Finally, in the case of N2O emissions (lower panel of Table 6), the 2007–2013 programme shows a significant reduction in per capita emissions for TER in column 1, while interaction terms for LDR are again positive and significant throughout.
+Regarding N2O intensity, a negative and significant coefficient is observed for the reference group in column 4, but positive and significant interactions with LDR offset this.
+For the 2014–2020 programme, the interaction with TER is negative and statistically significant in both column 1 and 2 (N2O per capita) and column 5 (N2O intensity).
+The interaction with LDR remains positive and significant only in column 1, in line with the pattern of increased emissions observed in the earlier programme.
+As regards the cycle component of CO2 emissions (upper panel of Table 7), the 2007–2013 programme is associated with a statistically significant reduction in per capita emissions for the reference group (developed regions) in models 2 and 3.
+Additionally, the interaction terms with less developed regions are consistently negative and statistically significant across all specifications.
+These suggest that low-carbon investments contributed to smoothing short-term fluctuations in CO2 emissions in these areas.
+The interaction terms with transition regions, however, are not statistically significant.
+For CO2 intensity, no significant effects are observed for the reference group.
+However, the interaction terms with LDR are negative and statistically significant in models 5 and 6.
+For the 2014–2020 programming period, none of the estimated effects are statistically significant for either the reference group or interaction terms.
+| Dependent variables by gas type | ||||||
+| CO2 per capita | CO2 intensity | |||||
+| Model | (1) | (2) | (3) | (4) | (5) | (6) |
+| Low Carbon Exp. pc | -0.010 | -0.066* | -0.081** | 0.034 | -0.007 | -0.018 |
+| (0.029) | (0.037) | (0.034) | (0.047) | (0.038) | (0.035) | |
+| Low Carbon Exp. pc | -0.174 | -0.224 | -0.237 | -0.137 | -0.203 | -0.217 |
+| (0.273) | (0.293) | (0.258) | (0.291) | (0.291) | (0.273) | |
+| Low Carbon Exp. pcTER | -0.206 | -0.163 | -0.166 | -0.148 | 0.015 | 0.008 |
+| (0.165) | (0.145) | (0.143) | (0.197) | (0.113) | (0.118) | |
+| Low Carbon Exp. pcTER | 0.325 | 0.090 | 0.127 | -0.113 | -0.478 | -0.459 |
+| (0.880) | (0.865) | (0.823) | (0.948) | (0.844) | (0.841) | |
+| Low Carbon Exp. pcLDR | -0.152** | -0.169*** | -0.199*** | -0.130* | -0.155*** | -0.173*** |
+| (0.065) | (0.062) | (0.061) | (0.068) | (0.054) | (0.051) | |
+| Low Carbon Exp. pcLDR | 0.295 | 0.350 | 0.410 | 0.181 | 0.278 | 0.316 |
+| (0.285) | (0.302) | (0.274) | (0.298) | (0.291) | (0.275) | |
+| CH4 per capita | CH4 intensity | |||||
+| Low Carbon Exp. pc | -0.085*** | -0.073*** | -0.062*** | -0.041 | -0.014 | 0.000 |
+| (0.025) | (0.023) | (0.022) | (0.027) | (0.028) | (0.031) | |
+| Low Carbon Exp. pc | 0.252** | 0.148 | 0.145 | 0.289*** | 0.169* | 0.165* |
+| (0.112) | (0.103) | (0.111) | (0.106) | (0.092) | (0.089) | |
+| Low Carbon Exp. pcTER | -0.109 | -0.062 | -0.035 | -0.050 | 0.117** | 0.140*** |
+| (0.094) | (0.065) | (0.058) | (0.127) | (0.047) | (0.053) | |
+| Low Carbon Exp. pcTER | -0.527 | -0.614* | -0.667** | -0.965** | -1.182*** | -1.252*** |
+| (0.354) | (0.337) | (0.334) | (0.411) | (0.304) | (0.337) | |
+| Low Carbon Exp. pcLDR | 0.019 | 0.000 | 0.010 | 0.040 | 0.014 | 0.036 |
+| (0.051) | (0.055) | (0.055) | (0.054) | (0.058) | (0.058) | |
+| Low Carbon Exp. pcLDR | -0.256 | 0.037 | -0.002 | -0.369** | -0.035 | -0.097 |
+| (0.172) | (0.173) | (0.184) | (0.170) | (0.178) | (0.177) | |
+| N2O per capita | N2O intensity | |||||
+| Low Carbon Exp. pc | -0.009 | -0.022 | -0.028 | 0.036 | 0.038 | 0.035 |
+| (0.021) | (0.029) | (0.031) | (0.026) | (0.025) | (0.028) | |
+| Low Carbon Exp. pc | 0.093 | 0.066 | 0.048 | 0.137 | 0.093 | 0.074 |
+| (0.150) | (0.149) | (0.127) | (0.166) | (0.146) | (0.142) | |
+| Low Carbon Exp. pcTER | -0.098** | -0.159*** | -0.163*** | -0.040 | 0.020 | 0.011 |
+| (0.044) | (0.056) | (0.057) | (0.069) | (0.039) | (0.046) | |
+| Low Carbon Exp. pcTER | 0.050 | 0.015 | 0.017 | -0.381 | -0.549*** | -0.563*** |
+| (0.259) | (0.268) | (0.253) | (0.245) | (0.198) | (0.212) | |
+| Low Carbon Exp. pcLDR | 0.113* | 0.129* | 0.121* | 0.134** | 0.143* | 0.147** |
+| (0.063) | (0.073) | (0.072) | (0.064) | (0.073) | (0.073) | |
+| Low Carbon Exp. pcLDR | -0.167 | -0.083 | -0.068 | -0.284* | -0.159 | -0.166 |
+| (0.153) | (0.162) | (0.148) | (0.165) | (0.158) | (0.156) | |
+| Notes: Estimated coefficients of control variables are omitted.
+Model specifications correspond to those in Table 2.
+Robust standard errors clustered at the NUTS 2 level are reported in parentheses.
+Significance levels: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01.
+| ||||||
+Turning to the cycle component of CH4 emissions (middle panel of Table 7), the 2007–2013 programme is associated with consistent and statistically significant reductions in per capita emissions for the reference group across all model specifications.
+In contrast, interaction terms with both TER and LDR are not statistically significant.
+In the case of the 2014–2020 programme, we observe a positive and statistically significant effect for the reference group in model 1 (CH4 per capita), and across all models for CH4 intensity (columns 4–6).
+The interaction terms with TER are negative and statistically significant in all models, except model 1.
+Interaction terms with LDR are not significant in most models, except CH4 intensity in column 4, where a negative effect is detected.
+Regarding the cycle component of N2O emissions (lower panel of Table 7), the 2007–2013 programme shows no significant effects for the reference group in terms of per capita emissions.
+However, all estimated interaction terms with TER are negative and statistically significant, while those with LDR are positive and significant (columns 1-3).
+For N2O intensity, the reference group shows no statistically significant response.
+However, the interaction terms with LDR are positive and statistically significant in models 4–6.
+For the 2014–2020 programming period, interaction terms with TER are again negative and statistically significant in models 5 and 6, and for LDR in model 1.
+5 Discussion
+The results from the baseline fixed effects model reveal heterogeneous impacts of low-carbon investments on both emission trends and cycles.
+Generally, these investments do not yield statistically significant outcomes in more developed regions.
+However, in less developed areas, low-carbon investments appear paradoxically associated with increases in the long-term trend of per capita emissions, and, in some specifications, also in emissions intensity.
+This suggests that, in these regions, investments stimulate economic growth that remains dependent on carbon-intensive technologies.
+In contrast, when examining short-term fluctuations, the results suggest a more positive role for these investments in reducing emissions volatility, especially in less developed regions, indicating that the funds may help stabilise emissions in the short run.
+When we employ instrumental variables to address potential endogeneity, the findings further reinforce the existence of regional disparities.
+The positive effect on per capita emissions in less developed regions is clearly confirmed, and additional significant effects emerge for transition regions that were not visible in the initial analysis.
+While this might suggest that estimations for transition regions were underestimated, the overall pattern is strengthened, reflecting heterogeneous regional impacts.
+Hence, the IV analysis enhances the credibility of a causal interpretation of the differentiated expenditure effects.
+When disaggregating the results by GHG type (CO2, CH4, and N2O), a more nuanced picture emerges.
+For trends, the increase in per capita emissions observed in less developed regions is primarily driven by CO2, the gas most closely linked to industrial activity and energy production.
+In contrast, the results for CH4 and N2O show mixed patterns, with adverse effects in some less developed regions and mitigating effects in transition areas.
+Regarding emission cycles, low-carbon investments show more consistently positive effects, particularly in reducing volatility for CO2 in less developed regions and for CH4 and N2O in transition regions.
+However, these results should be interpreted in light of the characteristics of the emission sources and policy contexts associated with each gas.
+Unlike CO2, which is closely tied to fossil fuel combustion and therefore directly targeted by ESI-funded low-carbon investments, CH4 and N2O emissions are largely linked to agricultural practices and specific industrial processes (Scheffler and Wiegmann, 2024).
+Their responsiveness to low-carbon funding may thus be more limited or dependent on complementary policies.
+In particular, the Common Agricultural Policy (CAP), which is funded separately from the ESI funds, has distinct objectives that may create overlaps with climate-related goals (European Court of Auditors, 2021).
+More broadly, the low-carbon economy fund from the ESI funds is one among several instruments in the EU’s climate policy architecture.
+Notably, the EU ETS has played a more central role in recent decades (Bayer and Aklin, 2020).
+Nevertheless, ESI funding remains important, as it aims to address regional disparities and supports place-based transitions.
+Altogether, our findings provide strong empirical support for our central hypothesis.
+While we find consistent evidence that low-carbon economy funding does not lead to long-term reductions in emissions indicators in more developed regions, the adverse effects observed in less developed regions are particularly concerning, given that these areas are the primary targets of ESI fund support.
+Nevertheless, the results suggest that while the funds may not be highly effective in driving long-term decarbonisation, they could still play a stabilising role by dampening short-term emissions volatility.
+It is important to note that most statistically significant effects are concentrated in the 2007-2013 programming period.
+The lack of robust results for the 2014-2020 period may be due to data limitations.
+Given that the environmental impacts of investments often take time to materialise, the imprecision likely reflects the short post-investment observation window.
+Moreover, in models using emissions intensity as the dependent variable, results are more frequently statistically not significant.
+Although this did not invalidate the central hypothesis of the study, it raises policy concerns.
+Specifically, the limited effects on emissions intensity, even in more developed areas, clearly hint that funding is not effectively supporting the decoupling of emissions from economic growth.
+These findings call for a reassessment of EU climate investment strategies.
+On the one hand, the failure to curb long-term emissions in less developed areas raises questions about allocative efficiency and the risk of reinforcing carbon lock-in (Unruh, 2000).
+On the other hand, mechanisms behind the heterogeneous impacts – such as institutional capacity gaps, infrastructure rigidity, or lack of policy complementarities – remain underexplored and merit further attention (Rodríguez-Pose and Garcilazo, 2015; Charron et al., 2014).
+In this respect, Iammarino et al.
 (2019) emphasise the need for tailored policies that are sensitive to the structural prospects of different regions, aligning with the Barca report’s advocacy for place-based strategies (Barca, 2009).
-These insights align with the broader literature on the Environmental Kuznets Hypothesis, which suggests that economic development can lead to improved environmental outcomes, but only if accompanied by effective governance and tailored policy interventions.
-Our research underscores the pressing need for tailored environmental policies to optimise interventions across European regions.
-The differential impacts of cohesion policy investments on regional emission trends and cycles highlight the necessity of moving away from one-size-fits-all solutions towards development-sensitive and emission-specific strategies.
-Special attention should be given to less developed regions, which, although not major emitters per capita currently, risk becoming significant pollution producers as they develop economically.
-Aligning investment strategies with regional profiles can enhance the effectiveness of emissions reduction efforts, contributing to a more nuanced policy dialogue on sustainable development.
-This underscores the critical intersection of regional growth, sustainability investments, and environmental outcomes within the EU’s policy framework.
+In addition, while the observed short-term stabilisation of emissions is encouraging, it may reflect counter-cyclical spending effects rather than deeper structural transformations, thereby questioning the long-term decarbonisation potential of these funds.
+Finally, although beyond this paper’s scope, some of the emissions increases in less developed regions may reflect industrial relocation rather than policy failure.
+From a broader EU or global perspective, such shifts may be environmentally neutral.
+These trade-offs also underline the need for more targeted, region-specific strategies.
+Future research should investigate the mechanisms behind regional disparities and identify the complementary policies needed to align the ESI spending with long-term climate and development objectives.
+6 Conclusions
+This paper examines the climate impact of one of the European Union’s key instruments for regional development: the “Low-Carbon Economy” funds allocated through the European Structural and Investment funds during the 2007-2013 and 2014-2020 programming periods.
+Focusing on per capita emissions and emissions intensity – both decomposed into trend and cyclical components – we estimate the structural (long-run) and cyclical (short-run) effects of low-carbon expenditures across regions.
+Our findings point to a substantial degree of heterogeneity.
+In more developed and transition regions, low-carbon funding shows no consistent long-term effect in reducing emissions, whether measured per capita or relative to GDP.
+However, such investments are associated with significant increases in long-term emissions trends in less developed regions, despite being the primary beneficiaries of these funds.
+These results suggest that, in lagging regions, climate-related investments may stimulate economic activity without achieving the intended decoupling from carbon-intensive processes.
+Instrumental variable estimations reinforce the robustness of our findings and lend support to a causal interpretation of the observed relationships.
+When disaggregating emissions by gas type (CO2, CH4, and N2O), results show further heterogeneity.
+In less developed regions, increases in emissions are mainly associated with CO2, whereas the effects on methane and nitrous oxide are less robust and vary across regions.
+Regarding short-run dynamics, we find some evidence that low-carbon investments contributed to reducing emissions volatility, especially related to expenditures of the 2007-2013 programme.
+Overall, the empirical results raise questions about how effectively the EU Cohesion Policy supports the broader decarbonisation agenda.
+While our findings point to some limitations in the environmental impact of low-carbon investments, they also highlight the regional challenges in turning funding into measurable emission reductions.
+Structural conditions, institutional capacity, and how policies are implemented matter.
+However, other elements, such as the types of projects financed, their interaction with other policy initiatives, and the time it takes for their effects to materialise, likely play a role as well.
+Our findings suggest that improving the climate effectiveness of the Cohesion Policy will require a more substantial alignment with regional contexts and development trajectories.
+Given the heterogeneous impacts observed across regions, it is essential to move beyond uniform approaches.
+Less developed regions deserve particular attention.
+Although they are not currently the highest emitters on a per capita basis, they face the risk of becoming major sources of emissions as they develop.
+We claim that aligning investment strategies with regional profiles can enhance the effectiveness of emissions reduction efforts.
 Acknowledgements
@@
 Estimation results modelled expenditures
+The cohesion data platform presents estimations of the expenditure by NUTS 2, which are aggregated for the most significant types of structural funds only.
+Therefore, to address our need for yearly data, we have opted for two approaches: the one presented in the main text, which assumes that expenditures are distributed evenly over time (dividing the reported expenditures over the programming period), and the another one, presented here, which assumes they follow the same pattern as the total aggregated expenditure reported by the cohesion data platform.888Data available at: https://cohesiondata.ec.europa.eu/Other/Historic-EU-payments-regionalised-and-modelled/tc55-7ysv/about_data.
+Both strategies yielded qualitatively similar conclusions, as a robustness check the results of the second strategy are shown in Tables A2 and A3.
 | Dependent variable: | ||||||
@@
 | (1) | (2) | (3) | (4) | (5) | (6) | |
-| Low Carbon Exp. pc | 0.182∗ | 0.167∗ | 0.004 | 0.185∗∗ | 0.213∗∗∗ | 0.134∗ |
+| Low Carbon Exp. pc | 0.057 | -0.034 | 0.111 | -0.530*** | -0.141 | -0.070 |
 | (0.231) | (0.190) | (0.134) | (0.182) | (0.199) | (0.164) | |
-| Low Carbon Exp. pc | 1.239∗ | 5.703 | 2.736 | 1.786∗∗∗ | 9.205∗∗ | 3.117 |
+| Low Carbon Exp. pc | -0.282 | -0.622 | 0.603 | -2.513 | -1.596 | -0.850 |
 | (1.654) | (2.160) | (1.928) | (2.621) | (2.037) | (1.896) | |
-| Low Carbon Exp. pcTER | 0.051 | 0.151 | 0.178 | 0.198 | 0.107 | 0.046 |
+| Low Carbon Exp. pcTER | -1.500 | -0.665 | -0.955 | 0.064 | -0.571 | -0.646 |
 | (0.951) | (1.117) | (1.213) | (0.818) | (0.748) | (0.809) | |
-| Low Carbon Exp. pcTER | 5.945∗∗∗ | 13.126 | 34.828∗∗∗ | 7.612∗∗∗ | 23.200∗∗ | 29.144∗∗∗ |
+| Low Carbon Exp. pcTER | -3.332 | -1.695 | -4.412 | 1.612 | -1.917 | -3.485 |
 | (4.645) | (4.536) | (4.570) | (4.472) | (4.400) | (4.136) | |
-| Low Carbon Exp. pcLDR | 0.792∗∗∗ | 0.768∗∗∗ | 0.588∗∗∗ | 0.394∗∗∗ | 0.439∗∗∗ | 0.387∗∗∗ |
+| Low Carbon Exp. pcLDR | 0.997*** | 0.879*** | 0.701*** | 0.548* | 0.544* | 0.447 |
 | (0.299) | (0.268) | (0.247) | (0.298) | (0.328) | (0.312) | |
-| Low Carbon Exp. pcLDR | 6.004∗∗∗ | 36.630∗∗∗ | 29.030∗∗∗ | 3.151∗∗∗ | 18.037∗∗∗ | 13.218∗∗∗ |
+| Low Carbon Exp. pcLDR | 2.964* | 2.601 | 1.533 | 1.442 | 1.766 | 1.011 |
 | (1.673) | (1.989) | (1.816) | (2.480) | (1.894) | (1.795) | |
-| GDP | 0.246∗∗∗ | 0.321∗∗∗ | 0.615∗∗∗ | 0.510∗∗∗ | ||
+| GDPt-5 | 0.240*** | -0.467*** | ||||
+| (0.065) | (0.102) | |||||
+| EPSt-5 | 0.001 | 0.002 | 0.003 | 0.002 | ||
 | (0.007) | (0.006) | (0.007) | (0.007) | |||
-| EPS | 0.010∗∗ | 0.013∗∗∗ | 0.005 | 0.010∗∗ | ||
+| Political Ideologyt-5 | 0.016 | 0.008 | -0.010 | -0.008 | ||
 | (0.010) | (0.009) | (0.016) | (0.014) | |||
-| Political Ideology | 0.017∗∗∗ | 0.001 | 0.002 | 0.007∗∗ | ||
-| (0.004) | (0.004) | (0.003) | (0.003) | |||
-| GVAA | 0.045∗∗∗ | 0.072∗∗∗ | ||||
+| GVAA,t-5 | 0.067** | 0.038 | ||||
 | (0.033) | (0.036) | |||||
-| GVAB-E | 0.222∗∗∗ | 0.084∗∗∗ | ||||
+| GVAB-E,t-5 | 0.246*** | -0.038 | ||||
 | (0.047) | (0.075) | |||||
-| GVAF | 0.023 | 0.095∗∗∗ | ||||
+| GVAF,t-5 | -0.136*** | -0.155*** | ||||
 | (0.041) | (0.057) | |||||
-| GVAG-J | 0.0004 | 0.110∗∗∗ | ||||
+| GVAG-J,t-5 | 0.129 | -0.205 | ||||
 | (0.090) | (0.131) | |||||
-| GVAK-N | 0.311∗∗∗ | 0.498∗∗∗ | ||||
+| GVAK-N,t-5 | -0.366*** | -0.392*** | ||||
 | (0.084) | (0.099) | |||||
-| GVAO-U | 0.345∗∗∗ | 0.334∗∗∗ | ||||
+| GVAO-U,t-5 | 0.167 | 0.302** | ||||
 | (0.109) | (0.134) | |||||
-| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
-| Adjusted R2 | 0.510 | 0.614 | 0.653 | 0.787 | 0.796 | 0.817 |
+| Observations | 3,872 | 3,408 | 3,408 | 3,872 | 3,408 | 3,408 |
 | Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
@@
 | (1) | (2) | (3) | (4) | (5) | (6) | |
-| Low Carbon Exp. pc | 0.049 | 0.020 | 0.013 | 0.031 | 0.024 | 0.030 |
+| Low Carbon Exp. pc | 0.034 | -0.021 | -0.031 | 0.056* | 0.038 | 0.026 |
 | (0.024) | (0.031) | (0.031) | (0.033) | (0.025) | (0.026) | |
-| Low Carbon Exp. pc | 0.378 | 6.016∗ | 5.025 | 0.306 | 5.767∗ | 4.607 |
+| Low Carbon Exp. pc | -0.671 | -0.713 | -0.779 | -0.145 | -0.569 | -0.602 |
 | (0.722) | (0.804) | (0.722) | (0.740) | (0.644) | (0.615) | |
-| Low Carbon Exp. pcTER | 0.163 | 0.376∗∗∗ | 0.426∗∗∗ | 0.162 | 0.385∗∗∗ | 0.453∗∗∗ |
+| Low Carbon Exp. pcTER | -0.290** | -0.270*** | -0.281*** | -0.224 | -0.091 | -0.095 |
 | (0.117) | (0.081) | (0.083) | (0.152) | (0.074) | (0.081) | |
-| Low Carbon Exp. pcTER | 0.076 | 25.584∗∗∗ | 30.159∗∗∗ | 0.072 | 25.857∗∗∗ | 23.884∗∗ |
+| Low Carbon Exp. pcTER | 1.702 | 1.176 | 1.261 | -0.321 | -0.915 | -0.914 |
 | (1.997) | (1.990) | (1.900) | (2.445) | (2.246) | (2.276) | |
-| Low Carbon Exp. pcLDR | 0.028 | 0.001 | 0.009 | 0.044 | 0.025 | 0.013 |
+| Low Carbon Exp. pcLDR | -0.143** | -0.141** | -0.157*** | -0.127** | -0.130*** | -0.138*** |
 | (0.058) | (0.056) | (0.056) | (0.053) | (0.046) | (0.046) | |
-| Low Carbon Exp. pcLDR | 0.225 | 5.407 | 5.581 | 0.230 | 5.182 | 2.318 |
+| Low Carbon Exp. pcLDR | 0.805 | 0.985 | 1.082 | 0.323 | 0.793 | 0.844 |
 | (0.757) | (0.809) | (0.745) | (0.768) | (0.668) | (0.642) | |
-| GDP | 0.460∗∗∗ | 0.519∗∗∗ | 0.535∗∗∗ | 0.464∗∗∗ | ||
+| GDPt-5 | -0.251*** | 0.080** | ||||
+| (0.037) | (0.035) | |||||
+| EPSt-5 | 0.009*** | 0.010*** | 0.011*** | 0.012*** | ||
 | (0.003) | (0.003) | (0.003) | (0.003) | |||
-| EPS | 0.026∗∗∗ | 0.031∗∗∗ | 0.018∗∗∗ | 0.018∗∗∗ | ||
+| Political Ideologyt-5 | 0.004 | 0.002 | 0.002 | 0.002 | ||
 | (0.003) | (0.003) | (0.003) | (0.003) | |||
-| Political Ideology | 0.003 | 0.003 | 0.0004 | 0.0004 | ||
-| (0.003) | (0.003) | (0.003) | (0.003) | |||
-| GVAA | 0.010 | 0.007 | ||||
+| GVAA,t-5 | -0.011 | -0.011 | ||||
 | (0.009) | (0.007) | |||||
-| GVAB-E | 0.115∗∗∗ | 0.078∗∗∗ | ||||
+| GVAB-E,t-5 | -0.011 | -0.004 | ||||
 | (0.018) | (0.018) | |||||
-| GVAF | 0.086∗∗∗ | 0.030∗∗ | ||||
+| GVAF,t-5 | -0.044*** | 0.017 | ||||
 | (0.014) | (0.017) | |||||
-| GVAG-J | 0.054∗∗ | 0.203∗∗∗ | ||||
+| GVAG-J,t-5 | 0.028 | 0.078** | ||||
 | (0.033) | (0.032) | |||||
-| GVAK-N | 0.076∗∗ | 0.134∗∗∗ | ||||
+| GVAK-N,t-5 | -0.049 | 0.024 | ||||
 | (0.040) | (0.038) | |||||
-| GVAO-U | 0.176∗∗∗ | 0.013 | ||||
+| GVAO-U,t-5 | -0.197*** | -0.057 | ||||
 | (0.042) | (0.042) | |||||
-| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
-| Adjusted R2 | 0.191 | 0.235 | 0.241 | 0.103 | 0.078 | 0.061 |
+| Observations | 3,872 | 3,408 | 3,408 | 3,872 | 3,408 | 3,408 |
 | Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
-Regression by type of polluting gas: trend components
-Trend components
-| Dependent variable: | ||||||
-| CO2pc | CO2i | |||||
-| (1) | (2) | (3) | (4) | (5) | (6) | |
-| Low Carbon Exp. pc | 0.170∗∗ | 0.161∗∗ | 0.002 | 0.172∗∗ | 0.193∗∗∗ | 0.094 |
-| (0.078) | (0.066) | (0.062) | (0.073) | (0.062) | (0.058) | |
-| Low Carbon Exp. pc | 0.678∗∗∗ | 0.833∗ | 0.234 | 0.926∗∗∗ | 1.337∗∗∗ | 0.617 |
-| (0.255) | (0.481) | (0.453) | (0.239) | (0.456) | (0.427) | |
-| Low Carbon Exp. pcTER | 0.003 | 0.268∗∗ | 0.013 | 0.084 | 0.097 | 0.110 |
-| (0.149) | (0.134) | (0.126) | (0.140) | (0.127) | (0.118) | |
-| Low Carbon Exp. pcTER | 1.718∗∗∗ | 0.049 | 2.866∗∗∗ | 2.102∗∗∗ | 0.847 | 2.222∗∗∗ |
-| (0.501) | (0.947) | (0.894) | (0.470) | (0.899) | (0.842) | |
-| Low Carbon Exp. pcLDR | 0.739∗∗∗ | 0.647∗∗∗ | 0.490∗∗∗ | 0.370∗∗∗ | 0.383∗∗∗ | 0.315∗∗∗ |
-| (0.096) | (0.083) | (0.078) | (0.090) | (0.078) | (0.074) | |
-| Low Carbon Exp. pcLDR | 2.603∗∗∗ | 3.441∗∗∗ | 2.423∗∗∗ | 1.514∗∗∗ | 2.267∗∗∗ | 1.621∗∗∗ |
-| (0.255) | (0.494) | (0.466) | (0.239) | (0.469) | (0.439) | |
-| GDP | 0.225∗∗∗ | 0.375∗∗∗ | 0.615∗∗∗ | 0.446∗∗∗ | ||
-| (0.019) | (0.022) | (0.018) | (0.021) | |||
-| EPS | 0.008 | 0.010∗∗ | 0.004 | 0.009∗ | ||
-| (0.005) | (0.005) | (0.005) | (0.005) | |||
-| Political Ideology | 0.022∗∗∗ | 0.002 | 0.008∗ | 0.006 | ||
-| (0.004) | (0.004) | (0.004) | (0.004) | |||
-| GVAA | 0.065∗∗∗ | 0.095∗∗∗ | ||||
-| (0.014) | (0.014) | |||||
-| GVAB-E | 0.244∗∗∗ | 0.055∗∗∗ | ||||
-| (0.017) | (0.016) | |||||
-| GVAF | 0.005 | 0.074∗∗∗ | ||||
-| (0.016) | (0.015) | |||||
-| GVAG-J | 0.032 | 0.143∗∗∗ | ||||
-| (0.033) | (0.031) | |||||
-| GVAK-N | 0.406∗∗∗ | 0.585∗∗∗ | ||||
-| (0.031) | (0.029) | |||||
-| GVAO-U | 0.430∗∗∗ | 0.435∗∗∗ | ||||
-| (0.037) | (0.035) | |||||
-| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
-| Adjusted R2 | 0.531 | 0.609 | 0.659 | 0.760 | 0.763 | 0.796 |
-| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
-| Dependent variable: | ||||||
-| CH4pc | CH4i | |||||
-| (1) | (2) | (3) | (4) | (5) | (6) | |
-| Low Carbon Exp. pc | 0.134∗ | 0.100 | 0.103 | 0.136∗ | 0.132∗∗ | 0.199∗∗∗ |
-| (0.079) | (0.064) | (0.064) | (0.075) | (0.065) | (0.066) | |
-| Low Carbon Exp. pc | 0.378 | 0.077 | 0.114 | 0.130 | 0.580 | 0.965∗∗ |
-| (0.258) | (0.471) | (0.468) | (0.247) | (0.473) | (0.483) | |
-| Low Carbon Exp. pcTER | 0.037 | 0.033 | 0.044 | 0.044 | 0.138 | 0.080 |
-| (0.151) | (0.131) | (0.130) | (0.145) | (0.132) | (0.134) | |
-| Low Carbon Exp. pcTER | 2.751∗∗∗ | 2.460∗∗∗ | 3.132∗∗∗ | 3.133∗∗∗ | 3.258∗∗∗ | 2.488∗∗∗ |
-| (0.507) | (0.928) | (0.923) | (0.487) | (0.933) | (0.952) | |
-| Low Carbon Exp. pcLDR | 0.406∗∗∗ | 0.429∗∗∗ | 0.352∗∗∗ | 0.038 | 0.165∗∗ | 0.177∗∗ |
-| (0.097) | (0.081) | (0.081) | (0.093) | (0.081) | (0.083) | |
-| Low Carbon Exp. pcLDR | 0.404 | 1.339∗∗∗ | 0.836∗ | 0.683∗∗∗ | 0.163 | 0.034 |
-| (0.258) | (0.484) | (0.481) | (0.248) | (0.486) | (0.496) | |
-| GDP | 0.172∗∗∗ | 0.100∗∗∗ | 0.668∗∗∗ | 0.722∗∗∗ | ||
-| (0.020) | (0.021) | (0.019) | (0.021) | |||
-| EPS | 0.026∗∗∗ | 0.019∗∗∗ | 0.030∗∗∗ | 0.021∗∗∗ | ||
-| (0.005) | (0.005) | (0.005) | (0.005) | |||
-| Political Ideology | 0.003 | 0.003 | 0.017∗∗∗ | 0.011∗∗ | ||
-| (0.004) | (0.004) | (0.004) | (0.004) | |||
-| GVAA | 0.023 | 0.007 | ||||
-| (0.015) | (0.015) | |||||
-| GVAB-E | 0.108∗∗∗ | 0.191∗∗∗ | ||||
-| (0.018) | (0.018) | |||||
-| GVAF | 0.136∗∗∗ | 0.214∗∗∗ | ||||
-| (0.016) | (0.017) | |||||
-| GVAG-J | 0.112∗∗∗ | 0.001 | ||||
-| (0.034) | (0.036) | |||||
-| GVAK-N | 0.010 | 0.189∗∗∗ | ||||
-| (0.032) | (0.033) | |||||
-| GVAO-U | 0.170∗∗∗ | 0.175∗∗∗ | ||||
-| (0.038) | (0.039) | |||||
-| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
-| Adjusted R2 | 0.345 | 0.473 | 0.490 | 0.683 | 0.716 | 0.711 |
-| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
-| Dependent variable: | ||||||
-| N2Opc | N2Oi | |||||
-| (1) | (2) | (3) | (4) | (5) | (6) | |
-| Low Carbon Exp. pc | 0.108 | 0.060 | 0.013 | 0.102∗ | 0.082 | 0.111∗∗ |
-| (0.071) | (0.055) | (0.054) | (0.061) | (0.053) | (0.055) | |
-| Low Carbon Exp. pc | 0.189 | 0.052 | 0.553 | 0.366∗ | 0.452 | 0.291 |
-| (0.233) | (0.403) | (0.397) | (0.199) | (0.386) | (0.402) | |
-| Low Carbon Exp. pcTER | 0.254∗ | 0.036 | 0.104 | 0.353∗∗∗ | 0.240∗∗ | 0.021 |
-| (0.136) | (0.113) | (0.110) | (0.116) | (0.108) | (0.111) | |
-| Low Carbon Exp. pcTER | 1.804∗∗∗ | 1.333∗ | 2.032∗∗∗ | 2.275∗∗∗ | 2.352∗∗∗ | 1.364∗ |
-| (0.457) | (0.795) | (0.781) | (0.391) | (0.760) | (0.792) | |
-| Low Carbon Exp. pcLDR | 0.693∗∗∗ | 0.344∗∗∗ | 0.290∗∗∗ | 0.331∗∗∗ | 0.085 | 0.116∗ |
-| (0.087) | (0.069) | (0.068) | (0.075) | (0.066) | (0.069) | |
-| Low Carbon Exp. pcLDR | 2.399∗∗∗ | 1.722∗∗∗ | 1.273∗∗∗ | 1.297∗∗∗ | 0.534 | 0.465 |
-| (0.233) | (0.414) | (0.407) | (0.199) | (0.396) | (0.412) | |
-| GDP | 0.103∗∗∗ | 0.135∗∗∗ | 0.759∗∗∗ | 0.714∗∗∗ | ||
-| (0.018) | (0.019) | (0.015) | (0.018) | |||
-| EPS | 0.006 | 0.010∗∗ | 0.010∗∗ | 0.012∗∗ | ||
-| (0.004) | (0.004) | (0.004) | (0.005) | |||
-| Political Ideology | 0.016∗∗∗ | 0.010∗∗∗ | 0.003 | 0.001 | ||
-| (0.004) | (0.004) | (0.003) | (0.004) | |||
-| GVAA | 0.091∗∗∗ | 0.062∗∗∗ | ||||
-| (0.013) | (0.013) | |||||
-| GVAB-E | 0.113∗∗∗ | 0.186∗∗∗ | ||||
-| (0.015) | (0.015) | |||||
-| GVAF | 0.040∗∗∗ | 0.039∗∗∗ | ||||
-| (0.014) | (0.014) | |||||
-| GVAG-J | 0.009 | 0.101∗∗∗ | ||||
-| (0.029) | (0.030) | |||||
-| GVAK-N | 0.188∗∗∗ | 0.369∗∗∗ | ||||
-| (0.027) | (0.027) | |||||
-| GVAO-U | 0.032 | 0.042 | ||||
-| (0.036) | (0.036) | |||||
-| Observations | 3,856 | 2,982 | 2,982 | 3,856 | 2,982 | 2,982 |
-| Adjusted R2 | 0.109 | 0.159 | 0.203 | 0.594 | 0.619 | 0.595 |
-| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
-Cycle components
-| Dependent variable: | ||||||
-| CO2pc | CO2i | |||||
-| (1) | (2) | (3) | (4) | (5) | (6) | |
-| Low Carbon Exp. pc | 0.072 | 0.003 | 0.008 | 0.059 | 0.001 | 0.004 |
-| (0.054) | (0.054) | (0.054) | (0.054) | (0.054) | (0.054) | |
-| Low Carbon Exp. pc | 0.162 | 0.482 | 0.394 | 0.152 | 0.463 | 0.315 |
-| (0.177) | (0.395) | (0.394) | (0.177) | (0.396) | (0.397) | |
-| Low Carbon Exp. pcTER | 0.096 | 0.308∗∗∗ | 0.374∗∗∗ | 0.089 | 0.311∗∗∗ | 0.377∗∗∗ |
-| (0.104) | (0.108) | (0.108) | (0.103) | (0.108) | (0.108) | |
-| Low Carbon Exp. pcTER | 0.115 | 1.745∗∗ | 1.918∗∗ | 0.134 | 1.871∗∗ | 1.668∗∗ |
-| (0.344) | (0.763) | (0.760) | (0.342) | (0.763) | (0.765) | |
-| Low Carbon Exp. pcLDR | 0.062 | 0.052 | 0.063 | 0.069 | 0.067 | 0.067 |
-| (0.067) | (0.068) | (0.068) | (0.067) | (0.068) | (0.068) | |
-| Low Carbon Exp. pcLDR | 0.024 | 0.681∗ | 0.721∗ | 0.030 | 0.674∗ | 0.494 |
-| (0.179) | (0.407) | (0.407) | (0.178) | (0.408) | (0.410) | |
-| GDP | 0.567∗∗∗ | 0.660∗∗∗ | 0.428∗∗∗ | 0.323∗∗∗ | ||
-| (0.034) | (0.042) | (0.034) | (0.042) | |||
-| EPS | 0.036∗∗∗ | 0.041∗∗∗ | 0.028∗∗∗ | 0.028∗∗∗ | ||
-| (0.004) | (0.005) | (0.004) | (0.005) | |||
-| Political Ideology | 0.001 | 0.0004 | 0.003 | 0.002 | ||
-| (0.003) | (0.003) | (0.003) | (0.003) | |||
-| GVAA | 0.006 | 0.010 | ||||
-| (0.010) | (0.010) | |||||
-| GVAB-E | 0.125∗∗∗ | 0.069∗∗∗ | ||||
-| (0.019) | (0.019) | |||||
-| GVAF | 0.100∗∗∗ | 0.044∗∗∗ | ||||
-| (0.015) | (0.015) | |||||
-| GVAG-J | 0.078∗∗∗ | 0.179∗∗∗ | ||||
-| (0.030) | (0.030) | |||||
-| GVAK-N | 0.125∗∗∗ | 0.087∗∗ | ||||
-| (0.039) | (0.039) | |||||
-| GVAO-U | 0.291∗∗∗ | 0.103∗∗ | ||||
-| (0.044) | (0.044) | |||||
-| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
-| Adjusted R2 | 0.206 | 0.261 | 0.270 | 0.044 | 0.042 | 0.041 |
-| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
-| Dependent variable: | ||||||
-| CH4pc | CH4i | |||||
-| (1) | (2) | (3) | (4) | (5) | (6) | |
-| Low Carbon Exp. pc | 0.198∗∗∗ | 0.174∗∗∗ | 0.184∗∗∗ | 0.185∗∗∗ | 0.173∗∗∗ | 0.172∗∗∗ |
-| (0.047) | (0.048) | (0.048) | (0.047) | (0.047) | (0.048) | |
-| Low Carbon Exp. pc | 0.075 | 0.188 | 0.168 | 0.085 | 0.207 | 0.247 |
-| (0.152) | (0.348) | (0.348) | (0.152) | (0.346) | (0.353) | |
-| Low Carbon Exp. pcTER | 0.152∗ | 0.237∗∗ | 0.256∗∗∗ | 0.146 | 0.240∗∗ | 0.259∗∗∗ |
-| (0.089) | (0.095) | (0.095) | (0.089) | (0.094) | (0.096) | |
-| Low Carbon Exp. pcTER | 0.699∗∗ | 1.353∗∗ | 1.455∗∗ | 0.718∗∗ | 1.479∗∗ | 1.206∗ |
-| (0.295) | (0.671) | (0.670) | (0.294) | (0.667) | (0.681) | |
-| Low Carbon Exp. pcLDR | 0.094 | 0.066 | 0.069 | 0.086 | 0.052 | 0.066 |
-| (0.058) | (0.060) | (0.060) | (0.057) | (0.060) | (0.061) | |
-| Low Carbon Exp. pcLDR | 0.199 | 0.198 | 0.124 | 0.194 | 0.190 | 0.105 |
-| (0.153) | (0.359) | (0.359) | (0.153) | (0.356) | (0.364) | |
-| GDP | 0.163∗∗∗ | 0.181∗∗∗ | 0.831∗∗∗ | 0.802∗∗∗ | ||
-| (0.029) | (0.037) | (0.029) | (0.037) | |||
-| EPS | 0.004 | 0.008∗∗ | 0.004 | 0.006 | ||
-| (0.004) | (0.004) | (0.004) | (0.004) | |||
-| Political Ideology | 0.002 | 0.001 | 0.002 | 0.001 | ||
-| (0.003) | (0.003) | (0.003) | (0.003) | |||
-| GVAA | 0.009 | 0.007 | ||||
-| (0.009) | (0.009) | |||||
-| GVAB-E | 0.048∗∗∗ | 0.146∗∗∗ | ||||
-| (0.016) | (0.017) | |||||
-| GVAF | 0.074∗∗∗ | 0.018 | ||||
-| (0.013) | (0.013) | |||||
-| GVAG-J | 0.002 | 0.260∗∗∗ | ||||
-| (0.026) | (0.027) | |||||
-| GVAK-N | 0.011 | 0.223∗∗∗ | ||||
-| (0.034) | (0.035) | |||||
-| GVAO-U | 0.045 | 0.233∗∗∗ | ||||
-| (0.039) | (0.039) | |||||
-| Observations | 3,872 | 2,996 | 2,996 | 3,872 | 2,996 | 2,996 |
-| Adjusted R2 | 0.034 | 0.040 | 0.032 | 0.258 | 0.243 | 0.216 |
-| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
-| Dependent variable: | ||||||
-| N2Opc | N2Oi | |||||
-| (1) | (2) | (3) | (4) | (5) | (6) | |
-| Low Carbon Exp. pc | 0.059 | 0.035 | 0.062 | 0.045 | 0.032 | 0.049 |
-| (0.056) | (0.063) | (0.063) | (0.055) | (0.063) | (0.064) | |
-| Low Carbon Exp. pc | 0.050 | 0.160 | 0.157 | 0.062 | 0.145 | 0.076 |
-| (0.181) | (0.462) | (0.461) | (0.180) | (0.458) | (0.465) | |
-| Low Carbon Exp. pcTER | 0.072 | 0.076 | 0.101 | 0.080 | 0.074 | 0.098 |
-| (0.106) | (0.126) | (0.126) | (0.105) | (0.125) | (0.127) | |
-| Low Carbon Exp. pcTER | 0.245 | 0.718 | 1.139 | 0.263 | 0.844 | 0.893 |
-| (0.351) | (0.891) | (0.889) | (0.348) | (0.883) | (0.896) | |
-| Low Carbon Exp. pcLDR | 0.191∗∗∗ | 0.214∗∗∗ | 0.193∗∗ | 0.184∗∗∗ | 0.200∗∗ | 0.190∗∗ |
-| (0.069) | (0.080) | (0.080) | (0.068) | (0.079) | (0.080) | |
-| Low Carbon Exp. pcLDR | 0.065 | 0.050 | 0.016 | 0.070 | 0.057 | 0.250 |
-| (0.183) | (0.476) | (0.476) | (0.181) | (0.472) | (0.480) | |
-| GDP | 0.140∗∗∗ | 0.157∗∗∗ | 0.856∗∗∗ | 0.826∗∗∗ | ||
-| (0.035) | (0.049) | (0.035) | (0.049) | |||
-| EPS | 0.0002 | 0.004 | 0.007 | 0.017∗∗∗ | ||
-| (0.005) | (0.005) | (0.005) | (0.005) | |||
-| Political Ideology | 0.009∗∗ | 0.007∗ | 0.005 | 0.005 | ||
-| (0.004) | (0.004) | (0.004) | (0.004) | |||
-| GVAA | 0.029∗∗ | 0.012 | ||||
-| (0.012) | (0.012) | |||||
-| GVAB-E | 0.039∗ | 0.156∗∗∗ | ||||
-| (0.022) | (0.022) | |||||
-| GVAF | 0.019 | 0.038∗∗ | ||||
-| (0.017) | (0.018) | |||||
-| GVAG-J | 0.114∗∗∗ | 0.147∗∗∗ | ||||
-| (0.035) | (0.035) | |||||
-| GVAK-N | 0.166∗∗∗ | 0.383∗∗∗ | ||||
-| (0.046) | (0.046) | |||||
-| GVAO-U | 0.031 | 0.154∗∗∗ | ||||
-| (0.052) | (0.052) | |||||
-| Observations | 3,856 | 2,982 | 2,982 | 3,856 | 2,982 | 2,982 |
-| Adjusted R2 | 0.018 | 0.011 | 0.020 | 0.241 | 0.175 | 0.156 |
-| Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
+Testing for instruments exogeneity
+| Dependent variable: | ||||
+| Trend residual | Cycle residual | |||
+| (3t) | (6t) | (3c) | (6c) | |
+| GHGEpc | GHGEi | GHGEpc | GHGEi | |
+| Labour productivity ppet-8 | 0.003 | 0.002 | -0.002 | 0.000 |
+| (0.004) | (0.004) | (0.003) | (0.003) | |
+| Consumption of fixed capitalt-8 | -0.001 | -0.002 | 0.004 | 0.001 |
+| (0.005) | (0.005) | (0.004) | (0.004) | |
+| Hours worked per capitat-8 | 0.000 | 0.000 | -0.000 | -0.000 |
+| (0.000) | (0.000) | (0.000) | (0.000) | |
+| Compensation of Employeest-8 | 0.000 | 0.002 | -0.004 | -0.001 |
+| (0.005) | (0.005) | (0.004) | (0.004) | |
+| TER | 0.002 | 0.003 | -0.003 | 0.000 |
+| (0.004) | (0.004) | (0.003) | (0.003) | |
+| LDR | 0.000 | 0.002 | -0.002 | -0.000 |
+| (0.003) | (0.003) | (0.002) | (0.002) | |
+| Constant | -0.031 | -0.039 | 0.036 | -0.001 |
+| (0.043) | (0.044) | (0.032) | (0.032) | |
+| Observations | 3,408 | 3,408 | 3,408 | 3,408 |
+| R-squared | 0.000 | 0.001 | 0.001 | 0.000 |
+| Notes: Standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||

### v3

@@
 We investigate the effectiveness of low-carbon expenditures from the European Structural and Investment Funds in reducing greenhouse gas emissions across EU regions.
-Using trend and cycle decomposition of two indicators – per capita emissions and emissions intensity – we find highly heterogeneous effects.
-In less developed regions, investments are linked to long-term increases in per capita emissions, while in transition and developed regions, effects are weak or not significant.
-Instrumental variable estimates confirm this pattern.
+Using trend and cycle decomposition of per capita emissions and emissions intensity, along with a panel data approach that incorporates long lags to mitigate reverse causality, we find highly heterogeneous effects.
+In less developed regions, investments are associated with long-term increases in per capita emissions, whereas in transition and developed regions, the effects are weak or not significant.
 When disaggregated by gas type, results remain inconsistent.
-Overall, our findings highlight the limited climate effectiveness of EU funding and the need for region-specific strategies.
+Our findings highlight that regional disparities challenge the effectiveness of EU climate efforts.
 Keywords: Greenhouse gas emissions; Green transition; Climate action; Low-carbon economy; European Structural and Investment Funds; Regional development
+JEL Codes: Q56; Q58
 1 Introduction
@@
 We focus on the relationship between economic growth and environmental sustainability, and, to that end, we decompose the emission time series of each region into trend and cycle components.
-This allows us to determine expenditures’ short-term and long-term effects, considering that some regions have significantly advanced in terms of sustainability – achieving a given degree of decoupling – whereas many others have yet to do so.
+This allows us to determine expenditures’ short-term and long-term effects, considering that some regions have significantly advanced in terms of sustainability—achieving a given degree of decoupling—whereas many others have yet to do so.
 Our analysis also accounts for the direct impacts of policy by considering the OECD Environmental Policy Stringency (EPS) index (OECD, 2016) and average regional political opinions as covariates.
+To address potential endogeneity and reverse causality, we rely on a design with long lags of all right-hand-side variables, different specifications with controls including region-specific industry-composition (proxied by sectoral Gross Value Added (GVA)), and region and year fixed effects.
 We find heterogeneous effects of low-carbon expenditures on GHG emissions across European regions.
@@
 In contrast, results for developed and transition regions do not reveal significant reductions in long-term emissions trends, although there is some evidence of short-term mitigating effects, particularly for the 2007–2013 programme.
-To address potential endogeneity, we implement a two-stage least squares (2SLS) model using lagged structural indicators and regional development categories as instruments.
-The instrumental variables estimates reinforce the main findings: low-carbon investments have heterogeneous effects across regions, with increases in per capita emissions in less developed regions and short-term reductions in developed regions.
-In transition regions, distinct patterns emerge that ultimately underscore the importance of the regional context.
 When disaggregating emissions by gas type, we consistently find marked regional heterogeneity.
-The results reveal no significant reductions in emissions for developed regions, while in less developed regions, emissions – especially CO2 and N2O per capita – increase significantly.
+The results reveal no significant reductions in emissions for developed regions, while in less developed regions, emissions—especially CO2 and N2O per capita—increase significantly.
 Effects in transition regions are mixed.
@@
 Regarding the drivers behind carbon footprints, the authors identified the relevance of household size, whether an area is urban or rural, the population’s education level, expenditure patterns, local temperature, resource availability, and the carbon intensity of the local electricity mix.
-Yet a significant strand of literature doubts that the downward-sloping relationships between higher-income countries and pollution levels – which might imply a decoupling of economic activity with emissions – are enough to face climate change efficiently.
+Yet a significant strand of literature doubts that the downward-sloping relationships between higher-income countries and pollution levels—which might imply a decoupling of economic activity with emissions—are enough to face climate change efficiently.
 Certainly, there are concerns about the possible shift in the production of pollution-intensive goods from developed to developing countries with less stringent environmental regulations (Hertwich and Peters, 2009; Levinson, 2023).
@@
 In this context, a growing body of literature indicates that international environmental agreements and climate policies are increasingly impacting emission trends, as reflected in the economic practices within countries, contributing to our understanding of the decoupling mechanism (Fischer and Heutel, 2013; Hildén et al., 2014; Haberl et al., 2020).
-Decoupling analyses aim to categorise the patterns of how emissions respond to economic growth – whether they increase at a slower rate, decrease, or stabilise as the economy grows (relative and absolute decoupling) or increase faster than the economy (coupling).
+Decoupling analyses aim to categorise the patterns of how emissions respond to economic growth—whether they increase at a slower rate, decrease, or stabilise as the economy grows (relative and absolute decoupling) or increase faster than the economy (coupling).
 Cohen et al.
@@
 Another fact is that emissions exhibit greater volatility compared to GDP, responding more dramatically to economic changes, and the extent to which emissions volatility exceeds GDP volatility decreases with the income level of the economies (Doda, 2014).
-These patterns are also observed at the subnational level: Dueñas and Mandel (2025), analysing NUTS 2 regions in Europe, further show that CO2 emissions growth rates follow an asymmetric and fat-tailed distribution and, similarly, that emissions volatility scales inversely with economic size of regions.
+These patterns are also observed at the subnational level: Dueñas and Mandel (2025), analysing NUTS 2 regions in Europe, further show that CO2 emissions growth rates follow an asymmetric and fat-tailed distribution and, similarly, that emissions volatility scales inversely with the economic size of regions.
 This implies that more stable emission patterns through economic cycles in developed regions could be linked to more efficient energy use, a greater share of services in the economy, and potentially more effective environmental regulations and policies.
@@
 The study notes progress in detaching economic growth from emissions under the EU ETS and production-based accounting, but they find no advancement in separating economic growth from consumption-based accounting emissions in either group.
-More directly related to the impacts of the EU Cohesion Policy, much of the literature has focused primarily on understanding its impact on regional development, specifically whether the funds have been effective in reducing regional disparities.
+While the above studies highlight the cyclical and long-term dynamics of emissions in response to policy, the Cohesion Policy literature has concentrated mainly on regional development objectives, which is its primary objective.
+Much of the literature has focused primarily on understanding its impact on regional development, specifically whether the funds have been effective in reducing regional disparities.
 For instance, very early on, Boldrin and Canova (2001) pointed out that these policies tend to be more redistributive than growth-oriented, reflecting political compromises over economic efficiency.
@@
 Institutional quality plays a relevant role in the effectiveness of the funds.
-Evidence shown that local and regional governance directly affects economic growth and the effectiveness of structural funds (Rodríguez-Pose and Garcilazo, 2015; Charron et al., 2014).
+Evidence shows that local and regional governance directly affects economic growth and the effectiveness of structural funds (Rodríguez-Pose and Garcilazo, 2015; Charron et al., 2014).
 Although relatively few studies have explicitly focused on the effect of EU structural funds on reducing emissions, some anticipate that many of the same structural patterns that determine growth performance may also shape the trajectory of sustainable growth.
+In particular, some contributions emphasise the importance of regional asymmetries in the green transition.
 Rodríguez-Pose and Bartalucci (2024) warn that while the green transition may benefit certain regions, it can also exacerbate existing territorial inequalities, particularly affecting less-developed regions.
-Nichifor et al.
-(2025) find that investments in renewables and environmental expenditures have positive effects on emissions reduction, but caution that these effects may exhibit diminishing returns beyond certain thresholds.
+They claim that less-developed, peri-urban, and rural regions in Southern and Eastern Europe are systematically more vulnerable, while metropolitan areas are better positioned to embrace green opportunities thanks to stronger infrastructure, skills, and institutional capacity.
+At the same time, the micro-foundations of the green upgrading further stress the role of regional capability-relatedness (Boschma et al., 2017).
+Santoalha and Boschma (2021) show that regions are more likely to diversify into green technologies when related capabilities are present.
+They also claim that political support does not exert a strong direct effect but interacts with capabilities: regional-level support strengthens, while national-level support may moderate the role of relatedness.
+Consistent with this territorial heterogeneity, Dueñas and Mandel (2025) show that European regions experience markedly different emission dynamics under the EU ETS, with less-developed regions displaying higher volatility and more uncertain decarbonisation pathways than their richer counterparts, suggesting that the growth–emissions nexus is mediated by regional characteristics, institutions, and historical trajectories.
+Complementing this evidence, Nichifor et al.
+(2025) find that in Eastern Europe, renewable investments and environmental expenditures reduce emissions on average, but their effectiveness is spatially uneven and may exhibit diminishing returns beyond policy-relevant thresholds, underscoring the need for coordinated, place-based strategies.
 In summary, in light of these findings, the effectiveness of the ESI funds in driving a low-carbon transition is conditional not only on structural factors such as institutional quality and regional development levels, but also on the phase of the economic cycle.
-Under Kuznets’ hypothesis, decoupling might appear problematic since regions would only begin to improve their environmental performance after reaching a certain level of wealth, implying a dependency between economic growth and environmental improvement.
-Therefore, seeking economic development without the corresponding environmental degradation is a great challenge for any policy framework.
-Therefore, the central hypothesis of this paper is as follows:
-Hypothesis: Low‑carbon economy expenditures from the European Structural and Investment Funds have heterogeneous impacts on GHG emissions across regions at different stages of development, influencing both the long‑run trend and the magnitude of the business‑cycle component of emissions.
-To test this hypothesis, rather than focusing on total emissions, we concentrate on per capita emissions and emissions intensity (relative to regional GDP) .
+Less-developed regions, characterised by higher vulnerability and weaker absorptive capacity, may struggle to convert expenditures into immediate emission reductions, potentially exhibiting weaker or delayed effects (and even short-run increases), while more developed regions are better positioned to leverage such funds into abatement.
+These asymmetries align with the logic of the environmental Kuznets hypothesis, suggesting that decoupling may be problematic, as regions tend to improve their environmental performance only after reaching a certain level of wealth.
+Therefore, seeking economic development while reducing disparities across European regions, and at the same time avoiding environmental degradation, remains a major challenge in the EU policy framework.
+Accordingly, the central hypothesis of this paper is as follows:
+Hypothesis 1: Low‑carbon economy expenditures from the European Structural and Investment Funds have heterogeneous impacts on GHG emissions across regions at different stages of development, influencing both the long‑run trend and the magnitude of the business‑cycle component of emissions.
+Building on recent contributions emphasising regional asymmetries of European regions (as in Santoalha and Boschma 2021 and Rodríguez-Pose and Bartalucci 2024), it can be further expected the following:
+Hypothesis 2: The effectiveness of low-carbon expenditures is conditional on regional development levels, with weaker—and in some cases emission-increasing—impacts in less-developed regions.
+To test these hypotheses, rather than focusing on total emissions, we concentrate on per capita emissions and emissions intensity (relative to regional GDP) .
 These indicators enhance comparability across regions by accounting for differences in population and economic scale, while also isolating environmental efficiency more effectively.
@@
 We aggregate investments at NUTS level 2 (NUTS 2), as defined in version 2021.
-To do this, we consider all related investments, regardless of their original level of geographic specificity in the original database – be it national, NUTS-1, or even more granular at NUTS-3.
+To do this, we consider all related investments, regardless of their original level of geographic specificity in the original database—be it national, NUTS-1, or even more granular at NUTS-3.
 Figure 1 shows the calculated distribution of total ESI funds expenditures reported by region for the periods 2007-2013 and 2014-2020, along with the eligibility of the regions to access these funds.
@@
 The aim of the expenditure can be tracked by the reported “intervention field” in the data.
-Specifically, related to TO-4, the intervention fields include renewable energy (wind/solar/biomass), energy efficiency renovation of public infrastructure, cycle tracks and footpaths, among others (see Table A1 in the Appendix).
+Specifically, related to TO-4, the intervention fields include renewable energy (wind/solar/biomass), energy efficiency renovation of public infrastructure, cycle tracks and footpaths, among others (see Appendix A1).
 For the 2007-2013 programme, there is no categorical variable explicitly associated with the low-carbon economy as in the 2014-2020 programme.
@@
 Priority labels related to this period include cycle tracks, intelligent transport systems, renewable energy (wind/solar/biomass/hydroelectric), energy efficiency, co-generation, energy management, and promotion of clean urban transport.
-The priority codes are reported in Table A1 of the Appendix.
+The priority codes are reported in Appendix A1.
 A critical aspect of the ESI funds expenditure data is the absence of detailed annual expenditure information at the regional level.
@@
 Therefore, considering the capital-intensive nature of the funded projects, we argue that their environmental effects are plausibly associated with the cumulative stock of investments.
-We further validate this modelling choice through robustness checks, including alternative cross-sectional estimations presented in Appendix Tables A2 and A3.
+We further validate this modelling choice through robustness checks, including alternative cross-sectional estimations presented in Appendix A2.
 Supplementary covariates
 Economic variables were obtained from the ARDECO (the Annual Regional Database of the European Commission’s Directorate General for Regional and Urban Policy), which contains annual data at the regional level and is maintained and updated by the Joint Research Centre of the European Commission.
-We use variables such as the real gross domestic product, the population, and the real gross value added (GVA) for all industry sectors according to NACE Rev.
+We use variables such as the real gross domestic product, the population, and the real gross value added for all industry sectors according to NACE Rev.
 2 definitions.
@@
 For instance, one might argue that environmental regulations or improvements in energy efficiency could, in principle, influence regional output and investment decisions.
-For these reasons, we have included a five-year lag for all explanatory variables, ensuring that the predictors precede the observed emissions outcomes by a meaningful temporal distance.
-Therefore, while our specification mitigates (but does not fully eliminate) simultaneity concerns, it relies on the timing assumption that lagged expenditures impact emissions.
 Moreover, the allocation of ESI funds is determined based on regional eligibility criteria related to development levels, making the investment variable exogenous to current emission levels.
@@
 Nonetheless, one could still argue that regions with higher emissions may have received more funds, representing a case of reverse causality or omitted variable bias.
-To test the robustness of our identification, we implement a 2SLS estimation using structural and institutional instruments plausibly uncorrelated with current emissions.
+For these reasons, we have included a five-year lag for all explanatory variables, ensuring that the predictors precede the observed emissions outcomes by a meaningful temporal distance.
+Therefore, our specification mitigates—though does not fully eliminate—simultaneity concerns and relies on the timing assumption that lagged expenditures impact emissions.
+A plausible strategy would have been to implement instrumental variables; however, given the nature of the impact of low-carbon investments—which take time to materialise—it is difficult to identify strong instruments.888As a robustness check, in Appendix A3, we implemented 2SLS estimations using structural and institutional variables (e.g.
+labour productivity, hours worked per capita, consumption of fixed capital, compensation of employees) as potential instruments.
+Despite these instruments proved to be valid, they are not strong enough to support a reliable specification, specifically the Kleibergen-Paap rk Wald F statistic is well below conventional thresholds.
+Thus, these IV results must be interpreted with caution.
 4 Results
@@
 As in the GHGEpc models, results for transition regions remain statistically not significant across all models.
+These results support H1, as the effects of low-carbon expenditures are heterogeneous across development levels and emission indicators.
+Moreover, the consistently positive and significant interactions for LDR alongside weak or null effects for DER/TER provide evidence in favour of H2.
 Control variables, particularly GDP and sectoral GVA, exhibit expected significant relationships with emissions trends.
@@
 As in the per capita models, the effects of the 2014–2020 programme are not statistically significant for any regional group.
+These findings reinforce H1 and H2, showing that low-carbon investments affect not only long-run trajectories but also the business-cycle component of emissions, with short-run impacts that vary heterogeneously across regions depending on their level of development.
 Lagged GDP is negatively associated with per capita emissions cycles and positively with emissions intensity cycles, consistent with short-run adjustments in the emissions-output relationship.
@@
 These results are in line with recent findings for European regions, which report that emissions growth rates are highly volatile and sensitive to policy phases, particularly within the EU ETS framework (Dueñas and Mandel, 2025).
-4.3 Instrumental variables estimations
-To address potential endogeneity concerns related to low-carbon investments, we complement our baseline fixed-effects specification with an instrumental variables (2SLS) approach.
-While the use of lagged explanatory variables in the baseline models helps mitigate simultaneity bias, endogeneity may persist due to omitted variables or reverse causality.
-For instance, structurally more proactive regions might both attract higher low-carbon funding and implement complementary environmental measures not fully captured by our controls, potentially biasing the estimated coefficients.
-We instrument per capita low-carbon expenditures from both programming periods (2007–2013 and 2014–2020) using regional structural indicators from the ARDECO database: (i) labour productivity per employee (log), (ii) consumption of fixed capital (log), (iii) hours worked per capita, and (iv) compensation of employees (log).
-These instruments are lagged by three years, not only to ensure temporal separation from emissions outcomes, but also based on the assumption that the allocation of funds is determined by pre-existing structural conditions.
-Additionally, we include regional development level dummies (TER and LDR, with DER as the reference category), which reflect the eligibility and allocation logic of the ESI funds.
-These variables are plausibly exogenous to short- and medium-term emissions dynamics.
-First-stage results, presented in Table 4, confirm the strength of our instruments, with F-statistics exceeding conventional thresholds and statistically significant coefficients in the expected direction.
-| Low Carbon Exp. pc07-13 | Low Carbon Exp. pc14-20 | |
-| Labour productivity ppet-3 | -1.033*** | -0.683*** |
-| (0.114) | (0.071) | |
-| Hours worked per capitat-3 | 0.003*** | 0.001*** |
-| (0.000) | (0.000) | |
-| Compensation of Employeest-3 | -1.836*** | 0.422*** |
-| (0.237) | (0.109) | |
-| Consumption of fixed capitalt-3 | 1.990*** | -0.585*** |
-| (0.237) | (0.110) | |
-| TER | 0.970*** | 0.559*** |
-| (0.117) | (0.066) | |
-| LDR | 1.527*** | 1.329*** |
-| (0.145) | (0.080) | |
-| Time Dummies | Yes | Yes |
-| F-statistic | 114.670 | 265.060 |
-| Prob F | 0.000 | 0.000 |
-| Observations | 1,694 | 1,687 |
-| Notes: Standard errors in parentheses. Significance levels: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. TER and LDR refer to transition and less developed regions, respectively; DER is the reference group. | ||
-For the second stage, we focus on the specifications that include the most comprehensive set of controls, those considering GVA disaggregated by sector.
-The column structure is intentionally aligned with the baseline results to facilitate direct comparison.
-Table 5 presents the second-stage results for both emissions performance indicators.
-Overall, the IV estimates reveal patterns that are qualitatively consistent with those obtained from the fixed-effects specification.
-To further support the validity of our instruments, we regress the second-stage residuals on the full set of instruments; the absence of significant coefficients suggests that the instruments are uncorrelated with the structural error term.
-Results of this test are reported in Appendix Table A4.
-| Dependent variable: | ||||
-| Trend | Cycle | |||
-| (3t) | (6t) | (3c) | (6c) | |
-| GHGEpc | GHGEi | GHGEpc | GHGEi | |
-| Low Carbon Exp. pc | -0.587 | -0.224 | -0.601* | -0.144 |
-| (0.393) | (0.389) | (0.345) | (0.201) | |
-| Low Carbon Exp. pc | 5.692 | 4.207 | 0.460 | 0.340 |
-| (3.468) | (3.847) | (0.815) | (0.846) | |
-| Low Carbon Exp. pcTER | 1.399* | 1.949*** | -0.300 | 0.092 |
-| (0.732) | (0.695) | (0.484) | (0.399) | |
-| Low Carbon Exp. pcTER | -0.779 | 0.392 | -0.551 | -0.639 |
-| (3.041) | (3.005) | (0.665) | (0.702) | |
-| Low Carbon Exp. pcLDR | 1.126*** | 0.539 | 0.357 | 0.116 |
-| (0.408) | (0.387) | (0.339) | (0.204) | |
-| Low Carbon Exp. pcLDR | -2.350 | -2.487 | -0.556 | -0.360 |
-| (3.189) | (3.462) | (0.740) | (0.772) | |
-| EPSt-5 | 0.001 | -0.001 | 0.009*** | 0.011*** |
-| (0.007) | (0.007) | (0.003) | (0.003) | |
-| Political Ideologyt-5 | 0.006 | -0.010 | 0.003 | 0.002 |
-| (0.010) | (0.013) | (0.003) | (0.003) | |
-| GVAA,t-5 | 0.076** | 0.044 | -0.012 | -0.011 |
-| (0.037) | (0.038) | (0.009) | (0.007) | |
-| GVAB-E,t-5 | 0.213*** | -0.064 | -0.008 | -0.003 |
-| (0.057) | (0.078) | (0.018) | (0.018) | |
-| GVAF,t-5 | -0.110** | -0.134** | -0.046*** | 0.015 |
-| (0.048) | (0.058) | (0.015) | (0.017) | |
-| GVAG-J,t-5 | 0.205* | -0.129 | 0.018 | 0.078** |
-| (0.113) | (0.129) | (0.035) | (0.032) | |
-| GVAK-N,t-5 | -0.389*** | -0.436*** | -0.027 | 0.030 |
-| (0.079) | (0.094) | (0.039) | (0.038) | |
-| GVAO-U,t-5 | 0.147 | 0.303** | -0.180*** | -0.045 |
-| (0.102) | (0.125) | (0.042) | (0.041) | |
-| Observations | 3,408 | 3,408 | 3,408 | 3,408 |
-| NUTS 2 fixed effects | Yes | Yes | Yes | Yes |
-| Time fixed effects | Yes | Yes | Yes | Yes |
-| Notes: Robust standard errors clustered at the NUTS 2 level are reported in parentheses.
-Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01.
-Low Carbon Exp.
-pc07-13 and Low Carbon Exp.
-pc14-20 are instrumented.
-| ||||
-For emissions per capita trends (Column 3t), low-carbon expenditures from the 2007–2013 programme are not statistically significant in developed regions, but the interaction term for less-developed regions remains positive and significant.
-This confirms that these regions experienced increases in long-term emissions per capita following higher investments.
-A key difference from the baseline is that the TER interaction becomes positive and significant, whereas it was previously negative and not significant.
-This suggests potential divergence in the response of transition regions under the IV approach.
-The 2014–2020 effects remain statistically not significant across development levels.
-Regarding emissions intensity trends (Column 6t), we find a significant positive effect for the 2007-2013 expenditures in transition regions.
-This result departs from the baseline findings, where all interaction effects were not statistically significant.
-Turning to the cyclical components, for emissions per capita (Column 3c), the 2007–2013 programme has a negative and statistically significant effect in developed regions, consistent with a smoothing effect on short-term volatility.
-However, the previously significant negative interaction effects for LDR are no longer significant under the IV specification.
-Similarly, for emissions intensity cycles (Column 6c), none of the coefficients related to expenditures are statistically significant, contrasting with the baseline results where some mitigating effects were observed.
-In summary, the IV results reinforce the main conclusion of our baseline specification: low-carbon investments have different effects on emissions across different development levels.
-While some coefficients change in significance or sign, the overarching pattern, particularly the long-term positive association in less-developed regions, remains robust.
-These findings enhance confidence in the causal interpretation of the estimated effects and underscore the importance of considering regional structural differences when assessing climate policy outcomes.
 4.3 Effect by type of gases
@@
 We adopt the same methodological approach used for total GHG emissions: decomposing the per capita and intensity time series for each gas into trend and cycle components, and estimating the effects of expenditures on these indicators following the specification in Eq.(3).
-Results are reported in Tables 6 and 7, corresponding to the trend and cycle components, respectively.
+Results are reported in Tables 4 and 5, corresponding to the trend and cycle components, respectively.
 For readability, the estimated coefficients of control variables are omitted from the tables.
@@
 For the 2014–2020 programming period, interaction terms with TER are again negative and statistically significant in models 5 and 6, and for LDR in model 1.
+Therefore, the gas-specific results further support H1, showing that the effects of low-carbon expenditures differ across levels of regional development and are also highly heterogeneous across emission types.
+At the same time, the evidence also appears to support H2, since in less-developed regions, the low-carbon expenditures are positively associated with CO2, CH4, and N2O trends, whereas in transition and developed regions, a negative relationship appears more common, although in many specifications this result is not statistically significant.
 5 Discussion
@@
 In contrast, when examining short-term fluctuations, the results suggest a more positive role for these investments in reducing emissions volatility, especially in less developed regions, indicating that the funds may help stabilise emissions in the short run.
-When we employ instrumental variables to address potential endogeneity, the findings further reinforce the existence of regional disparities.
-The positive effect on per capita emissions in less developed regions is clearly confirmed, and additional significant effects emerge for transition regions that were not visible in the initial analysis.
-While this might suggest that estimations for transition regions were underestimated, the overall pattern is strengthened, reflecting heterogeneous regional impacts.
-Hence, the IV analysis enhances the credibility of a causal interpretation of the differentiated expenditure effects.
 When disaggregating the results by GHG type (CO2, CH4, and N2O), a more nuanced picture emerges.
@@
 Nevertheless, ESI funding remains important, as it aims to address regional disparities and supports place-based transitions.
-Altogether, our findings provide strong empirical support for our central hypothesis.
-While we find consistent evidence that low-carbon economy funding does not lead to long-term reductions in emissions indicators in more developed regions, the adverse effects observed in less developed regions are particularly concerning, given that these areas are the primary targets of ESI fund support.
-Nevertheless, the results suggest that while the funds may not be highly effective in driving long-term decarbonisation, they could still play a stabilising role by dampening short-term emissions volatility.
+Overall, the results are consistent with Hypothesis 1, indicating that low-carbon expenditures have heterogeneous impacts across different levels of regional development.
+Moreover, while we find consistent evidence that low-carbon economy funding does not lead to long-term reductions in emission indicators in more developed regions, the adverse effects observed in less developed areas are particularly concerning, given that these regions are the primary targets of ESI fund support.
+This pattern aligns with Hypothesis 2, which states that the effectiveness of such expenditures is conditional on regional development.
+Nevertheless, when examining the cyclical component of the emissions time series, we find that low-carbon investments may play a stabilising role by dampening short-term emissions volatility.
 It is important to note that most statistically significant effects are concentrated in the 2007-2013 programming period.
@@
 On the one hand, the failure to curb long-term emissions in less developed areas raises questions about allocative efficiency and the risk of reinforcing carbon lock-in (Unruh, 2000).
-On the other hand, mechanisms behind the heterogeneous impacts – such as institutional capacity gaps, infrastructure rigidity, or lack of policy complementarities – remain underexplored and merit further attention (Rodríguez-Pose and Garcilazo, 2015; Charron et al., 2014).
+On the other hand, mechanisms behind the heterogeneous impacts—such as institutional capacity gaps, infrastructure rigidity, or lack of policy complementarities—remain underexplored and merit further attention (Rodríguez-Pose and Garcilazo, 2015; Charron et al., 2014).
 In this respect, Iammarino et al.
@@
 Future research should investigate the mechanisms behind regional disparities and identify the complementary policies needed to align the ESI spending with long-term climate and development objectives.
+It is worth noting that the identification strategy, based on long lags and fixed effects, is aimed to mitigate concerns of reverse causality; however, it may not fully address endogeneity.
+Instrumental variable estimations, using regional structural and institutional variables as instruments, resulted in weak instruments.
+Therefore, future work should seek to strengthen causal inference, for instance by exploiting exogenous quasi-natural experiments (e.g., changes in eligibility thresholds or policy allocation rules).
 6 Conclusions
 This paper examines the climate impact of one of the European Union’s key instruments for regional development: the “Low-Carbon Economy” funds allocated through the European Structural and Investment funds during the 2007-2013 and 2014-2020 programming periods.
-Focusing on per capita emissions and emissions intensity – both decomposed into trend and cyclical components – we estimate the structural (long-run) and cyclical (short-run) effects of low-carbon expenditures across regions.
+Focusing on per capita emissions and emissions intensity—both decomposed into trend and cyclical components—we estimate the structural (long-run) and cyclical (short-run) effects of low-carbon expenditures across regions.
 Our findings point to a substantial degree of heterogeneity.
@@
 These results suggest that, in lagging regions, climate-related investments may stimulate economic activity without achieving the intended decoupling from carbon-intensive processes.
-Instrumental variable estimations reinforce the robustness of our findings and lend support to a causal interpretation of the observed relationships.
+These findings are consistent across specifications with different sets of controls.
 When disaggregating emissions by gas type (CO2, CH4, and N2O), results show further heterogeneity.
@@
 We claim that aligning investment strategies with regional profiles can enhance the effectiveness of emissions reduction efforts.
-Acknowledgements
+Acknoledgment
 The authors acknowledge the financial support of the DECIPHER (Decision-making framework and processes for holistic evaluation of environmental and climate policies) project, funded by the Horizon Europe: Research and Innovation Programme under grant agreement No.
@@
 Appendix
-Low-carbon economy definitions by ESI programmes
+Appendix A1 Low-carbon economy definitions by ESI programmes
 | 2014-2020 funding program | |
@@
 | 25-52 | Promotion of clean urban transport |
-Estimation results modelled expenditures
+Appendix A2 Estimation results modelled expenditures
 The cohesion data platform presents estimations of the expenditure by NUTS 2, which are aggregated for the most significant types of structural funds only.
-Therefore, to address our need for yearly data, we have opted for two approaches: the one presented in the main text, which assumes that expenditures are distributed evenly over time (dividing the reported expenditures over the programming period), and the another one, presented here, which assumes they follow the same pattern as the total aggregated expenditure reported by the cohesion data platform.888Data available at: https://cohesiondata.ec.europa.eu/Other/Historic-EU-payments-regionalised-and-modelled/tc55-7ysv/about_data.
+Therefore, to address our need for yearly data, we have opted for two approaches: the one presented in the main text, which assumes that expenditures are distributed evenly over time (dividing the reported expenditures over the programming period), and the another one, presented here, which assumes they follow the same pattern as the total aggregated expenditure reported by the cohesion data platform.999Data available at: https://cohesiondata.ec.europa.eu/Other/Historic-EU-payments-regionalised-and-modelled/tc55-7ysv/about_data.
 Both strategies yielded qualitatively similar conclusions, as a robustness check the results of the second strategy are shown in Tables A2 and A3.
@@
 | Notes: NUTS 2 cluster robust standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||||
-Testing for instruments exogeneity
+Appendix A3 Instrumental variables (IV/2SLS)
+An instrumental variables approach can be used to address potential endogeneity concerns related to low-carbon investments as a complement to our baseline fixed-effects specification.
+While the use of lagged explanatory variables in the baseline models helps mitigate simultaneity bias, endogeneity may persist due to omitted variables.
+For instance, more proactive regions might both attract higher low-carbon funding and implement complementary environmental measures not fully captured by our controls, potentially biasing the estimated coefficients.
+However, one of the main challenges in identifying valid instruments for low-carbon investments is that, while several structural variables can plausibly explain the allocation of such funds without being directly correlated with emissions, the delayed nature of their effects substantially weakens the instruments’ explanatory power.
+As a result, the fitted values of investment exhibit limited variation, which reduces the likelihood of detecting statistically significant effects in the second stage.
+We present these results below.
+We first identified a set of potential instruments for per capita low-carbon expenditures based on regional structural indicators from the ARDECO database, including labour productivity per employee (log), consumption of fixed capital (log), hours worked per capita, and compensation of employees (log).
+Then, we employed two-stage least squares estimations (2SLS), including year, regional, and eligibility fixed effects, clustering standard errors at the regional level, and controlling for lagged sectoral value added and structural indicators.
+After testing various specifications, we found that, according to the underidentification test, these variables performed reasonably well when considered either individually or in combination.
+Subsequently, we observed that using three- to five-period lags yielded comparatively better first-stage results.
+It is not surprising that lagged versions of these variables perform well, as pre-existing structural conditions are likely to influence the allocation of funds.
+However, these lags are added to the five-period lag already applied to the dependent variable, weakening their potential as strong instruments.
+In Table A4, we report the results for two specifications: the trend of GHG emissions per capita (third column) and the trend of GHG emissions intensity (fourth column), along with their corresponding first-stage regressions (first two columns).
+We present these estimations as an example to illustrate the limitations of our empirical approach.
+In these specifications, the set of instruments is reasonably correlated with the endogenous regressor, as indicated by the Kleibergen-Paap rk LM statistic (greater than 10, with Chi-sq.
+-value 0.05), and there is no evidence against their validity according to the Hansen J test (-value 0.1).
+However, the instruments consistently fail to pass the weak identification tests: the Kleibergen-Paap rk Wald statistic remains well below conventional thresholds ().
+Because this statistic is robust to heteroskedasticity and clustering, it provides the appropriate benchmark in our panel setting.
+After implementing different specifications to exhaustively explore the possibility of addressing potential endogeneity concerns, we therefore conclude that the IV estimates cannot provide reliable evidence and must be interpreted with caution.
+| First stage | Second stage | |||
 | Dependent variable: | ||||
-| Trend residual | Cycle residual | |||
-| (3t) | (6t) | (3c) | (6c) | |
-| GHGEpc | GHGEi | GHGEpc | GHGEi | |
-| Labour productivity ppet-8 | 0.003 | 0.002 | -0.002 | 0.000 |
+| Low Carbon Exp. pc | Low Carbon Exp. pc | GHGE per capita | GHGE intensity | |
+| Low Carbon Exp. pc | -0.604 | -0.101 | ||
+| (0.588) | (0.615) | |||
+| Low Carbon Exp. pc | 3.071 | -7.180 | ||
+| (2.064) | (4.764) | |||
+| Low Carbon Exp. pcTER | 1.558* | 2.676** | ||
+| (0.888) | (1.272) | |||
+| Low Carbon Exp. pcTER | -3.332 | -3.195 | ||
+| (4.136) | (5.033) | |||
+| Low Carbon Exp. pcLDR | 0.535 | 0.446 | ||
+| (0.693) | (0.726) | |||
+| Low Carbon Exp. pcLDR | 2.078 | 7.318 | ||
+| (2.252) | (4.569) | |||
+| EPSt-5 | 0.018*** | 0.003** | -0.004 | -0.015 |
 | (0.004) | (0.001) | (0.010) | (0.011) | |
-| Consumption of fixed capitalt-8 | -0.001 | -0.002 | 0.004 | 0.001 |
+| Political Ideologyt-5 | 0.016*** | 0.004*** | -0.007 | -0.016 |
 | (0.006) | (0.001) | (0.014) | (0.017) | |
-| Hours worked per capitat-8 | 0.000 | 0.000 | -0.000 | -0.000 |
+| GVAA,t-6 | -0.005 | -0.001 | 0.069* | 0.033 |
 | (0.020) | (0.003) | (0.035) | (0.039) | |
-| Compensation of Employeest-8 | 0.000 | 0.002 | -0.004 | -0.001 |
+| GVAB-E,t-6 | 0.023 | 0.010* | 0.142** | -0.078 |
 | (0.028) | (0.005) | (0.070) | (0.092) | |
-| TER | 0.002 | 0.003 | -0.003 | 0.000 |
+| GVAF,t-6 | -0.017 | -0.001 | -0.134*** | -0.126 |
 | (0.018) | (0.007) | (0.051) | (0.079) | |
-| LDR | 0.000 | 0.002 | -0.002 | -0.000 |
+| GVAG-J,t-6 | -0.084 | -0.031*** | 0.376*** | -0.058 |
 | (0.059) | (0.011) | (0.134) | (0.167) | |
-| Constant | -0.031 | -0.039 | 0.036 | -0.001 |
+| GVAK-N,t-6 | 0.051 | 0.057*** | -0.444*** | -0.257** |
 | (0.035) | (0.013) | (0.104) | (0.102) | |
+| GVAO-U,t-6 | -0.128** | 0.015 | 0.112 | 0.233* |
+| (0.052) | (0.014) | (0.109) | (0.137) | |
+| Labour productivity ppet-8 | 0.238*** | 0.029* | ||
+| (0.063) | (0.017) | |||
+| Labour productivity ppet-8TER | -0.217*** | -0.024 | ||
+| (0.066) | (0.021) | |||
+| Labour productivity ppet-8LDR | -0.149** | -0.027 | ||
+| (0.060) | (0.017) | |||
+| Consumption of fixed capitalt-8 | -0.047 | -0.059*** | ||
+| (0.034) | (0.013) | |||
+| Consumption of fixed capitalt-8TER | 0.174*** | 0.046*** | ||
+| (0.030) | (0.012) | |||
+| Consumption of fixed capitalt-8LDR | 0.091*** | 0.009 | ||
+| (0.025) | (0.010) | |||
+| Hours worked per capitat-8 | 0.000 | 0.000 | ||
+| (0.000) | (0.000) | |||
+| Hours worked per capitat-8TER | -0.000 | 0.000 | ||
+| (0.000) | (0.000) | |||
+| Hours worked per capitat-8LDR | 0.000 | 0.000 | ||
+| (0.000) | (0.000) | |||
+| Compensation of Employeest-8 | 0.125*** | 0.011 | ||
+| (0.037) | (0.016) | |||
+| Compensation of Employeest-8TER | -0.047 | -0.013 | ||
+| (0.058) | (0.024) | |||
+| Compensation of Employeest-8LDR | -0.156*** | -0.005 | ||
+| (0.052) | (0.021) | |||
+| -1.384** | -0.315** | |||
+| (0.633) | (0.157) | |||
 | Observations | 3,408 | 3,408 | 3,408 | 3,408 |
-| R-squared | 0.000 | 0.001 | 0.001 | 0.000 |
-| Notes: Standard errors are in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01 | ||||
+| Kleibergen-Paap rk LM (p-value) | 26.844 (0.000) | 26.844 (0.000) | ||
+| Cragg–Donald Wald F | 12.957 | 12.957 | ||
+| Kleibergen-Paap rk Wald F | 2.676 | 2.676 | ||
+| Hansen J (p-value) | 7.542 (0.274) | 9.196 (0.163) | ||
+| Notes: Robust standard errors clustered at the NUTS 2 level are reported in parentheses. Significance level: ∗p0.1; ∗∗p0.05; ∗∗∗p0.01. | ||||